A Complete Bibliography of Publications in Annales de Physique (Paris, France)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.01

Title word cross-reference

1 [BBL07, Gog94]. 10^{−5} [DDL^{+}07]. 2 [Byc94, PF93, RMT^{+}07, dCP07]. 4 [PF93]. 5d(5/2)_{J=3}^{0} [BSMS92]. ^{∗} [BRG^{+}97, SVDL^{+}94]. ^{+} [HJK^{+}07, KRVV95, NLM^{+}94]. ^{+3} [BDT^{+}95]. ^{127} [RP95b]. ^{129} [RP95a]. ^{2∗} [JLS^{+}95]. ^{2+} [MPT^{+}07]. ^{3+} [HJK^{+}07]. ^{4} [BDT^{+}95]. ^{4+} [NVJ^{+}95]. ^{7+} [LS94]. 8+ [LS94]. _{1−x} [AFF^{+}94]. _{1−x−y} [GGM^{+}97]. _{13/2} [BDT^{+}95]. _{2} [AGF94, BWDR95, FABC88, GCT^{+}97, HJK^{+}07, Kop88, LDCBA92, MRG^{+}88, MMN^{+}94, RP95a, RP95b, TFdFR89, Ver88]. _{3} [GJJR95, MRG^{+}88, Min94, Ver88]. _{4} [AGF94, BHS^{+}94, Kop88, MMN^{+}94, TFdFR89, XZS^{+}07]. _{6} [AGF94]. _{7−δ} [MRG^{+}88]. _{c} [BAGR88, GFC88, OBU^{+}88]. _{x} [AFF^{+}94, GGM^{+}97, Ver88]. _{y} [GGM^{+}97]. A [CMF^{+}95]. A + BC → ABC → AB + C [RB91]. ≈ 10 [MNO^{+}94]. δ [MRG^{+}88]. E||B [ZB90]. γ [LD91]. μ [HST97, NVJ^{+}95]. Q [NBCB^{+}07]. × [GMS^{+}94].

-C [AGF94]. -doped [HJK^{+}07, MPT^{+}07]. -photon [dCP07]. -xMxCuO [Kop88].

/O [LDCBA92].

10 [LRRVT92]. 100 [GMS^{+}94, Sau95]. 10J [GMS^{+}94]. 1960’s [CT85a]. 1987a [LD91].

20 [KEH07]. 200-300 [PV92].

50J [SGZ97].

86-km [LDC^{+}07].

90 [Ano90]. 94’ [Bré94].

a-t-on [AVC92]. Ablation [DAB^{+}94, GGA^{+}97, HTBL94, BCKV97, GGCC97, GGA^{+}94, HFF97, MKM97, VHBL97]. ablation/ionisation [MKM97]. absolus [Tro94]. absorber [DBD07]. Absorption [BDT^{+}95, Dum85, MRG^{+}88, Pet99, Rom95, vB89]. AC [Reu95]. accélérateur [LM00]. Accélérateurs [Lag00]. accélération [Bob86]. accelerator [LM00]. accelerators [Lag00]. accès [Qué92]. Accord [GPL95, dM95]. Achievements [Kam91]. acides [MB97]. ACO [BBS94]. Acoustic [AS85]. acoustiques [DKL86]. activité [SM89]. actuel [ATE92, Dei97, Del00, Sch00]. adaptatif [DM94, DMFR94]. adaptées [Sau92]. adiabatique [BLW94]. adsorption [Cra85, SVS98]. advanced [Fer02]. affinement [GPL95]. AFM [lS08]. agents [BRJ^{+}07]. aggregates [Cal99]. aggregating [BJ88]. agrégats [Cal98, Cal99, MNO^{+}94]. aide [ABB94, BL92, GDA^{+}95, TPD^{+}97]. aimantation [Gen85]. air [DLKP89]. aléatoire [Clé08, Mon95]. aléatoires [DKL86]. Alfred [Sal85, Bén85a, CT85a, Cur85, DG85, ski85]. Algorithmes [Bla03]. Algorithms [Bla03]. algue’ [AGF94]. Alkali [AB84, BTP85, Bou02]. Alkali-rare [AB84]. All-fibered [FFPM07]. Allocution [Leh85]. alpins [Tou85]. alternative [EQ85]. amas [Pie00]. ambiants [VHBL97]. Amélioration [CB95]. améliorations [Del00]. amorcée [LBH92]. amorphe [FHF92, PLP^{+}97, dUF97]. amorphes [MR91a, MR91b]. amorphous [Wya05]. Ampère [Bén85a]. Amplificateur [RB95, MRIC^{+}95]. amplificateurs [GPL95, PCS^{+}95, TLG95]. amplification [JJC84, Sau95]. Amplifications [Bor95]. amplitudes [Gia84]. Analyse [CMF^{+}95, CAG^{+}97, HFF97, Muz95, BSMS92, BDL^{+}92]. Analysis [Dés95, CCH^{+}87, JWG^{+}07]. analytique [Gal85]. ancestrale [SM89]. Anchor [LEPMS00]. anchored [LEPMS00]. Ancrage [LEPMS00]. ancré [LEPMS00]. Anderson [Fle09]. André [Fri85, Tou85]. anesthésié [MLJC97]. angulaires [Van01]. Angular [Ale87, Van01]. anisotrope [Ger04]. anisotropic [Ger04]. anisotropie [PBD^{+}95]. anisotropies [Pas86, PVJ^{+}95]. Annales [Ano90]. anneau [Cou91]. années [Hei89]. antennae [Bri85]. antennas [Vin85]. Antiproton [CV89]. antiquark [BB85]. appariés [AGF94]. appel [ML94]. appelle [Fra91]. applicables [BH92]. Application [BBH95, BBG^{+}85, AAC97, Ben92, Bén85b, CDMC92, Fer02, GGA^{+}94, KLC^{+}97, Lar06, MLCJ94, NLM^{+}94, PCFC07, FOF95, VGG97]. Applications [AEA^{+}94, BH94, Det99, EB94, GMS^{+}94, Mon95, NG94, AA88, Ayr95, BA92, BL86, BCB^{+}95, BR88, Dhe90, Gio95, IMCD97, Kan09, Kli94, Méj05, PNCA94, Rob85, SFE95, Suz84, ZAC^{+}97, dUF97, Bre93]. applied [Bra99, ST84]. appliqué [Bra99]. appliquée [CBOf95, MKM97]. Apport [FMOS97]. apporter [Ren92]. approach [LB02]. Approche [CCJ^{+}94, Bab86, Gia84, LB02]. après [Enc88, KZAK^{+}92]. Ar-SiH [BHS^{+}94]. arc [CFHF94]. architectures [Bla03]. ArF [LDCBA92]. aromatic [Ld89]. array [RMT^{+}07]. art [Ayr95, Dei97, FTGB95]. artificial [LPPC99]. ascension [LB06]. aspect [DM94]. Aspects [Buc85, Cho96, Cil94, Cug96, MDF88, Cug86, Sch00, SV87b]. assisted [LM00]. assistés [LM00]. associé [ALE95]. astrochimiques [RC97]. astrophysics [AR90, AS01, BSB91, Bom91, Des01, SB84]. astrophysique [AS01, Des01, Pec85]. Astrophysiques [Sau94]. asymétrique [SM89]. asymétriques [AGF94]. asymmetric [Hua06]. atmosphere [Méj05, Méj05]. atmospheres [BHWS91, dB84, Ano84]. atmosphérique [Bla85, CAG^{+}97]. Atom [dLMDP07, BJZ^{+}07, Cra85, Dor84, JBTV07, vA08]. atome [Bou85, FSBD95]. Atomes [CT85b, Bou02, Gaë09, Har85, LLC^{+}95, VLL^{+}95]. Atomic [Kle85, SB84, BSB91, Deh85, Esa86, Gal99, Lah06, MV92, RMT^{+}07, WIB^{+}85, CT85a]. atomique [Gal99, Lah06, Lem92, Min95]. atomiques [Bor95, SAD^{+}95, Van85]. atoms [ABS84, BCP^{+}07, Bou02, DZB89, DZB92, Gaë09, MRG^{+}88, Met85, Phi85]. au-dessus [GABG97]. auprès [BBS94]. aurorales [Gen87]. autocompression [GHZ^{+}95]. autoionisants [KSF^{+}94]. autour [CMF^{+}95]. aux [Bal96, BH94, Bob05, BH92, Fer02, ML94]. avalanche [GJJR95]. avancées [Fer02]. avancés [Pet94]. Avant [Ano84, Ran84, SK85, WHPD94]. Avant-propos [Ano84, Ran84, SK85, WHPD94]. avril [Ano97b]. axe [CMF^{+}95].

bande [BSM^{+}97, BWDR95]. bandes [Cas98, MR91a, MR91b]. bandgap [Cas98]. bands [AK88]. barrière [MMN^{+}94, SM89]. barriers [MRC^{+}03]. Baryon [Pet99]. baryonique [Pet99]. Base [Reu00, LM00, Qui95, Tor04]. based [Dav06, LBF^{+}07, Tor04, ZWJ^{+}07]. basée [Bla94]. basées [Dei97]. Basic [LM00, Reu00]. basse [Bar00, KVC^{+}97, Mey84, ZD94]. basses [RC97]. BBO [GHZ^{+}95]. be [RH85]. beams [Deu86]. become [Whi91]. before [PFC91]. belong [LWMG91]. bénéfique [FMOS97]. besoin [AVC92]. between [ABS84, Gil90, Hua06, Kop88]. beyond [Str85]. bi [Lab00, LCSY88]. bi-dimensionnels [Lab00]. Bi-Sr-Ca-Cu-O [LCSY88]. bidimensionnel [Fle09]. bilan [BM89]. binaries [Las90]. bio [Fou92]. bio-cristallographie [Fou92]. biologique [Bén85b]. biologiques [Gal85]. Biology [CLM^{+}97]. biophysics [STC^{+}07]. bioxyde [PDAYS89]. biréfringinces [PBD^{+}95]. bistabilité [FMSZ95]. bistability [DBD07]. Bitter [Str85]. black [Mey91]. Blandin [Fri85, Tou85]. Blériot [EB94]. bleu [Dei97, FTGB95, GJJR95]. blocage [CB95, Gaë09]. Boltzmann [DDL^{+}07]. bonne [FMOS97]. bore [AAC97]. Bose [Clé08, Ger04, Sch03, vA08]. Bragg [AB94, IMCD97, SLI^{+}94]. Breakage [HJK^{+}07]. bref [Min95]. Broadening [GSH^{+}84]. Brossel [Str85]. Brownian [Mon95]. brownien [Mon95]. bruit [GF95, Rey90, Tor02]. bubble [LB06]. bubbles [LB02]. Building [BDLM09]. bulle [LB06]. bulles [LB02]. bursting [LB06].

C [AGF94, GGM^{+}97, HHP89]. C1 [Ano92, Ano94, Ano97b, Ano90]. C1-90 [Ano90]. Ca [LCSY88, KRVV95]. cadence [LBB^{+}94, VKCP97]. CaF [HJK^{+}07]. calcium [MPT^{+}07, dSLR^{+}07]. Calcul [Bra99]. calculation [ST84]. Calculs [RP95a]. ‘calibration’ [Ben92]. caméra [IS94]. Can [RH85]. canonical [Diu90]. capillaires [Pou97, ZD94]. capteurs [Men94]. Caractérisation [BBF^{+}95, CPDMR94, FOF95, MKG^{+}92, MMN^{+}94, PNCA94, PV92]. caractéristiques [LM00, WCB92]. carbon [KBW^{+}07, Kro89, Ryt91]. Carbonaceous [Ryt91]. carbone [GGA^{+}94, GGM^{+}97]. Carcinogenesis [Ave97a]. cardiologie [Com92]. cas [PDAYS89]. Cascades [GHZ^{+}95]. case [FABC88, Fri92]. cathode [LBS^{+}92]. cavité [Har85, Maa97, MVCF95, NNH95, Rom95]. Cavités [SAD^{+}95, PVJ^{+}95]. cavities [BBDC07, VCB^{+}07]. cavity [Maa97]. CBr [AGF94]. CCD [Men94]. Cd [SKGRF95]. CdTe [BBF^{+}95]. CdZnTe [GBC^{+}07]. CEA [SA97]. CEA-DAM [SA97]. cell [Pat96]. cells [BRJ^{+}07]. cellule [Pat96]. cellules [BBD95]. celui [Hul95]. Centre [Sal85, EB94]. centres [ZWJ^{+}07]. céramique [SFB94]. céramiques [ASP94, PNCA94]. certaines [Mon95]. cerveau [MLCJ94]. ces [Fra91]. CH [MMN^{+}94, TFdFR89]. chaîne [Bla94]. chaînes [Bob05]. Chains [Bob05]. champ [CVR^{+}95, Dar84, LSB94, Maa97, Mey03, TVW^{+}95]. champagne [LB02, LB06, LB02]. champs [Bor95]. Chaos [BLW94, Cil94, JCL07, RB91]. characteristics [LM00]. Characterization [PRG94]. charge [BAGR88, Bon84, Kop88]. charged [Bon84]. Charles [Bou96]. chauds [AA91, CGST92]. Chemical [GZH89, AAB^{+}85]. chez [MLJC97]. Chimie [Rau89]. chimique [BM89, PDAYS89]. chimiques [LB02, VL89]. chip [vA08]. chlore [PDAYS89]. chlorine [EQ85]. choc [OSA94]. choice [Jac07, JWG^{+}07]. choix [Jac07]. chromium [BCP^{+}07]. cible [AAC97, CCJ^{+}94, GGCC97, GGA^{+}94, HTBL94, CDMC92]. cibles [AGG^{+}94, Fré00, HFF97, VHBL97]. Cinétique [BSMS92, Van89, PDAYS89, SM97]. Cl [AGF94]. classical [Ste85]. classique [Ala86]. clustering [BJ88]. clusters [BJ88, Cal98, Pie00, TFdFR89]. CNRS [FDSF92]. CO [MMN^{+}94, TFdFR89]. codes [Fré00]. coefficients [Reu00, Vol97e]. coexistence [Gul04]. coherence [AHO92, Det99, Ger04, Det99, Gio95, HAC^{+}95b, Ger04]. Coherent [GBC^{+}07, Our07, PCFC07, Lem92, Pou97]. cohérente [KCK^{+}92, LB92]. cohérentes [FDSF92, Our07]. col [Bré94]. collaboration [Str85]. collage [LSD^{+}97]. collectifs [Gia84]. Collective [Cho96]. collectives [Mey84]. Collision [Blo85, Ami85, Bal96, DC87, Gry85]. Collision-induced [Blo85]. Collisional [MPW84]. collisionnel [FS92]. Collisions [Bab86, Cug86, Cug96, BSB91, Bon84, BRTG90, Bor92, DZB92, Qué92, Sch84]. Colloque [Ano90]. COLOQ’10 [DG07]. colour [ZWJ^{+}07]. combinée [WCB92]. combs [MSS^{+}07]. cométaire [Enc88]. cométaires [BM89]. comète [Enc88]. comets [GZH89]. comme [BSM^{+}97]. Commun [EB94]. communication [JCL07]. communications [CdBB^{+}99]. compact [VKCP97]. compacte [RCK^{+}94]. comparison [Mot85]. Competition [Kop88]. composante [Ala86]. composants [FMOS97, MB97]. composites [ASD93]. composition [BM89]. compounds [Jus88, Ryt91]. compression [FGD07]. comprimée [SRF95]. comptage [Cub88]. computation [Bra99]. computers [Bla03]. concept [Dar84]. concepts [Gry85, MR91a]. concernant [RP95b]. Condensat [Ger04, Clé08]. condensate [Clé08, Ger04, Sch03]. conductance [Reu95]. conducteur [NNH95, VPF95]. conducteurs [LFTV95, MR91a, MR91b]. conduction [ML94]. configuration [dM95]. configurations [Cin88]. confinée [OSA94]. confinement [Deu86]. confocal [Bra99]. confocale [Bra99]. conforme [Gra92]. Congrès [Bro85]. Congress [Bro85]. Conjugaison [Bri98]. conjugation [Bri98]. conjuguée [NNH95]. connecté [Saf97]. connected [Saf97]. conservatif [RB91]. consistent [Ale87]. constant [DDL^{+}07]. constantes [RP95a, RP95b]. Construction [BDLM09]. containing [DBD07, TFdFR89]. content [Pet99]. contenu [Pet99]. Continu [PC92, Gal99, KDC^{+}95]. Continuous [GR94, Gal99, LBF^{+}07]. continuum [BSMS92]. Contrôle [MMZ^{+}97, SFB94, Hul95, SFE95]. contraintes [Cla99, GGM^{+}97]. Control [AAB^{+}85, GBC^{+}07]. Controlling [Gay01]. convection [Abi94]. conventional [MLT85]. Conversion [Boc92, GJJR95, Lem92, Wyo92]. cooling [Lah06, Phi85]. copper [Jef88, MRG^{+}88]. core [LDR07]. corpuscle [Jac07]. corpuscule [Jac07]. Correction [AFH94]. corrections [PT88]. Correlated [AK88, Bob05]. Correlated-electron [AK88]. correlating [KBW^{+}07]. correlation [BPJ^{+}07, BRJ^{+}07, FF99, PTC99, Vol97e, PTC99, Qué92, Vol97e, DBG89]. Corrélations [SB88, Jef88, JDS^{+}95]. corrélés [Bob05]. cosmology [Not88]. Couches [AFF^{+}94, AAC97, CAA^{+}94, CDMC92, FHF92, GGA^{+}94, NFV^{+}94, dUF97]. couleurs’ [KSF^{+}94]. Coulomb [Fle09]. coulombienne [Fle09]. Couplage [KDC^{+}95, HSD^{+}94, KSF^{+}94]. Coupled [Fau94]. couplés [Lef90]. coupling [VCB^{+}07]. Couplonic [BBL07]. courte [SB88]. courtes [Rib97, SFE95]. Cr [NVJ^{+}95]. Crack [Pon07]. cratère [GGA^{+}97]. créé [LSB94]. créée [CPDMR94, KVC^{+}97]. creep [Ovc94]. créés [Ben92, Gau92, HFF97, NG94, Sau92]. creuse [LBS^{+}92]. Cristal [DM94, DMFR94, GJJR95]. Cristallisation [PLP^{+}97]. cristallographie [Fou92]. cristaux [BBS94, BDT^{+}95, CMF^{+}95, Lab00, LBFdR94, Tor04, Vil03]. critical [LM00, OBU^{+}88]. critique [DBG89]. critiques [LM00]. Croissance [Pel04]. crossed [BCP^{+}07]. Cryogenic [lS08]. cryoprotecteurs [VP85]. crystal [BBL07, DSS87, Kam91, LDR07]. crystal-laser [Kam91]. crystals [Dav06, Kat94, Lab00, Tor04, Vil03]. CS [BWDR95]. Cu [LCSY88, MRG^{+}88, Ver88]. cubiques [CGM87]. cuivre [EB97, GGCC97]. Cummings [AHOE95]. CuO [FABC88]. cuprate [Bob05]. cuprates [Bob05]. Current [Ber85, Del00, Hua06, Sch00]. Curves [Mon95]. CVD [ZWJ^{+}07]. CVD-grown [ZWJ^{+}07]. cycle [Del00, FFPM07, MLCJ94, Sch00].

D [BBL07, Byc94, Gog94, RMT^{+}07]. D.B.D. [MHN^{+}97]. D.R.A.S.C. [SL95]. d’ablation [AFF^{+}94]. d’absorption [PDGM97, Pet99, Rom95]. d’AEROSPATIALE [EB94]. d’alcalin [DVT^{+}97]. d’alcalins [Bou02]. d’Alfred [Rob85]. d’alimentation [PPEZ97]. d’alliages [AGF94]. DAM [SA97]. d’Anderson [DKL86, Fle09]. d’anneaux [Reu95]. dans [Abi94, AA91, AFH94, ACG^{+}97, BWDR95, Ben92, Boc92, BLW94, Bou02, Bou85, BDT^{+}95, Bou00, Bri98, CGST92, Cal98, Cal99, Cam99, CGM87, Cla99, Clé08, CT85b, Dan85, Dei97, DKL86, DBG89, FSBD95, Fle09, FDSF92, GHZ^{+}95, GMS^{+}97, GDA^{+}95, GABG97, Ger04, GF95, Gia84, GGM^{+}97, GJJR95, Har85, HSD^{+}94, HAC^{+}95b, KRL^{+}95, KSF^{+}94, KVC^{+}97, LS96, LDCBA92, LLC^{+}95, LFTV95, LBH92, LSD^{+}97, Lem92, LL85, LEPMS00, Maa97, Maz89, Méj05, Miq97, MR91a, MR91b, ML94, MKG^{+}92, MHN^{+}97, Nau02, Nen92, NLNA94, Pas86, Pec85, Per05, PDGM97, Pet99, Pet94, Pie01, PJ97, PC92, PV92, PBD89, Que84, Rab94, Rau89, Reu00, RAR^{+}95, RKCP97, RS90, Sch92, SCSN93, SV87a, SV87b, SM89, SVDL^{+}94, Tor02, Tor04, Tro94, VHBL97]. d’argon [RMA94]. data [STC^{+}07]. d’atomes [Bor95, HAC^{+}95b, JDS^{+}95, PJ85, RAR^{+}95]. d’avancement [Ste92]. d’azote [HTBL94]. décharge [BE92, FHF92, KVC^{+}97, LDCBA92, MKG^{+}92, MMN^{+}94, PP92a, PPEZ97, PP92b, SCR^{+}94]. décharges [Pou97, ZD94]. d’échelle [FMSZ95]. déchets [Ler00, Sch00]. decipher [Pon07]. déclenchés [PBL^{+}94]. deduced [DGLR90]. d’effets [GHZ^{+}95]. Défis [Mat93]. deformation [Gal99]. déformations [Gal99]. degrees [Mat91]. delay [JCL07]. delayed [Jac07, JWG^{+}07]. delayed-choice [JWG^{+}07]. delays [TD85]. d’électrons [ABB94, Fle09]. delta [RS90]. d’émission [GFCB99, MLJC97, PC92, SVDL^{+}94]. demolition [PRG94]. Demonstration [Cal99]. d’énergie [RKCP97]. dense [GGP84, LMpk01, Rei05]. densité [AFH94]. densities [SF84]. density [BSB91, IFRS94]. Dépôt [BCKV97, CAA^{+}94, RKCP97, FHF92]. dépendance [SKGRF95]. dependence [Ber87]. dependent [Sch85]. dépolarisée [Que84]. depolarizing [BSB91]. déposées [AFF^{+}94]. deposition [SVS98]. deposition/adsorption [SVS98]. dépots [NFV^{+}94]. d’équilibre [Ala86]. Derivation [BSB91, Bom91]. Description [Cla99, LEPMS00, LEPMS00, SV87a]. désordonnés [LS96, ML94]. dessus [GABG97]. destructives [GGPR95]. detachment [Esa86]. d’état [Sur91]. d’états [AGF94, KSF^{+}94, Our07]. Detection [BHS^{+}94, LDCBA92, BBDC07, Bre93, Per05, Bor85, Per05]. determination [MRG^{+}88]. determining [EQ85]. d’étoiles [Lum85]. détonation [DBG89, DLKP89, PBD89, VL89]. detonations [DWD89, PDAYS89, Van89]. detuning [AHO92]. deutérium [AFH94]. deux [Ala86, DNB^{+}95, Gaë09, GNMS^{+}95, JLS^{+}95, Let95, WCB92, KSF^{+}94]. deuxième [BSMS92]. development [Fri91]. developments [Des01, SB84]. développement [Bén85a]. Développements [Des01, Miq97]. devices [Bou00]. d’excitation [SV87a, SV87b]. d’harmoniques [ACG^{+}97, LB92, LSB94]. d’hydrodynamique [Ano84]. diagnostic [CFHF94, SGFE97]. diagnostics [BDL^{+}92, Tro02]. diagram [AGTS91]. diamètre [DBG89]. Diamond [FSC^{+}07]. diatomic [Ber84]. Dielectric [Rei05, Fer02]. diélectrique [Ber88, Fer02, FSBD95, MMN^{+}94]. diélectriques [NFV^{+}94]. Different [Sch00, BRJ^{+}07]. Différentes [VPF95, VHBL97]. differents [Sch00, AGG^{+}94, VHBL97]. Diffraction [KLC^{+}97, KBW^{+}07]. diffuse [VLL^{+}95]. Diffusion [Mic91, BWDR95, Can04, Mon95, Que84, Vol97e, Vol97e]. d’images [PTC99]. dimension [Let95]. Dimensional [Cou96a, Cou96b, Mon95, NF94, FABC88, Fle09, Lab00, vA08]. dimensionnels [Lab00]. dimensions [Ala86, Let95]. dimères [Bou02]. dimers [Bou02]. d’impulsions [GDA^{+}95, GARG94, MRIC^{+}95, Méj05, Sau95]. d’indice [VPF95]. d’installation [SA97]. d’intensité [JDS^{+}95]. d’intensités [Qué92]. d’interaction [SGFE97, Tro02]. d’interface [PTZR95]. d’interpolation [RP95b]. d’introduction [Gra92]. d’iode [RP95a, CB95, GB95, RB95]. diodes [Dei97, MVCF95]. d’ionisation [Gau92]. d’ions [Cug86, KRVV95]. dipole [BCP^{+}07]. Direct [BLTH94, Maa97]. directe [TBLH^{+}94]. directionnelle [AGF94]. discharge [BHS^{+}94]. discreet [Gal99]. discret [Gal99]. disjoncteur [CFHF94]. disk [Bra99]. Disorder [Cho96, Jus88]. disordered [LS96, Pon07]. d’isospin [GG87]. dispositifs [Bou00, Cub88]. disque [Bra99]. dissemination [LDC^{+}07]. dissipative [BRTG90]. distances [BGP84]. distribution [Diu90, LBF^{+}07, Vol97e]. distributions [Ale87, Muz95]. divers [DLKP89]. Dix [Hei89]. do [Whi91]. d’observations [BM89]. Does [LWMG91]. domain [BRTG90, Sch92]. domaine [Ben92, FDSF92, GDA^{+}95, LSD^{+}97, Miq97, Nen92, Pet94, PV92, Sch92, Tro94]. d’onde [PBD89]. d’ondes [Bou02, Bri98, KDC^{+}95]. dopage [MDF88]. dopé [Bla94, GJJR95]. doped [HJK^{+}07, MPT^{+}07]. dopés [BDT^{+}95, SKGRF95, VPF95]. Doppler [BBD95, DNB^{+}95]. d’or [TVW^{+}95]. d’ordre [ACG^{+}97, LB92]. d’ordres [MMZ^{+}97]. Dorsalis [MLJC97]. doublage [MRIC^{+}95]. double [Ser85, zP85]. doubled [BMV^{+}07]. d’ouverture [Leh85]. down [Rom95]. d’oxydation [Pet94]. d’oxyde [FHF92]. driven [Deu86]. driving [Sch84]. dualité [Jac07]. duality [Jac07]. due [DZB89]. d’un [BSM^{+}97, BBC^{+}95a, BM89, BDLM09, Bra99, CAG^{+}97, CFHF94, Clé08, CB95, JDS^{+}95, KRVV95, KLC^{+}97, LB92, LSB94, Lah06, LDCBA92, LEPMS00, MRIC^{+}95, MMN^{+}94, NNH95, RMA94, Saf97, SGZ97]. d’une [Bla94, BLD^{+}95, CPDMR94, GGCC97, GGA^{+}94, HTBL94, KVC^{+}97, LB06, MMK94, MKG^{+}92, Per05, PPEZ97, VL89]. dur [GGA^{+}94]. durant [SFB94]. Durées [BRG^{+}97]. during [BHWS91]. dust [GZH89]. duty [FFPM07]. duty-cycle [FFPM07]. d’UV [PP92b]. d’UV-X [PP92b]. dwarfs [Mey91]. Dye [BPJ^{+}07]. Dye-substrate [BPJ^{+}07]. dynamic [Clé08, QdR98]. dynamical [IFRS94]. Dynamics [Dor84, Fau94, HR92, Rig95, Bou02, CCH^{+}87, Cal98, Cha00, Gay01, JCL07]. Dynamique [Bou02, BLD^{+}95, Cal98, Cha00, Gau92, Lef90, MMK94, RB91, TLG95, KZAK^{+}92, Mat93, Qué92, QdR98]. Dynamiques [Rab94, Clé08, FMSZ95].

E.N.S. [Roc85]. early [BHWS91]. Earthquakes [GR94]. eau [Bén85b, VP85]. eau-cryoprotecteurs [VP85]. échelle [BL92, BGP84, Chi92, Maz85, Pas86]. éclair [MB97]. éclairs [Vit92]. éclatement [LB06]. École [CT85a]. écoulement [Abi94, Pet91]. Éditeurs [Ave97b]. Effect [AHO92, BRJ^{+}07, BP87, Ber87, Hua06, JWBB99, Fle09]. Effective [Mey03]. effectives [Mey03]. Effects [MN84, TD85, Voi97, Ami85, BdR91, Bre93, DW85, Jus88, LR87, VCB^{+}07]. effervescence [LB02]. Effet [Fle09, HSD^{+}94, AFH94, DNB^{+}95, JLS^{+}95, Noz85]. Effets [FMSZ95, Lum85, KZAK^{+}92]. efficace [Bra99]. Efficiency [dCP07, Dav06, NBHC^{+}07]. Efficient [Bra99]. effondrement [Gou93]. eigenmodes [Bre93]. eigenstates [Bre93]. Einstein [Clé08, Ger04, Sch03]. Elaboration [BCB^{+}95]. élargissement [NLNA94]. Elastic [ARC96, GR94, DZB89, Vol97b, VGM97, Vol97c, Vol97d, Vol97e, VGG97]. élastique [Vol97b, VGM97, Vol97c, Vol97d, Vol97e, VGG97]. Elasto [Vin85, dC85]. elasto-optical [dC85]. électrique [PP92a, PP92b]. électriques [AFH94]. électroactifs [MDF88]. electrodynamic [DW85]. electrodynamics [Ste85]. Electrodynamique [Har85]. electrolyte [KTB90]. Electromagnetic [PBG^{+}84, dC85]. Electron [ABS84, Esa86, Jef88, AK88, Byc94, Cou94, Deh85, Fle09, Hug85, KBW^{+}07, Pie01, Pie01, Pie01]. électron-électron [Pie01]. Electronic [MRG^{+}88, Kle85, Per05]. electronics [Bou00]. électronique [Bou00, BBF^{+}95]. électroniques [MR91a, MR91b, ML94, MFG^{+}97, Per05]. electrons [Cal98, Cam99, Hua06, RMR88, Cal98, Cam99, Cou91, Hul95]. Electrotechnical [BR88]. Élements [Gra92]. élevé [ACG^{+}97, LB92]. élevée [VKCP97]. élévés [MMZ^{+}97]. elliptique [LPvWR95]. EMC [BP87, Ber87]. Emission [ZD94, BMV^{+}07, Bon84, BBDC07, GFCB99, Gay01, GBC^{+}07, HHP89, JJC84, Jol93, XZS^{+}07, Bor95, Gau92, GPL95, KCK^{+}92, LS94, LRRVT92, RMA94, TLG95, WBC^{+}94]. émissions [Gen87]. emprisonnement [GMS^{+}97]. endoénergétiques [BRG^{+}97]. énergétique [BDLM09]. énergie [Bab86, Bar00, Cug86, IS94, Mey84, MBIA92, PDGM97]. energies [DZB92]. energy [Bar00, Bon84, BRTG90, Bor92, BDLM09, CG84, GT86, PCFC07]. enhanced [LR87]. enhancement [PCFC07]. entangled [Our07]. Entanglement [Gaë09]. entre [AAC97, Bab86, BWDR95, FSBD95, LRRVT92]. environment [Per05]. environments [Bri98]. environnement [Pet94]. environnements [Rau89]. epitaxial [BdR91]. époque [Roc85]. épreuve [Dep87, Jac07]. equation [BSB91, Bom91, LMpk01, Sur91]. Equations [NF94, Sur96, Bal96, Vol97e, Bal96, Vol97e]. Equilibre [Gou93, Bal96]. equilibrium [AAB^{+}85, Bal96]. Errata [Ano91a, Ano91b]. Erratum [Ano90, Cla84a, RB91]. establishing [DWD89]. Étalonnage [BBS94, LBFdR94]. Étalonnages [Tro94]. étalons [Van85]. Étapes [Laf91a]. état [Ayr95, BSMS92, BDT^{+}95, Dei97, FTGB95, FSBD95, LPvWR95, ATE92, Ste92]. états [Eph84, Fra91, BRB^{+}95]. étés [Tou85]. étirées [Sou94]. étoiles [Gou93]. étroite [BSM^{+}97]. étude [BBG^{+}85, CDMC92, CGM87, Eph84, MKM97, MLCJ94, NG94, Sau92, MR91b, AAC97, GMS^{+}97, GGCC97, GGA^{+}94, JDS^{+}95, KZAK^{+}92, KVC^{+}97, Lab00, Let95, MRIC^{+}95, Mon95, NVJ^{+}95, NA97, Our07, PBD^{+}95, PBL^{+}94, Per05, PPEZ97, PV92, PBD89, RC97, SV87a, SV87b, TBLH^{+}94, VHBL97]. Études [Abi94, VP85]. étudiées [Del00]. EU205 [Ste92]. Eurêka [Ste92]. évanscente [HAC^{+}95b]. evaporated [Ale87]. évaporation [Lah06]. Evaporative [Lah06]. evidenced [BPJ^{+}07]. evolution [GZH89, Laf91b, Mic91, VL89, Laf91a]. evolved [DNQR91]. Exaltation [Tor04]. exchange [Bon84, Hap85]. excimère [BL92, GMS^{+}94, HSD^{+}94, JLS^{+}95, PBL^{+}94, SFB94, SGZ97]. Excimères [DVT^{+}97, ASP94, ATE92, AEA^{+}94, BA92, BE92, DC92, Dam94, EB97, GGCC97, GGM^{+}97, HFF97, LBB^{+}94, LSD^{+}97, LBDM94, Miq97, PLP^{+}97, Ste92, dUF97]. exciplexe [BRG^{+}97, MBIA92, SVDL^{+}94]. Excitation [DZB92, JLS^{+}95, KZAK^{+}92, KBW^{+}07, MFM^{+}07b, MLJC97, SM97, KSF^{+}94, Tor04]. Excitations [Mey84, BAGR88, Koh85]. excité [BSMS92, BDT^{+}95, FSBD95]. excited [ABS84]. excités [BE92, CRKP94]. Exciton [CGM87]. Excitonic [BGC^{+}07]. Excitons [Com95]. exemples [ASD93, BH92]. exoergic [Dor84]. expansion [AAC97]. Expérience [KSF^{+}94, Dep87, Jac07]. expériences [BDB^{+}97, PP85, Tro02, RS90]. experiment [Jac07, JWG^{+}07, Owe86]. Experimental [JWG^{+}07, Bre93, Our07, VGM97, Vol97d, Lab00, Let95]. expérimentale [AGF94, Bab86, Eph84, Lab00, Let95, Our07, PBL^{+}94]. expérimentaux [MDF88, SV87b, VGM97, Vol97d]. Experiments [GR94, MRG^{+}88, Tro02]. exploration [Enc88]. explosif [PBD89]. explosifs [DLKP89]. explosive [Lum85]. exponentielles [Sou94]. extensive [Vol97b, VGM97, Vol97c, Vol97d, Vol97e, VGG97]. extra [MVCF95]. extra-cavité [MVCF95]. extraction [Dav06, Sch00]. extragalactiques [BGP84]. extraterrestres [Hei89].

fabrication [Dhe90, Qui95]. Fabry [PVJ^{+}95, Bou96, TD85]. Fabry-Perot [PVJ^{+}95]. Facteurs [LBS^{+}92]. faible [IS94]. faibles [PVJ^{+}95]. faisant [ML94]. faisceau [CT85b]. faisceaux [MVCF95]. fascinant [MMK94]. Fe [Jol93]. feedback [MJM^{+}07]. femtosecond [Méj05, Cam99]. femtoseconde [Bla94, CMF^{+}95, Cam99, GHZ^{+}95, GRA^{+}94, GABG97, Hac95a]. femtosecondes [MRIC^{+}95, Méj05, RMA94, Sal92, Sau95]. Fermi [BRTG90, Bor92, LMpk01, Sch03, TVW^{+}95]. ferromagnetic [Mél97]. ferromagnétiques [Mél97]. février [Ano97b]. fibered [FFPM07]. fibers [QdR98]. fibre [LBF^{+}07, LDC^{+}07]. fibres [IS94, LDR07, QdR98]. field [BSB91, Bom91, GFC88, JWBB99, Maa97, Mey03, OBU^{+}88, Sch84]. fil [Saf97]. filled [Byc94]. films [AGF94, BdR91, BCKV97, GGM^{+}97, TBLH^{+}94]. fils [Pie01, Saf97, Tor02]. filters [FF99]. Filtrage [NNH95]. fine [vB89]. finesse [PVJ^{+}95, VCB^{+}07]. finite [DW85]. first [LB02, VGM97, Vol97d]. Fission [Bar00, Ale87, HR92]. flash [CRKP94, LBH92, PV92, RCK^{+}94]. fluctuations [Hua06, IFRS94]. fluides [SV87a, SV87b]. fluids [And84]. Fluorescence [CRKP94, JLS^{+}95, MLCJ94, BPJ^{+}07, BRJ^{+}07, HSD^{+}94, LDCBA92, MPT^{+}07, MJM^{+}07]. fluoride [dSLR^{+}07]. Flux [DY88]. following [Bon84]. Fonction [Ber88, PPEZ97]. fonctionnelles [Mon95]. fonctions [Qué92, Vol97e]. fond [GGA^{+}97]. fondamentaux [Eph84, MR91a, PDGM97]. forêt [Cha00]. force [LLC^{+}95]. Foreword [Laf91b]. formant [VGG97]. formation [Kro89, Maz85, Mey91]. forme [Pel04, AAC97]. Forming [NF94, VGG97]. formules [RP95b]. forte [LBB^{+}94, SBE92, SGZ97]. fortes [Ste92]. four [Blo85, Dor84]. four-wave [Blo85]. fractales [Muz95]. fracture [Pon07]. fragment [HHP89]. Fragmentation [BRTG90]. fragments [Ale87]. Franco [Bré94]. Franco-Russe [Bré94]. Free [Cou94, Deh85, RMR88]. freedom [Mat91]. French [Ano97b, Abi94, AGF94, ASD93, AA91, Ala86, ALE95, AB84, AVC92, AFH94, AB94, ABB94, AGG^{+}94, AAC97, Ano84, Ano90, ACG^{+}97, AFF^{+}94, AS01, ASP94, Ave97b, ATE92, AEA^{+}94, Ayr95, Bab86, BSM^{+}97, Bal96, BH94, BWDR95, BBC^{+}95a, Bar00, BBH95, BA92, BBS94, BBC^{+}95b, BCKV97, Ben92, Bén85a, Bén85b, BSMS92, Ber88, BL86, BBG^{+}85, Bla03, Bla85, Bla94, Bob86, Bob05, Boc92, BM89, BH92, BCB^{+}95, BLW94, BL92, Bor85, Bor95, BGP84, Bou96, Bou02, Bou85, BDT^{+}95, BDLM09, BDL^{+}92, BRG^{+}97, BDB^{+}97, Bou00, Bra99, BE92, Bre97, BBF^{+}95, Bré94, BBD95, Bri98, Bro85, BRB^{+}95, BLD^{+}95, CCJ^{+}94, CRKP94, Cag85, CGST92, Cal98, CMF^{+}95, Cal99, Cam99, Can04, CBOf95, Cas98, Cas94, CAA^{+}94, Cav00b, Cav00a, CAG^{+}97, CDMC92, Cha00, CGM87, CPDMR94]. French [Chi92, CVR^{+}95, CFHF94, Cla99, Clé08, CT85b, Com92, CB95, Cou91, Cou96a, Cou96b, Cub88, Cug86, Cur85, DC92, Dam94, Dan85, Dar84, Dav92, DAB^{+}94, Dei97, DVT^{+}97, Del00, DKL86, Dep87, DBG89, Des01, Det99, DM94, Dum85, DLKP89, DNB^{+}95, Enc88, Eph84, EB94, EB97, FOF95, FGVV89, FTGB95, Fer02, FMSZ95, FSBD95, Fle09, FMOS97, FS92, FDSF92, Fou92, Fra91, Fré00, FHF92, Gaë09, GG87, Gal99, Gal85, GHZ^{+}95, GMS^{+}97, GDA^{+}95, GFCB99, GGA^{+}97, GGCC97, Gau92, GRA^{+}94, GABG97, GARG94, Gen85, Gen87, Ger04, GGA^{+}94, GF95, Gia84, Gio95, GPL95, GMS^{+}94, GNMS^{+}95, GSH^{+}84, Gou93, Gra92, GGM^{+}97, Gry85, GJJR95, GB95, Hac95a, HG94, Har85, Hei89, HSD^{+}94, HFF97, HAC^{+}95b, Hen94, HST97, HTBL94, Hul95, IMCD97, IS94, Jac07]. French [JLS^{+}95, Joy92, JDS^{+}95, KRL^{+}95, KZAK^{+}92, Kan09, KSF^{+}94, KDC^{+}95, KCK^{+}92, KVC^{+}97, Kli94, KRVV95, KLC^{+}97, LB92, LSB94, Lab00, LBB^{+}94, LS96, Laf91a, Lag00, Lah06, LDCBA92, LS94, LLC^{+}95, LFTV95, LBH92, LBS^{+}92, LRRVT92, LSD^{+}97, Lav89, LBDM94, Lef90, LPvWR95, LBFdR94, Leh85, Lem92, Ler00, Let95, LL85, LB02, LB06, LM00, Lor98, Lum85, LEPMS00, Maa97, MRIC^{+}95, Mat93, Maz85, Maz89, Méj05, Mél97, MVCF95, Men94, Mét00, MMZ^{+}97, Mey84, Mey03, MB97, MKM97, Min95, Miq97, MMK94, MDF88, MR91a, MR91b, ML94, MBIA92, Mon95, MFG^{+}97, MKG^{+}92, MMN^{+}94, MHN^{+}97, MLCJ94, MLJC97, MNO^{+}94, Muz95, Nau02, NNH95, Nen92, Nen97, NLNA94, NVJ^{+}95, NG94, NA97, NFV^{+}94, Noz85, OSA94, Our07, PDAYS89, PCS^{+}95, PBD^{+}95, Pas86, PBL^{+}94, Pat96, PTC99, PP85]. French [Pec85, Pel04, PF93, PTZR95, PP92a, Per05, Pet91, PDGM97, Pet99, Pet94, Pie00, Pie01, PJ85, PVJ^{+}95, PJ97, Pou92, PPEZ97, PC92, PV92, PP92b, Pou97, PNCA94, PBD89, PLP^{+}97, Qué92, Que84, QdR98, Qui95, Rab94, Rau89, RB91, RP95a, RP95b, Ren92, Reu95, Reu00, RMA94, Rey90, Rib97, RAR^{+}95, RCK^{+}94, RKCP97, Rob85, Roc85, RB95, Rom95, Rou92, RC97, RS90, Saf97, SM97, SBE92, Sau92, Sau95, Sau94, SKGRF95, Sch00, SL95, Sch92, SA97, SRF95, Sen03, SCR^{+}94, SCSN93, SB88, SV87a, SV87b, Sou94, SLI^{+}94, SM89, SFB94, SFE95, SGFE97, Ste92, SGZ97, Str85, SVDL^{+}94, Sur91, SPJP92, SAD^{+}95, TLG95, TBLH^{+}94, VL89, TVW^{+}95, Tor02, Tor04, Tos85, Tou85, Tro94, TPD^{+}97, Tro02, Van89, VPF95, Van85, VLL^{+}95]. French [VP85, VKCP97, Vil03, Vit92, VHBL97, Vol97a, Vol97b, VGM97, Vol97c, Vol97d, Vol97e, VGG97, WBC^{+}94, WCB92, Wyo92, ZAC^{+}97, ZD94, dM95, dUF97, Van01]. fréquance [BWDR95]. fréquence [BBC^{+}95a, Lem92, MRIC^{+}95, PTZR95, Van85, Wyo92]. fréquence-somme [PTZR95]. Frequency [LDC^{+}07, CHH^{+}07, FDS94, MSS^{+}07]. Fresnel [AB94, IMCD97, SLI^{+}94, TPD^{+}97]. froids [JDS^{+}95]. front [DBG89, PBD89]. frontière [MMK94]. frustration [HSD^{+}94]. fullerènes [MKM97]. function [MRG^{+}88]. Functionals [Mon95]. functions [Ber87, Vol97e]. Fundamental [Diu90, LDC^{+}07]. fusion [Deu86, dUF97]. futur [Nen92]. future [Hap85]. futures [Hei89].

Gabor [JP99]. galaxies [Pie00]. gallery [NBCB^{+}07]. gamma [Ham85]. GaN [BGC^{+}07, Dav06]. GaN-based [Dav06]. gas [AB84, DZB89, Fle09, vA08]. gases [Cla84a, Cla84b, Sch03]. Gaussian [Our07]. gaussiennes [KDC^{+}95]. gaussiens [Our07]. gaz [ACG^{+}97, CRKP94, DVT^{+}97, DBG89, Fle09, GMS^{+}97, KVC^{+}97, LBH92, Lem92, MNO^{+}94, RKCP97, SM97, VHBL97, ZD94]. gazeux [DLKP89, PBD89]. Ge [AFF^{+}94, GGM^{+}97, HFF97]. géant [Lum85]. géantes [AA91, Fra91]. Gels [BCB^{+}95]. générale [Gou93, Vol97a]. Généralités [Hac95a]. generated [BHS^{+}94]. générateur [VKCP97]. Générateurs [PCS^{+}95]. Generation [GCT^{+}97, ACG^{+}97, CGST92, Lem92, PP92b, Pou97, Sau95, SCR^{+}94, Tor04, dC85, LB92, LSB94, PTZR95, Tor04]. genèse [Dar84]. géométrie [OSA94]. geophysical [And84]. giant [Vil03]. giant-spin [Vil03]. giants [Whi91]. Glass [Rig95, VGG97]. glasses [Col85, GFC88, Sou85]. Glassy [Mél97]. grain [Kro89]. grand [CMF^{+}95, Vil03]. grande [BL92, LBB^{+}94, Maz85, MBIA92, Pas86, PVJ^{+}95]. grandes [Gia84, Maz85]. Grands [Dep87]. granulaires [Cla99]. granular [Cla99]. graphite [GGA^{+}94]. Gravitational [Not88, BDT85, Bri85, dC85]. gravitationnel [Gou93]. gravitationnelles [Bor85]. gravure [BDB^{+}97, DAB^{+}94]. GRECO [GGP84]. group [Can04, CT85a]. groupe [Can04, FDSF92]. Groupement [Bén85a]. grown [ZWJ^{+}07]. growth [DSS87, Pel04]. guide [MMdC^{+}07, dCP07, Lah06]. Guided [BJZ^{+}07, Lah06]. GW [LRRVT92]. gyrometry [Bre93]. gyroscope [Sch08].

H [Fre85, TFdFR89]. half [Byc94]. half-filled [Byc94]. Halley [Enc88]. Hanle [Blo85]. Harmonic [GCT^{+}97, XZS^{+}07]. harmonics [Tor04]. harmonique [Tor04]. harmoniques [MMZ^{+}97]. haut [RCK^{+}94]. haute [AFH94, Bab86, Bla94, CRKP94, Cug86, HG94, Lag00, PPEZ97, PC92, RKCP97, Rom95, WBC^{+}94]. hautes [JLS^{+}95]. Heavy [BRTG90, Cug96, Rem96, Sur96, Bor92, Cin88, GT86]. Heavy-Ion [Cug96, Sur96, BRTG90, Bor92, Cin88]. Heidelberg [zP85]. Hele [Pat96]. helium [MN84, NLM^{+}94, Sch85, BL86, LL85]. heterogeneous [Pat96]. hétérogènes [Pat96]. Hétérostructures [CVR^{+}95]. Hg [MKG^{+}92]. hiérarchique [Sou94]. High [Dav06, JP99, MMdC^{+}07, NBCB^{+}07, AK88, BAGR88, DWD89, FFPM07, GFC88, JCL07, KEH07, MSS^{+}07, NBHC^{+}07, OBU^{+}88, REI^{+}88, VCB^{+}07, Lag00]. High- [NBCB^{+}07]. high-efficiency [NBHC^{+}07]. high-finesse [VCB^{+}07]. high-quality [FFPM07]. high-T [OBU^{+}88]. high-temperature [AK88]. highly [Bon84]. historical [Col85]. historique [Dan85]. history [vB89]. hole [FABC88]. holes [Mey91]. hollow [LDR07]. holograms [JWBB99, JP99]. homme [Cur85]. horizontal [Abi94]. hors [Bal96]. host [HJK^{+}07]. hot [Bor92]. hybrides [BCB^{+}95]. hydrazine [DLKP89]. hydrocarbures [DLKP89]. hydrocarbures/air [DLKP89]. hydrodynamics [Bed91]. hydrodynamiques [NA97]. Hydrogen [Cra85, SF84]. hydrogène [DLKP89]. hydroxyapatites [MPT^{+}07]. hyper [Gal99]. hyperfines [RP95a, RP95b]. hyphe [Pel04]. Hz [GMS^{+}94].

idées [BBG^{+}85]. II [BHWS91, CVR^{+}95, Cou96b, Jol93, MR91b, SV87b, VGM97]. III [FMOS97, Jus88, ML94, Qui95, Vol97c]. illustration [Gry85]. ILM [GGP84]. image [Bra99, PTC99]. Imagerie [AGG^{+}94, BG95, IS94, TPD^{+}97, Men94]. Imaging [CLM^{+}97, Lar06, STC^{+}07, LWMG91]. IMFM [FDSF92]. immediate [Maa97]. immobiles [PJ85]. imparfaite [Pat96]. imperfect [Pat96]. implantation [GGM^{+}97, MDF88]. implementation [LPPC99]. Implications [RC97]. importance [DGLR91]. imprimeur [Rab94]. Improvement [MJM^{+}07, Del00]. improvements [Del00]. impulsionnel [MBIA92]. impulsionnelle [KVC^{+}97, MHN^{+}97, RCK^{+}94]. impulsions [RMA94, Rib97, SFE95, SGFE97]. Impuretés [Bob05, Noz85, Dan85]. Impurities [Bob05]. In-line [MLP99]. incoherent [FF99, Pou97]. incohérentes [FDSF92]. index [JBTV07]. indice [SPJP92]. indirectes [RB91]. individuelles [Mey84]. Induced [Ave97a, Bar00, Blo85, MFM^{+}07a]. induit [BLW94]. induite [BLD^{+}95, JLS^{+}95, LDCBA92, MLCJ94, PBD^{+}95]. induites [Bar00, Gry85]. induits [BDB^{+}97, NA97]. industriel [BCKV97, CFHF94, Ste92]. industrielle [Dav92]. industrielles [EB94]. industry [Fri92]. Inelastic [ARC96]. Inertial [Deu86]. infinite [IFRS94]. Influence [PDAYS89, Rob85]. Infrared [MSS^{+}07, PFC91]. InGaAsP [Bou93]. ingénierie [CBOf95]. inhomogènes [Ber88]. initial [Mey91]. Injection [BBC^{+}95a, Bou93]. injection-locking [Bou93]. innovantes [Del00]. innovative [Del00]. inorganiques [BCB^{+}95]. inspiration [DG85]. instabilité [Pel04, Rab94]. instabilités [Abi94]. instabilities [Par84]. instability [Pel04]. instantanée [IS94]. instantanées [SL95]. instrumentation [Sch92]. Instruments [Lor98]. insulator [Let95]. insulators [LS96]. intense [Deu86, LB92, LSB94, LRRVT92, Sch84]. intenses [Sal92]. intensifiée [IS94]. Intensité [Pou92, Bla94, Bra99, Lag00]. intensity [Bra99, Lag00]. Intérêt [Com92, Rou92]. interacting [LMpk01]. Interaction [BA92, Bob86, CDMC92, FSBD95, Gil90, AB84, AAC97, Boc92, CCJ^{+}94, CV89, Fle09, GARG94, LS94, SBE92, Tro02]. Interactions [DZB89, Hua06, Mey03, Pie01, BPJ^{+}07, PDGM97, PBD^{+}95]. interconnexions [MFG^{+}97]. interdites [Cas98]. interface [dSLR^{+}07]. interfaces [FDS94, SHDT88]. interfaciales [Rab94]. interference [Bou02, Bou02]. interféromètres [PJ97]. Interferometric [Bri85]. Interféromètrie [Min95, SPJP92]. interférométrique [CBOf95]. interferometry [PFC91]. intérieur [VL89]. interlayer [AK88]. Intermediate [GT86, Ale87, Mot85]. intermediate-mass [Ale87]. intermultiplet [DZB92]. internal [MRC^{+}03]. International [Sal85]. interne [FHF92]. interpretation [AB84, SB84]. interstellar [Bed91, Ber84, GZH89, Kro89, Ld89]. intracavité [KDC^{+}95, Rom95]. intracavités [PVJ^{+}95]. intracellulaire [Ott97]. Intrication [Gaë09]. intriqués [Our07]. Introduction [ASD93, Bén85b, BDT85, Cha99, Dan85, Lal85, Rey90, Col85]. invariance [Gra92]. Inverse [GFCB99]. inverses [GFCB99]. inversion [Cas94]. inversions [JJC84]. Investigation [KEH07, XZS^{+}07]. iode [RP95b]. Ion [Cug96, Sur96, BRTG90, Bor92, BDT^{+}95, CHH^{+}07, Cin88, Deu86, GT86, MN84]. ionique [Gal85, GGM^{+}97]. ioniques [DVT^{+}97]. Ionisation [RMA94, MKM97]. ionization [MPW84, NLM^{+}94]. Ionizing [Ave97a]. Ions [Rem96, Bab86, Ber84, Bon84, Esa86, LS94, WIB^{+}85]. irradiation [GGP84, MPW84, TBLH^{+}94]. irradié [RMA94]. Irreversible [SVS98]. isolant [Let95]. isolants [LS96]. isolated [Reu95]. isolées [Reu95]. isotopes [Sch03]. Isotopic [Cav00a]. isotopique [Cav00a]. issues [MRIC^{+}95]. IV [Vol97d].

J [MNO^{+}94]. Jaynes [AHOE95]. jet [Lah06, RMA94]. Jets [Sau94]. jonction [Per05]. Jones [Det99]. Josephson [Hua06]. juin [Ano92]. junction [Per05].

Kastler [Sal85, Bén85a, CT85a, Cur85, DG85, Rob85, Roc85, Str85, ski85]. Kerr [ALE95, MFM^{+}07a]. keV [GABG97, LRRVT92]. key [LBF^{+}07]. kHz [KEH07]. Kinetic [LMpk01, Sur96]. Kinetics [MPT^{+}07, GBC^{+}07]. km [LBF^{+}07, LDC^{+}07]. Kondo [Noz85]. Kr [BRG^{+}97, SVDL^{+}94]. KrF [CDMC92, GGCC97, MBIA92]. krypton [HSD^{+}94, KZAK^{+}92]. KTiOPO [XZS^{+}07]. kW [GMS^{+}94].

l [Kan09]. laboratoire [Kli94, Mat93, Roc85]. Lagrangian [CCH^{+}87]. Lamb [Owe86]. lamellaire [BSM^{+}97]. lamellaires [ASD93, KLC^{+}97]. lames [PJ97]. lampe [PPEZ97]. lampes [PV92]. Landau [Byc94]. Large [Ld89, BWDR95, GPL95, NNH95, TLG95]. Laser [Cou91, GMS^{+}94, GB95, MBIA92, Phi85, SGZ97, AGG^{+}94, AAC97, AFF^{+}94, ASP94, BBC^{+}95a, BMV^{+}07, BA92, BCKV97, Ben92, BJZ^{+}07, Bla94, Bob86, Boc92, BL92, BLTH94, BDB^{+}97, Bou93, Bri98, BLD^{+}95, CCJ^{+}94, CAG^{+}97, CDMC92, CPDMR94, CB95, Cou94, DC92, Dam94, DAB^{+}94, DBD07, Dum85, EB97, FTGB95, GGA^{+}97, GGCC97, Gau92, GRA^{+}94, GGP84, GARG94, GGA^{+}94, GGM^{+}97, HFF97, HTBL94, HJK^{+}07, JJC84, KRL^{+}95, Kam91, KZAK^{+}92, KCK^{+}92, Kli94, KRVV95, LB92, LSB94, LDCBA92, LBH92, LBDM94, MMdC^{+}07, MB97, MKM97, MFM^{+}07b, MPW84, MFG^{+}97, MLCJ94, MLJC97, MNO^{+}94, NLM^{+}94, NNH95, NG94, NA97, OSA94, PDGM97, PCFC07, PNCA94, PLP^{+}97, RMA94, RB95, Rom95, SBE92, Sau92, SA97, Sch08, SFB94, SFE95, SGFE97, Ste92, TBLH^{+}94, Tro02, VHBL97, WCB92, ZAC^{+}97, dCP07]. Laser-cooling [Phi85]. laser-matière [BA92, SBE92, Tro02]. laser-plasma [BLTH94]. laser-surface [PDGM97]. laser-UV [TBLH^{+}94]. Lasers [LBB^{+}94, Ano90, ATE92, AEA^{+}94, BA92, BH92, BDL^{+}92, BE92, Bre93, Bri98, BRB^{+}95, Cas94, EB94, FDSF92, GF95, GPL95, KDC^{+}95, KSG^{+}90, LSD^{+}97, MMdC^{+}07, MVCF95, Miq97, NLNA94, PBL^{+}94, Rib97, Sal92, TLG95, dUF97]. lasers/amplificateurs [GPL95, TLG95]. Late [Laf91b, Mic91]. lateral [BBDC07]. latéraux [NNH95]. law [LS96]. laws [DGLR90]. leads [Saf97]. lecture [Bra99]. Lectures [And84, Cla84a, Cla84b]. LEDs [Dav06]. lensing [Not88]. lentes [DNB^{+}95]. lentilles [TPD^{+}97]. LEP [PT88]. leur [CMF^{+}95, Fré00, Maz85, Muz95]. leurs [MB97, SM97]. level [Byc94, DDL^{+}07]. libre [MMK94]. libres [Cou91]. liés [Ano84, LB02]. life [GZH89]. LIGA [BBH95]. light [Blo85, Dav06, FF99, Gil90, HHP89, MFM^{+}07a, Our07, XZS^{+}07, dLMDP07]. limit [Ste85]. limitations [LS96, Ren92]. limites [HG94]. limits [LS96, Bor92]. Line [Kat94, BSB91, Bom91, MLP99, SB84, SF84]. linéaire [ALE95, Cal98, CB95, Gry85, LEPMS00, PTC99]. linéaires [CBOf95, FOF95, GHZ^{+}95, Wyo92]. linear [Kro89, LEPMS00, TD85]. linéarité [ALE95]. linéarités [SKGRF95]. linearity [MLP99]. lines [GSH^{+}84, Jol93, LD91, MN84, MRC^{+}03, Pet99]. Liquid [VGG97, Cal99, Kat94]. liquide [CCJ^{+}94, Cal99, VGG97]. liquides [CMF^{+}95, Que84]. liquids [CG84, Dyu94]. lithium [Sch03]. lithographie [HG94, HST97]. LMJ [SA97]. LO-phonon [GBC^{+}07]. local [PTC99, RMR88]. locale [PTC99]. Localisation [DKL86, Fle09]. localization [Fle09, Lav88]. locked [NLM^{+}94]. locking [Bou93, Gaë09]. ‘logarithmic [DGLR91]. loi [FMSZ95, LS96]. loin [Eph84, Mey84]. Long [Koh85]. longitudinal [Bre97]. lors [AAC97, VPF95]. Louis [EB94]. Louis-Blériot [EB94]. lourds [Bab86, Cug86]. Low [GFC88, Bon84, BHS^{+}94, FFPM07, PBG^{+}84, SF84, Bar00]. Low-energy [Bar00]. lumière [CFHF94, Dei97, Our07, Que84, SRF95]. luminescence [SVDL^{+}94]. luminescents [RAR^{+}95]. lumineuse [Pou92]. lumineux [CT85b]. Lumino [Tos85]. Lumino-réfrigération [Tos85]. luminosity [AGTS91].

m [HST97, NVJ^{+}95]. machine [LRRVT92]. macroscopiques [Pet91]. magnéetique [Reu95]. magnésium [KSF^{+}94]. Magnetic [Gen85, Met85, BSB91, Bom91, MRC^{+}03, Ram85, Reu95, LWMG91]. magnetically [Lah06]. magnetically-guided [Lah06]. magnétique [Gen85, TVW^{+}95]. magnétiquement [Lah06]. magnétiques [Dan85, Noz85]. Magnetism [Vil03]. Magnétisme [Vil03]. magnetization [Gen85]. Magnéto [Bou00, BBL07, Bou00]. Magneto-electronics [Bou00]. Magnéto-électronique [Bou00]. magneto-photonic [BBL07]. maille [VL89]. MALIX [LRRVT92]. man [Cur85]. Manifestations [Cal99]. manquante [Maz89]. marée [Lum85]. Markovian [Cam99]. markovien [Cam99]. mars [Bré94]. mass [Ale87]. masse [CDMC92, HST97, Maz89]. masses [Mey03, Mey91]. master [BSB91]. Materials [Mot85, Cas98, Cla99, Pon07]. Matériaux [Cas98, ASD93, AGG^{+}94, ASP94, BBC^{+}95b, CBOf95, Cla99, EB97, FOF95, LSD^{+}97, NA97, OSA94, PNCA94, Wyo92]. matière [AVC92, BA92, Boc92, GARG94, SBE92, Tro02]. matrices [Det99]. matrix [BSB91]. matter [Tro02]. mean [Mey03]. mean-field [Mey03]. means [XZS^{+}07]. Measurement [JBTV07, DDL^{+}07, MLP99]. measurements [OBU^{+}88, PRG94, SB84]. measuring [Saf97]. mécanique [Bal96, Van01]. Mécanismes [SKGRF95, ML94]. mechanical [Ber84]. mechanics [Bal96, Diu90, DGLR90, DGLR91, Van01]. médecine [ATE92]. Media [Mon95]. Medium [GR94, Bed91, Ld89]. méesoscopiques [Reu95]. mégajoule [SA97]. mélange [Bri98, BLD^{+}95, PBD89]. mélanges [DLKP89, GMS^{+}97, SM97, VP85]. mélasses [Cou96a, Cou96b]. membrane [SM89]. mémoire [KRL^{+}95]. memoriam [ski85]. mercure [PPEZ97, ZD94]. mercure-gaz-rare [ZD94]. mercury [PCFC07]. mérite [LBS^{+}92]. mesonic [Mat91]. mesoscopic [Hua06, Pie01, Reu95, Tor02, lS08]. Mesoscopics [Lar06]. mésoscopiques [Pie01, Tor02]. Mesure [ALE95, MHN^{+}97, SPJP92, Saf97]. Mesures [PVJ^{+}95, SL95, GGPR95]. metal [Cal98, Let95, SHDT88, Let95]. métal-isolant [Let95]. metal-semiconductor [SHDT88]. metallic [BdR91]. métallique [LS94]. métalliques [ASP94, Cal98, EB97, Lem92, OSA94]. metals [Mot85]. Metamaterials [Kan09]. métamatériaux [Kan09]. metastable [BCP^{+}07]. métaux [Dan85]. method [EQ85]. Méthodes [Bor85, BL86, BBG^{+}85]. methods [Dhe90, MLT85]. métrologie [Lor98]. metrologist [Bou96]. métrologue [Bou96]. metrology [CHH^{+}07, Lor98]. Mg [HHP89]. MHD [Par84]. Micro [DC92, Dam94, MFG^{+}97, SFE95, SGFE97, BBH95, BBDC07, Kan09, Miq97, Pou97, SFB94]. micro-cavities [BBDC07]. micro-décharges [Pou97]. micro-électroniques [MFG^{+}97]. micro-ondes [Kan09]. micro-optique [BBH95]. Micro-usinage [DC92, Dam94, MFG^{+}97, SFE95, SGFE97, Miq97, SFB94]. Microcavité [TVW^{+}95, RAR^{+}95, Sen03]. microcavities [Gay01, NBCB^{+}07]. Microcavity [Sen03]. microchamps [AFH94]. microélectronique [Rou92]. Microfield [ST84]. microfocalisation [SLI^{+}94]. Micromanipulation [Ott97]. micromirrors [HAC^{+}07]. micronique [TPD^{+}97]. Microscopie [Joy92, Bra99, Gal85]. microscopiquex [Gia84]. microscopy [Bra99, SHDT88, STC^{+}07]. microseconde [PV92]. microstructuration [Fer02]. microstructuring [Fer02]. microtechniques [BH94]. microwave [JCL07]. microwaves [Kan09]. milieu [BG95, Mon95, Pec85, Per05]. milieux [Ber88, Bri98, SV87a, SV87b, SM89]. minces [AGF94, AAC97, AFF^{+}94, BCKV97, BBD95, CAA^{+}94, CDMC92, FHF92, NFV^{+}94, dUF97]. miniatures [LBS^{+}92]. minima [SCSN93]. Miroir [VLL^{+}95, CB95, LLC^{+}95]. miroirs [Dav92]. mise [Pet91]. Mixing [GCT^{+}97, Blo85, Bri98]. mixte [Abi94]. mixture [RMR88]. Mixtures [Sch03]. mJ [IS94]. mode [MMZ^{+}97, NLM^{+}94, NBHC^{+}07, PPEZ97]. mode-locked [NLM^{+}94]. Model [AHOE95, BHWS91, NF94, GSH^{+}84]. Modèle [Ren92]. modeled [Pel04]. modelées [Pel04]. modèles [ML94]. modeless [MMdC^{+}07, MFM^{+}07b]. Modélisation [BE92, NLNA94, PP92a, AB84, PBL^{+}94]. Modelling [IFRS94, KSG^{+}90]. Models [Rem96, Ale87, AA88, BB85]. Modern [KTB90, Kam91, CCH^{+}87]. modes [CB95, NNH95, NBCB^{+}07, Vol97b, VGM97, Vol97c, Vol97d]. modulation [BWDR95]. modulator [dLMDP07]. Molasses [Cou96a, Cou96b]. moléculaire [Lef90, Lem92, PP85, Per05]. moléculaires [SV87a, SV87b]. Molecular [dB84, Cin88, Cra85, Per05]. molecule [MJM^{+}07, RP95a]. Molecules [DNQR91, ABS84, Ber84, FDS94, Kro89, Ld89, TFdFR89, Vil03, BBG^{+}85, BLW94, CMF^{+}95, DNB^{+}95, Vil03]. moments [EQ85, Van01]. momentum [PBG^{+}84, Van01]. monde [BDLM09]. monochromateur [BSM^{+}97]. monomodes [PCS^{+}95]. Morphogénèse [Hen94, MMK94]. morphologie [AGF94]. Motion [Mon95, MN84]. Motions [Cho96]. Mott [Eli94, LS96, Lav88]. mou [ABB94]. mouillabilité [Lav89]. Mouillage [Pat96, QdR98]. mous [Joy92, PP92b, SPJP92]. mouvement [Mon95]. Mouvements [Gia84]. moyen [Mey03, PVJ^{+}95]. moyenne [LBB^{+}94, SBE92]. moyennes [Ste92]. multiaxes [BRB^{+}95]. multichargés [LS94]. multicouche [AB94, BSM^{+}97]. Multicouches [BH92, ABB94, Dav92, KLC^{+}97]. multidrug [BRJ^{+}07]. multifragmentation [Bor92]. multilayered [Dhe90]. multimode [KRL^{+}95]. multipassage [RB95]. multiphonons [Fra91]. Multiphoton [AHOE95]. Multiphotonic [STC^{+}07]. multiphotonique [SM97]. multiples [Cag85]. multiplex [MSS^{+}07]. multiplexage [Gio95]. Multipole [BTP85]. Multivariate [Dés95]. muon [Hug85]. muonium [Owe86].

N [GCT^{+}97, LDCBA92, LS94]. nanocristaux [SKGRF95]. nanocrystals [XZS^{+}07]. nanodiamonds [ZWJ^{+}07]. nanometer [BHS^{+}94]. nanométriques [BH92]. nanoparticles [FSC^{+}07]. nanoporeux [Fer02]. Nanoporous [Fer02]. nanoprobes [FSC^{+}07]. nanoseconde [Ayr95]. nanosphere [RMT^{+}07]. nanostructures [Nau02]. nanotube [KBW^{+}07]. naturel [Pec85]. Nd [MVCF95, BBC^{+}95a, CB95, GB95]. near [JWBB99]. nebulae [AGTS91, Ryt91]. nébuleuses [Pas86]. negative [Esa86]. néon [PC92, WBC^{+}94]. networks [Fri92]. Neural [Fri92]. neurochirurgie [AEA^{+}94]. neutral [MN84, Met85, Phi85]. Neutron [Ram85, Bar00, Mey91]. neutron-induced [Bar00]. neutronics [Reu00]. neutronique [Reu00]. neutrons [Bar00]. nickel [ZWJ^{+}07]. nickel-related [ZWJ^{+}07]. nitride [BLTH94, Tor04]. nitride-based [Tor04]. nitrure [AAC97, FHF92]. nitrures [Tor04]. niveau [Sau95]. nm [CDMC92, KZAK^{+}92, MLJC97, MNO^{+}94, OSA94, PV92]. No [Ano92, Ano94, Ano95a, Ano95b, Ano97a, Ano97b, Ano05, Ano06, Ano07a, Ano07b, Ano08a, Ano08b, Ano09a, Ano09b, LDCBA92]. Noble [MRC^{+}03, DZB89]. noicissement [VPF95]. noir [Lum85]. noise [RH85, Tor02]. Non [MLT85, ALE95, Bal96, BLW94, Cal98, Cam99, Can04, CBOf95, CB95, FOF95, GHZ^{+}95, GGPR95, Gry85, MLP99, MLJC97, Our07, PTC99, PRG94, SKGRF95, TD85, Vol97b, VGM97, Vol97c, Vol97d, Vol97e, VGG97, Wyo92, Vol97b, VGM97, Vol97c, Vol97d, Vol97e, VGG97]. non-adiabatique [BLW94]. Non-conventional [MLT85]. non-demolition [PRG94]. non-destructives [GGPR95]. non-equilibrium [Bal96]. non-extensive [Vol97b, VGM97, Vol97c, Vol97d, Vol97e, VGG97, Vol97b, VGM97, Vol97c, Vol97d, Vol97e, VGG97]. non-Gaussian [Our07]. non-gaussiens [Our07]. non-linéaire [Cal98]. non-linéaires [FOF95]. non-linéarité [ALE95]. non-linéarités [SKGRF95]. non-Markovian [Cam99]. non-markovien [Cam99]. non-perturbative [Can04]. nondegenerated [Dyu94]. Nonlinear [Fau94, Cal98, PTC99]. normal [AHO92]. Normale [CT85a]. normes [Mét00]. norms [Mét00]. Note [Ave97b]. Nouveaux [Wyo92, FDSF92]. nouvelle [LB92]. Nouvelles [FS92, RP95b, GFCB99]. novel [NBHC^{+}07]. noyau [Gal99, Mat93, MLJC97]. Noyaux [Eph84, Mey84, AA91, BM89, Gia84, RS90]. ns [IS94]. nuage [KRVV95]. nucléaire [AS01, Bar00, Des01, GG87, Qué92, Sur91]. nucléaires [Cav00b, Mey03, Reu00]. Nuclear [AR90, Ber87, CLM^{+}97, Dés95, LD91, Rem96, Bar00, Cin88, Des01, HR92, Kle85, Mey03, Reu00, Suz84, AS01, Cav00b]. Nucléation [LB06, LB06]. Nuclei [Cho96, AGTS91, Bor92, Gul04, Mat91, MV92, Ott85, Bal96]. nucleiques [MB97]. nucleon [Bre97, Bre97]. nucleonic [IFRS94]. nucleus [CV89, Gal99]. nuclides [EQ85]. numériques [BDB^{+}97]. nutations [Maa97].

O [LDCBA92, MRG^{+}88, Ver88, LS94, LCSY88, TFdFR89]. oblique [DWD89]. observables [BTP85]. Observation [AGF94, LLC^{+}95, VCB^{+}07]. observational [Not88]. observed [AGTS91]. observée [BLD^{+}95]. Obstruction [JWBB99]. obtenu [MMN^{+}94]. obtenue [VPF95]. octobre [Ano94]. offertes [PJ97]. OH [MHN^{+}97]. ombroscopie [CCJ^{+}94]. onde [HAC^{+}95b, Jac07]. ondelettes [Muz95]. ondes [Bor85, BLD^{+}95, DKL86, Kan09]. One [Cou96a, Cou96b, Mon95, vA08, FABC88]. One-Dimensional [Cou96a, Cou96b, Mon95, vA08]. onto [BRJ^{+}07]. OPO [Ayr95]. OPOs [KEH07]. Optical [CdBB^{+}99, Cou96a, Cou96b, DBD07, HLS85, Ott85, Sch84, WIB^{+}85, zP85, AAB^{+}85, Ami85, BCP^{+}07, Ber85, Bra99, CHH^{+}07, DDL^{+}07, DG85, FGD07, GSH^{+}84, JCL07, JP99, LDC^{+}07, Rei05, Ser85, Str85, STC^{+}07, Vin85, dC85, Lor98]. Optically [Sch85, AS85, BTP85]. optics [BJZ^{+}07, Dhe90, HAC^{+}07, Kan09, PRG94, dLMDP07]. optimisation [FOF95, PV92]. options [Del00]. Optique [AB94, Ano90, BBH95, BBC^{+}95b, BBG^{+}85, BG95, BH92, BCB^{+}95, Bou85, Bra99, Dav92, Dum85, FMSZ95, GHZ^{+}95, GPL95, GGPR95, Gry85, Hen94, HST97, Kan09, LL85, Lor98, Pec85, PF93, PJ85, Rob85, SM89, TLG95, Van85]. Optiques [IMCD97, SLI^{+}94, Bor85, Cou96a, Cou96b, GDA^{+}95, Hac95a, Ott97, PCS^{+}95, Sau92, SKGRF95, dM95]. opto [BBF^{+}95]. opto-électronique [BBF^{+}95]. Optoelectronic [JCL07, LPPC99]. optoélectronique [Qui95]. Optomechanical [VCB^{+}07]. Orbital [Reu95]. orbitale [Reu95]. orbite [HSD^{+}94]. Order [Cho96]. Ordering [Kan85, Kop88]. ordinary [Sou85]. ordinateurs [Bla03]. ordre [BLD^{+}95, GHZ^{+}95]. organic [BMV^{+}07]. organique [Rau89]. organiques [BBS94, BCB^{+}95, LBFdR94]. organiques-inorganiques [BCB^{+}95]. originating [GSH^{+}84]. origine [SM89]. origines [VPF95]. origins [GZH89]. OrthoTerPhenyl [VGG97, VGG97]. oscillateur [GHZ^{+}95, GF95].

Oscillateurs [Hac95a, dM95, Lef90]. Oscillations [Maa97]. Oscillators [Fau94, JCL07]. OTP [VGG97, VGG97]. outil [MNO^{+}94]. ouvert [Abi94]. ouverture [GPL95, TLG95]. overlapping [MN84]. oxidation [Ver88]. oxide [DY88, Rao88]. oxides [Jef88, OBU^{+}88]. Oxygen [FABC88]. oxygène [DLKP89].

packet [Bou02]. paires [ML94]. pairing [AK88, FABC88]. pairs [HJK^{+}07, RMR88]. panache [AGG^{+}94, AAC97, CDMC92]. paquets [Bou02]. parabolic [FGD07]. parallel [CdBB^{+}99]. parameters [BBL07]. paramétrique [GHZ^{+}95, GF95, KVC^{+}97]. paramétriques [Hac95a, PCS^{+}95, dM95]. parité [Bou85]. part [Cou96a, Cou96b]. Particle [Rem96, Deh85]. particles [Ale87, BJ88, BHS^{+}94]. particules [Bob86, SB88]. particulières [LM00]. partie [Cou96a, Cou96b]. partir [BSMS92, Muz95]. passante [BSM^{+}97]. passive [Lar06]. past [Hap85]. Pattern [NF94]. Paul [KRVV95]. Pb [MPT^{+}07]. peg [Pel04]. pendular [TD85]. penetration [DY88]. performances [CB95, Joy92]. performant [MNO^{+}94]. peripheral [BRTG90]. Perot [PVJ^{+}95, TD85]. Personal [CT85a]. perspectives [Dei97, Ler00, Pou97, SGZ97, PJ97]. Perte [HAC^{+}95b, KRL^{+}95]. Pertinence [LS96]. perturbatif [Can04]. perturbations [Koh85, PBD89]. perturbative [Can04]. peu [SB88]. Phase [AHOE95, Gul04, AHO92, ALE95, Bri98, BLD^{+}95, CMF^{+}95, FF99, Ger04, MLP99, PTC99, VGG97, dM95, Bri98, Ger04]. phase-shift [MLP99]. phases [LCSY88]. phenomena [DSS87]. phénomène [Abi94]. phénomènes [MR91b, NA97]. Phenomenological [Cla99, LEPMS00]. phénoménologique [Cla99, LEPMS00]. phenomenon [MV92]. phonon [GBC^{+}07]. phosphorites [Jul89]. photo [BBDC07, PBL^{+}94, VPF95]. photo-déclenchés [PBL^{+}94]. photo-detection [BBDC07]. photo-noicissement [VPF95]. photoablation [AGG^{+}94, CAA^{+}94]. photochimie [MB97]. photochimiques [NFV^{+}94]. photoemission [MRG^{+}88]. photolithographie [BH94]. photoluminescence [ZWJ^{+}07]. photoluminescent [FSC^{+}07]. photométrie [Pou92]. Photon [Bon84, Jac07, NLM^{+}94, NBHC^{+}07, Per05, RH85, ZWJ^{+}07, dCP07]. photonic [BBL07, Dav06, Fer02, Lab00, LDR07, Tor04, Cas98]. photoniques [Cas98, Fer02, FS92, Lab00, Tor04]. photons [CT85b, Cub88, DNB^{+}95, FMOS97, GNMS^{+}95, GJJR95, Hul95, Jac07, JLS^{+}95, Per05]. photopolymères [BBC^{+}95b]. photostability [MJM^{+}07]. physical [AB84, DGLR91]. physicians [LWMG91]. physicien [Cav00b]. physicist [Cav00b]. physicists [LWMG91]. physico [LB02]. physico-chimiques [LB02]. physicochemical [LB02]. Physics [Cla99, CT85a, Dés95, Sur96, AS01, Kam91, LB02, LDC^{+}07, SB84, Suz84, lS08, Fré00, Sal85]. Physique [Fré00, Ano84, AS01, Chi92, Cla99, Enc88, Lem92, LB02, MDF88, PP85, Qué92, Roc85, Tou85, Ano90]. physiques [VL89, VPF95, VP85]. picosecond [FFPM07]. picoseconde [CMF^{+}95, GRA^{+}94, GB95, Hac95a, LPvWR95]. piège [Ger04, KRVV95]. piézoélectrique [CVR^{+}95]. piles [Cla99]. pillar [NBCB^{+}07]. pinces [Ott97]. pinch [AFH94, CFHF94]. Pinning [Ovc94]. placés [RAR^{+}95]. plain [Sou85]. planaire [RAR^{+}95]. Planar [Mon95]. plane [Cla84a, Cla84b]. planes [FABC88]. planétaires [Pas86, Rau89]. planetary [AGTS91, Ryt91, dB84]. planètes [Gen87]. plasma [Ala86, AFH94, AAC97, Bob86, BLTH94, CAG^{+}97, FGVV89, GRA^{+}94, GABG97, LDCBA92, LBH92, MMN^{+}94, VHBL97]. plasmas [Ben92, CGST92, Gau92, GGP84, HFF97, NG94, PP92a, PC92, Rei05, Sau92]. PLGS [MFM^{+}07b]. plus [BM89]. plusieurs [SCSN93]. Poincaré [Det99]. point [Cav00b, Pie00]. Poland [DG85]. polarisation [Det99, JDS^{+}95, LPvWR95]. polarisé [BL86, LL85]. polariton [CGM87]. Polaritons [Sen03]. polarization [BSB91, Bom91, Det99, Kle85, SB84]. polarized [Fre85]. polarons [ML94]. polychromateur [BSM^{+}97, KLC^{+}97]. polychromatic [dCP07]. polycristalin [BBF^{+}95]. polymères [BL92, LBDM94, MDF88]. Pompage [Bou85, BBG^{+}85, Dum85, FS92, Gry85, LL85, Pec85, PJ85, Van85]. pompé [CB95, GB95, NVJ^{+}95, RB95]. pompés [MVCF95]. Population [JJC84, Cas94]. Porte [Bré94]. portée [SB88]. position [AGTS91]. possible [Qué92]. post [LDCBA92]. post-décharge [LDCBA92]. postulate [Diu90]. potential [AB84, BB85, Clé08]. potentials [AB84]. potentiel [BCKV97, Clé08]. potentielles [AEA^{+}94, BA92]. potentiels [BRG^{+}97, SCSN93]. pour [AB94, Bla03, BH92, CFHF94, Dav92, DMFR94, FMOS97, Fou92, FHF92, LRRVT92, Lor98, Men94, MFG^{+}97, MNO^{+}94, NFV^{+}94, Qui95, Rou92, SBE92, SLI^{+}94, Ste92, SGZ97, Tro02, VHBL97, Wyo92, dUF97]. power [MMdC^{+}07]. Préambule [Laf91a]. prébiotique [Rau89]. predictions [Ber84]. Préface [Bor85, Bré94]. premier [BLD^{+}95]. première [LB02]. premières [ZAC^{+}97]. Premiers [VGM97, Vol97d]. préparation [MKM97]. préparés [GGM^{+}97]. presence [BSB91, Bom91, Sch84, HTBL94]. present [Hap85, Nen92]. Présentation [Bro85, Vol97a, Bro85]. pression [CRKP94, CAG^{+}97, KZAK^{+}92, KVC^{+}97, LBH92, PPEZ97, PC92, RKCP97, ZD94]. pressure [BHS^{+}94]. Principes [Bla94, LM00, Mét00, Dep87, Gio95, IMCD97]. principles [Dhe90, LM00, Mét00]. probe [Sch85]. problem [Byc94, Jol93]. Problèmes [Ano84, GFCB99]. problems [GFCB99]. procédé [BBH95, SFE95]. Procédés [Pet94, SFB94]. process [BJ88, Can04]. processes [BRTG90, Buc85, DC87, LB02, PBG^{+}84, Per05, SVS98, DZB89]. processing [PTC99, STC^{+}07]. processors [CdBB^{+}99]. Processus [Can04, PDGM97, BRG^{+}97, LB02, Per05]. proche [Dei97]. produced [GGP84]. Production [Eph84, GARG94, GABG97, HST97, Rou92, SBE92]. produit [ABB94, CAG^{+}97, VHBL97]. produits [PP92a, TBLH^{+}94]. Professor [Sal85]. profil [SVDL^{+}94]. profile [KBW^{+}07]. profiles [MN84, ST84, SF84]. profils [CGST92]. profonde [BH94]. profonds [MLCJ94]. programme [Ste92]. Progrès [FDSF92, GABG97, KCK^{+}92, Pou97]. project [Nen97, Lor98]. projection [FGVV89]. projet [Lor98, Nen97, SA97]. Propagation [Méj05, Cla99, Pon07, VHBL97, Méj05]. Properties [AHOE95, Mon95, ZB90, Clé08, Ger04, GFC88, IFRS94, Rei05]. propos [Ano84, Ran84, SK85, WHPD94]. propres [BRB^{+}95]. Propriétés [Ala86, Clé08, MR91a, MR91b, ML94, RAR^{+}95, Saf97, Ger04, Mon95]. Prospects [Ler00, BR88]. protection [Mét00]. protéines [MB97]. Protonated [Ber84]. Pseudo [Rem96]. Pseudo-Particle [Rem96]. puisé [RMA94]. puisées [LBS^{+}92]. puissance [FDSF92, LBB^{+}94, SBE92, SGZ97]. puissances [Ste92]. Puits [Qui95]. pulse [FFPM07, KDC^{+}95]. pulsées [FDSF92]. pulses [Méj05, VKCP97]. pumped [AS85, BTP85, Sch85]. pumping [AAB^{+}85, Ami85, Ber85, DG85, HLS85, Ott85, Ser85, WIB^{+}85, zP85]. pur [MKG^{+}92]. purity [JCL07].

quadrupole [EQ85]. qualité [FMOS97]. quality [FFPM07]. quantification [Maa97]. quantique [Ano90, Dar84, GF95, Nau02, Rey90, Saf97, Tor02, Van01]. quantiques [Bla03, GGPR95, Maa97, Our07, Qui95]. Quantitative [CLM^{+}97, Rom95]. quantization [Maa97]. Quantum [DW85, HAC^{+}07, LBF^{+}07, BdR91, Ber84, Bla03, Dyu94, GBC^{+}07, Gog94, Our07, PRG94, Saf97, Ste85, Tor02, VCB^{+}07, Van01, Nau02]. quantum-mechanical [Ber84]. quark [BB85]. quark-antiquark [BB85]. quarks [Mat93]. quasar [Pet99]. quasars [Jol93, Pet99]. Quasi [NF94, FABC88, RMR88, dM95]. quasi-accord [dM95]. quasi-one [FABC88]. Quasi-Two-Dimensional [NF94]. quatres [BLD^{+}95]. que [Fra91]. quelles [AVC92, Ren92]. Quelques [PP85, Tro02, BH92, LBS^{+}92, Mon95, RP95a].

Rabi [Maa97, TVW^{+}95]. radial [MRC^{+}03]. Radiation [Ave97a, Buc85, NLM^{+}94, dC85]. radiative [Bom91, DC87, PT88]. radiative-collision [DC87]. radiatives [RAR^{+}95]. radicalaires [RC97]. radii [Mey03]. radio [Gen87]. radioactifs [Com92]. Radioactivities [HHP89]. radiofrequency [BHS^{+}94, MLT85]. radiological [Mét00]. radiologique [Mét00]. radius [Deh85]. raies [CGST92, GMS^{+}97, Pet99]. rainforest [Cha00]. Raman [Cou96a, Cou96b, KBW^{+}07, LDR07, LR87, MFM^{+}07a]. Random [Mon95, Clé08]. range [JJC84]. Raphe [MLJC97]. rapide [CRKP94, KVC^{+}97, MKG^{+}92]. rare [AB84, DVT^{+}97, LBH92, MNO^{+}94, ZD94]. rares [CRKP94, GMS^{+}97, RKCP97, SM97]. rat [MLCJ94, MLJC97]. rate [FFPM07]. rates [KEH07]. ray [GGP84, JJC84, JP99, KSG^{+}90, Las90, LD91, MRG^{+}88, vB89]. Rayleigh [BWDR95, Cou96a, Cou96b]. Rayonnement [ABB94, LSB94, AB94, BH94, BA92, BBS94, BL92, DMFR94, GABG97, LB92, Lem92, Nen92, NA97, Pet94, PPEZ97, Pou97, RKCP97]. rayonnements [PNCA94, SCR^{+}94, WCB92]. rayons [Fou92, GARG94, HG94, Joy92, Mey03, SBE92, SLI^{+}94, VKCP97]. Rb [AS85]. Réacteur [FHF92]. Réacteurs [Cav00b, Reu00]. reacting [Cla84a, Cla84b]. reaction [Ver88, Can04, Can04]. réaction-diffusion [Can04, Can04]. Reactions [Rem96, AR90, ABS84, Bar00, Cin88, Cug00, Dor84, GT86, Bar00, RB91, RC97]. réactive [HTBL94]. réactivité [Lef90, Reu00]. reactivity [Reu00]. reactors [Cav00b, Reu00]. reading [Bra99]. Réalisation [Dav92, AAC97, Bla94, CDMC92, FMOS97, GGA^{+}94, KLC^{+}97]. realization [JWG^{+}07]. récents [ACG^{+}97, Des01, FDSF92, GABG97, Miq97, Pou97]. reception [dC85]. Recherches [Bla85, EB94, Hei89, KCK^{+}92]. recollections [CT85a]. Recombinaison [MVCF95]. recombination [KSG^{+}90]. reconstruction [JWBB99, JP99]. recristallisation [GGM^{+}97, dUF97]. Recuit [BDB^{+}97, SGZ97]. redistribution [Sch84]. reduced [RH85]. Réduction [GF95, Rey90]. reflection [dSLR^{+}07]. réflectivité [CGM87]. réflexion [FSBD95, GNMS^{+}95, HAC^{+}95b, VLL^{+}95]. Reformation [EB97]. refraction [JBTV07, SPJP92, VPF95]. réfrigération [Tos85]. refroidi [DMFR94]. Refroidissement [KRVV95, Lah06]. régénératif [MRIC^{+}95]. régime [Ayr95, Cam99, GHZ^{+}95, LPvWR95]. régimes [Hac95a, KDC^{+}95]. règle [TVW^{+}95]. régularité [DBG89]. related [LB02, ZWJ^{+}07]. Relation [BWDR95, SVDL^{+}94]. relativistes [ABB94]. Relativistic [Cug96, BB85]. relativité [Gou93]. Relaxation [Cam99, LFTV95, Bén85b, BGC^{+}07, CG84, GGM^{+}97, KZAK^{+}92, PDGM97, Sen03]. relevance [Mat91, LS96]. Renormalisation [Sou94, Can04]. renormalization [Can04]. répartition [PBD89]. repetition [FFPM07, KEH07, RCK^{+}94]. réponse [LEPMS00]. réponses [Ren92]. research [Ber85, BDT85]. réseau [BSM^{+}97, JDS^{+}95]. réseaux [DKL86, KLC^{+}97]. resistant [BRJ^{+}07]. Resistive [Par84]. résolue [Dum85]. resolution [JP99, HG94, WBC^{+}94]. Resonance [LWMG91, BBD95, Gen85, AS85, Gen85, Ram85, Ser85, Str85, zP85, GMS^{+}97]. Resonances [Cin88, AA91, Blo85, Fra91, Gry85, Lef90, SKGRF95]. résonnante [FSBD95, Har85]. résonnantes [Cag85]. response [LEPMS00]. rest [Deh85]. résultats [ACG^{+}97, WCB92]. Results [ARC96]. retardé [Jac07]. retinas [LPPC99]. rétroinjection [NNH95]. ‘Réunion [Bré94]. revealed [BRJ^{+}07, Pet99]. révélé [Pet99]. review [Owe86, PBG^{+}84]. revue [FGVV89]. rigidity [Wya05]. ring [Bre93, Rom95, Sch08]. rings [Reu95]. rise [LB06]. role [Kro89]. rotation [CMF^{+}95]. rotational [Vol97b, VGM97]. rotationnelle [MHN^{+}97]. rotationnels [Vol97b, VGM97]. Rotations [Van01, PF93]. rouge [GJJR95]. rouge/bleu [GJJR95]. rules [Suz84]. rupture [Lum85]. Russe [Bré94]. Rydberg [Gaë09, GSH^{+}84, Har85].

S [SKGRF95]. S.T.M. [SHDT88]. sable [Cla99]. same [KBW^{+}07]. sample [BLTH94]. sand [Cla99]. sans [Cas94, DNB^{+}95, PJ97]. saphir [Bla94, MRIC^{+}95]. saptiale [TLG95]. saturable [DBD07]. saturation [NVJ^{+}95]. saturée [Dum85, KCK^{+}92]. saut [ML94]. savant [Cur85]. scale [BHS^{+}94]. scales [Bor92]. scaling [GFC88]. Scanning [SHDT88]. Scattering [ARC96, Buc85, DZB89, LDR07, LR87, MFM^{+}07a]. schémas [MR91a, MR91b]. sciences [Rob85]. scientifiques [Ave97b]. scientist [Cur85]. scintillantes [IS94]. Se [SKGRF95]. sea [Sch03]. Search [BdR91, Ale87]. second [Tor04, XZS^{+}07]. second-harmonic [XZS^{+}07]. seconde [GHZ^{+}95]. seconde-ordre [GHZ^{+}95]. sein [Per05]. seismic [Lar06]. Selected [Cho96, Gog94]. Sélection [DNB^{+}95, PJ85]. Selective [dSLR^{+}07, Ami85, FSBD95, GNMS^{+}95, KZAK^{+}92]. sélectives [JDS^{+}95]. Self [BMV^{+}07]. Self-doubled [BMV^{+}07]. Semi [BB85, LFTV95, MR91a, MR91b, NNH95, VPF95]. semi-conducteur [NNH95, VPF95]. semi-conducteurs [LFTV95, MR91a, MR91b]. Semi-relativistic [BB85]. semiconducteurs [Bou00, Cam99, CGM87, CVR^{+}95, Dei97, GPL95, TLG95, TVW^{+}95]. Semiconductor [Voi97, Bou00, Bou93, Gay01, SHDT88]. Semiconductors [Rig95, Cam99, HLS85]. Semimagnetic [Rig95]. Sensibilité [DLKP89, Rom95]. sensitivity [MSS^{+}07]. separation [Cav00a, Cav00a]. séparatrices [PJ97]. ses [HG94, ZAC^{+}97]. set [LBF^{+}07]. set-up [LBF^{+}07]. setups [KEH07]. Several [Cil94, TFdFR89]. SHADOW [LBFdR94]. shape [Pel04]. Shaw [Pat96]. shift [MLP99, Owe86]. shock [Bed91]. Si [AFF^{+}94, GGM^{+}97, HFF97]. siècle [BM89]. siège [MMK94]. SiGe [HFF97]. signaux [Hei89]. SiH [BHS^{+}94]. silicium [BDB^{+}97, Fer02, FHF92, PLP^{+}97, SGZ97, dUF97]. silicon [Fer02]. similaritons [FGD07]. simple [GSH^{+}84, VGM97, Vol97d, Vol97b, Vol97c]. simples [ASD93]. simplifiée [Vol97b, VGM97, Vol97c, Vol97d]. Simulation [dUF97, LBFdR94, ST84]. simulations [BDB^{+}97]. Single [Deh85, Jac07, CHH^{+}07, KBW^{+}07, MJM^{+}07, NBHC^{+}07, ZWJ^{+}07]. single-mode [NBHC^{+}07]. single-molecule [MJM^{+}07]. single-photon [ZWJ^{+}07]. size [BdR91]. Slow [MFM^{+}07a]. small [Bre93]. smectique [CMF^{+}95]. SN [LD91]. SNe [BHWS91]. Sodium [MFM^{+}07b, Bla85, MMdC^{+}07]. soft [JJC84]. SOLEIL [Nen97]. Solid [Sch08, Bri98, Cal99, Cra85, SVS98]. Solid-state [Sch08]. solide [Cal99, KRL^{+}95]. solide-liquide [Cal99]. solides [Bri98, ML94, NA97, Rib97]. solidification [AGF94]. solids [CCH^{+}87, Wya05]. soliton [MFM^{+}07a]. Solubilité [GGM^{+}97]. solutions [Del00, KTB90]. Some [Mon95, LR87, Tro02]. somme [PTZR95]. sommeil [MLCJ94]. son [PPEZ97]. soumis [Abi94, PNCA94]. Source [GRA^{+}94, LBH92, NFV^{+}94, RCK^{+}94, SBE92, Vit92, CPDMR94, IS94, KVC^{+}97, LB92, NBHC^{+}07, PCFC07, ZWJ^{+}07]. Sources [Jac07, Kli94, AVC92, Dei97, FFPM07, FS92, FDSF92, Fou92, KSF^{+}94, LBS^{+}92, NG94, Rou92, WCB92]. sous [LBH92, LM00, MLJC97, SM97, TVW^{+}95]. sous-critiques [LM00].

Souvenirs [Cou85]. space [Deh85, DW85, Sou85]. Spallation [Cug00, Fré00]. spatial [MMZ^{+}97, MFM^{+}07a, dLMDP07]. spatio [Cha00]. spatio-temporelle [Cha00]. Spatiotemporal [Cil94, Cha00]. specific [LM00]. spécifique [BBD95]. spécifiques [AVC92, BDL^{+}92]. Spectral [Bou93, FGD07, SB84]. spectrale [BWDR95, GPL95, HFF97, KRL^{+}95, TLG95]. spectre [Pet99]. spectres [FSBD95]. Spectrométrie [PJ97, WBC^{+}94, CDMC92]. spectroscopic [Men94]. Spectroscopie [CBOf95, Cou96a, Cou96b, GDA^{+}95, GNMS^{+}95, MB97, MLJC97, PTZR95, Rom95, SM97, SRF95, Bén85b, DNB^{+}95, MHN^{+}97, MNO^{+}94]. spectroscopique [CAG^{+}97, MMN^{+}94]. Spectroscopy [Cou96a, Cou96b, Bon84, BPJ^{+}07, BRJ^{+}07, FDS94, GGP84, MSS^{+}07, MLT85, Rao88, RMT^{+}07, dB84, dSLR^{+}07]. spectrum [Jol93, Pet99]. spéculaire [VLL^{+}95]. sphere [Det99, Det99]. sphérique [Gou93]. spheroïdal [Kro89]. SPHINX [VKCP97]. Spin [Fre85, Hap85, Rig95, Sou85, Bre97, BGC^{+}07, Col85, HSD^{+}94, Sch85, Tou85, Vil03]. spin-orbite [HSD^{+}94]. spin-polarized [Fre85]. splitting [TVW^{+}95]. spontanée [Sen03]. Spontaneous [BBDC07, Gay01, JJC84, Sen03]. square [AB84]. square-well [AB84]. squeezing [AHO92]. Sr [LCSY88]. stabilisation [BBC^{+}95a]. Stabilité [LPvWR95, Eph84, Mey84]. stabilization [Bou93, Kle85]. stade [ZAC^{+}97]. Stages [Fri91, BHWS91, Laf91b, Mic91]. Standard [Ren92]. Standardissimo [Ren92]. standing [ZB90]. star [Laf91b]. Stark [MN84, NLNA94, ST84, Voi97]. stars [DNQR91, Fri91, MMdC^{+}07, Mey91, Ryt91, Whi91, dCP07]. state [Sch08]. states [GSH^{+}84, Our07]. Static [Clé08, IFRS94]. statiques [Clé08]. statistical [Bal96, Diu90, DGLR90, DGLR91]. statistique [Bal96]. stellaire [Laf91a]. stellaires [Ano84]. stellar [Mey91, Mic91]. step [JJC84]. stereo [PTC99, PTC99]. stigmatic [JP99]. Stimulated [LDR07, MFM^{+}07a, Sen03, Cou96a, Cou96b]. stimulée [BWDR95, Bor95, Cou96a, Cou96b, Sen03]. STM [lS08]. stochastic [LPPC99]. stockage [Cou91, Sch00]. stoichiometry [Rao88]. storage [Sch00, STC^{+}07]. stored [WIB^{+}85]. stress [Cla99]. strong [MPW84]. strongly [LMpk01]. Structural [TFdFR89]. Structuration [BL92]. Structure [Rao88, Ber87, Bre97, DBG89, GG87, MRG^{+}88, PDAYS89, Pet91, Van89, vB89]. structures [Fer02, Maz85]. Studied [Del00]. Studies [Ser85, BSB91, Bom91, DG85]. Study [PCFC07, Bre93, Fri92, KTB90, Lab00, Let95, Our07, TFdFR89, IFRS94, Per05]. sub [BBD95, GRA^{+}94, LM00]. sub-critical [LM00]. sub-Doppler [BBD95]. sub-picoseconde [GRA^{+}94]. subdiffusion [MRC^{+}03]. submicronique [BL92]. submicroseconde [PV92]. substrate [BPJ^{+}07]. Sum [FDS94, Suz84]. Sun [Nen97]. Super [Gal99, BBS94]. Superconducting [LCSY88, Bla03, BR88, OBU^{+}88]. Superconductivity [AA88, BAGR88, Lav88, Min94, RMR88, Eli94, Kop88]. superconductor [DY88, GFC88]. superconductors [AK88, BAGR88, Bob05, Gil90, LEPMS00, Ovc94, Rao88, REI^{+}88]. supercontinua [MSS^{+}07]. superdeformation [MV92]. Superlattice [Jus88]. Superlattices [Voi97]. supernova [Bed91, LD91]. supernovae [Jul89]. superpositions [Our07]. superréseaux [Qui95]. supraconducteurs [Bla03, Bob05, LEPMS00]. supraconductrice [CDMC92]. Surface [LR87, BL92, BLTH94, FSBD95, LS94, NA97, PBD^{+}95, PDGM97, PNCA94, SCR^{+}94]. Surface-enhanced [LR87]. surfaces [Cra85, EB97, Kan85, Koh85, Pat96, Pon07, SHDT88, SVS98]. sursauts [Ham85]. survol [Min95]. Susceptibilitée [Reu95]. susceptibility [Reu95]. suspensions [Pet91]. switchcubes [CdBB^{+}99]. symbiotic [Fri91]. Symétrie [GG87, Dep87, Gou93]. synchronisés [MVCF95]. Synchronization [NLM^{+}94]. synchrotron [BH94, BBS94, NLM^{+}94, Nen92, WCB92]. synthèse [TBLH^{+}94]. synthesis [BLTH94]. synthetic [JWBB99]. system [BMV^{+}07, IFRS94, LCSY88]. Systèmes [Mél97, ASD93, Bob05, LM00, SB88]. Systems [NF94, Bob05, BJ88, Byc94, LMpk01, LM00, Mél97, REI^{+}88, dC85].

t [AVC92, BAGR88, GFC88, OBU^{+}88]. T.H.T. [MHN^{+}97]. Taille [DBG89, SKGRF95]. tardives [Laf91a]. targets [Fré00]. tas [Cla99]. taux [RCK^{+}94]. technique [AFF^{+}94, DWD89]. Techniques [HSD^{+}94, Joy92, KTB90]. technologie [FOF95]. technologique [DM94]. télécommunications [FOF95]. TEM [ZB90]. temperature [AGTS91, AK88, REI^{+}88, MHN^{+}97, SVDL^{+}94]. température-profil [SVDL^{+}94]. températures [JLS^{+}95, RC97, SL95]. temporelle [BSMS92, Cha00]. temps [Dum85]. tendances [GFCB99]. tens [TFdFR89]. terahertz [CHH^{+}07, GDA^{+}95]. terawatt [Méj05]. térawatts [Méj05]. test [Jac07, Maa97]. tests [Dhe90, LDC^{+}07, VGM97, Vol97d]. TeV [Chi92]. TF [PJ97]. their [Fré00, Suz84]. Theoretical [Bre93, Sal85, Des01, JWG^{+}07, Cug96, Our07]. Théorie [Det99, Vol97b, VGM97, Vol97c, Vol97d, Vol97e, VGG97, ASD93, Kan09, Vol97c]. theories [Mey03, Mey03]. théorique [Our07, SV87a]. théoriques [Cug86, Des01, Mat93, MDF88, Ren92]. Theory [Dyu94, BJ88, Gog94, Kan09, Owe86, Vol97b, VGM97, Vol97c, Vol97d, Vol97e, VGG97, Det99]. thermal [DZB92]. thermalization [Bor92]. thermique [dUF97]. thermiques [Abi94]. thermodynamic [Ger04]. Thermodynamics [Ver88, DGLR90]. thermodynamiques [Ger04]. thirty [Str85]. thought [JWG^{+}07]. thousands [TFdFR89]. THT [KVC^{+}97]. thulium [GJJR95]. Ti [BLTH94, MRIC^{+}95]. Time [Bor92, Buc85, MJM^{+}07, TD85]. Time-scales [Bor92]. tissus [Gal85, MLCJ94]. titane [Bla94, HTBL94]. titanium [BLTH94]. tomographie [GFCB99]. tomography [GFCB99]. tool [BJZ^{+}07, CCH^{+}87, STC^{+}07]. topics [Gog94]. toroidal [MRC^{+}03]. trace [MJM^{+}07]. traceurs [Com92]. Traitement [LBDM94, OSA94, PTC99, PNCA94]. Traitements [ASP94, FMOS97]. transfer [ABS84, BAGR88, Bom91, DZB92, PBG^{+}84]. transfert [SV87a, SV87b]. transformation [AVC92, DGLR91]. transformée [Muz95]. transistors [dUF97]. Transition [Cil94, BLW94, BLD^{+}95, Cal99, Eli94, Let95, Mot85, Sou85]. transitions [Cag85, DZB92, RP95a]. translational [Vol97c, Vol97d]. translationnels [Vol97c, Vol97d]. transmission [DBG89]. Transmutation [Ler00]. Transport [LL85, Mon95, Nau02, Tor02, Bal96, MRC^{+}03, MR91b, ML94, Saf97, Tor02]. transversal [Gen85]. transversale [Gen85]. transverse [PCS^{+}95]. trap [BCP^{+}07, Ger04]. Trapping [BCP^{+}07, Phi85]. traps [Met85]. traversant [ABB94]. trends [GFCB99]. très [HG94, KRL^{+}95, Rom95, RC97, WBC^{+}94]. triatomiques [DVT^{+}97]. Triggered [ZWJ^{+}07]. trois [Ala86, Tou85]. tropicale [Cha00]. trou [Lum85]. trous [LFTV95]. tube [Vit92]. tunnel [Per05]. Tunneling [REI^{+}88, SCSN93, SHDT88]. turbide [BG95]. TW [Sau95]. Two [NF94, Fle09, Gaë09, Lab00, Let95, NLM^{+}94]. two-dimensional [Fle09, Lab00]. two-photon [NLM^{+}94]. type [RB91].

U.S. [Str85]. U.V [CDMC92]. U.V. [AGG^{+}94, BDL^{+}92, EB94, GGA^{+}94]. ultimes [HG94]. ultra [Sal92, VKCP97]. ultra-compact [VKCP97]. ultra-intenses [Sal92]. ultracold [Sch03]. ultracourtes [GDA^{+}95, GARG94]. ultrarapide [LFTV95]. Ultrarelativistic [Cug96]. ultrasound [Lar06]. ultrathin [BdR91]. ultraviolet [AAC97, KZAK^{+}92]. UM34 [FDSF92]. unexplained [LR87]. unidimensionnelle [Mon95]. unidimensionnelles [Cou96a, Cou96b]. unidimensionnels [DKL86]. unique [Mat93]. uniques [Jac07]. univers [Pet99, Maz85, Maz89]. Universe [Pet99]. unsolved [Jol93]. unstable [Ott85]. Updatable [FF99]. Upper [OBU^{+}88]. UPt [Min94]. uranium [Sch00]. usage [Ste92]. usages [Qui95]. users [PBG^{+}84]. usinage [DC92, Dam94, Miq97, MFG^{+}97, SFB94, SFE95, SGFE97]. using [Gaë09, LBF^{+}07, MJM^{+}07]. utilisant [Gaë09, IS94]. Utilisation [BSM^{+}97, CFHF94, LSD^{+}97, Men94, WCB92, ATE92, GABG97, Pet94, Lem92]. UV [AB94, NA97, Ave97a, BH94, Ben92, CCJ^{+}94, CGST92, CFHF94, Cou94, Dei97, Dhe90, FDSF92, KSF^{+}94, KLC^{+}97, MFM^{+}07b, MHN^{+}97, MLJC97, Pet94, PPEZ97, PC92, PNCA94, Rib97, Sch92, SCR^{+}94, SGFE97, TBLH^{+}94, Tro94, ZAC^{+}97, ZD94]. UVX [LB92, LSB94].

V [Vol97e, FMOS97, Jus88, Qui95, Vol97e]. valence [Cal98, Mot85]. valency [MRG^{+}88]. validation [Fré00]. vallée [Eph84, Mey84]. value [Deh85]. vapeur [EB97]. vapeurs [Lem92]. vaporization [PCFC07]. vapour [AS85, BTP85, dSLR^{+}07]. variables [LBF^{+}07, VL89]. variation [ALE95, VPF95]. veille [MLCJ94]. veille-sommeil [MLCJ94]. velocities [DWD89]. Velocity [Ami85]. Velocity-selective [Ami85]. Vents [Sau94]. verre [GB95]. verres [SKGRF95, Tou85, VPF95]. vertical [BBDC07]. very [Dor84]. VI [CVR^{+}95, VGG97]. Vibrational [CG84]. vibrationelle [PTZR95]. vibrationnal [FDS94]. vibrations [Fra91, Jus88]. vide [CT85b, KVC^{+}97, VHBL97]. vie [BRG^{+}97]. view [Cav00b]. vins [LB02]. violation [Bou85]. Virgo [Lor98]. visco [Vol97b, VGM97, Vol97c, Vol97d, Vol97e, VGG97]. visco-elastic [Vol97b, VGM97, Vol97c, Vol97d, Vol97e, VGG97]. visco-élastique [Vol97b, VGM97, Vol97c, Vol97d, Vol97e, VGG97]. Viscoélasticité [Que84]. viscosités [Vol97e]. viscosity [Vol97e]. Visible [PFC91, NLM^{+}94]. vitesse [SL95]. vitreuse [VGG97]. vitreux [Mél97]. vivants [SM89]. VLT [PFC91]. Vol. [Ano92, Ano94, Ano95a, Ano95b, Ano97a, Ano97b, Ano05, Ano06, Ano07a, Ano07b, Ano08a, Ano08b, Ano09a, Ano09b]. vortex [LEPMS00]. vortices [LEPMS00]. vue [Cav00b]. VUV [FS92, FDSF92, FHF92, HSD^{+}94, Lem92, MNO^{+}94, NLM^{+}94, NFV^{+}94, PCFC07, PJ97, PC92, SCR^{+}94, Vit92].

Waals [FSBD95, LLC^{+}95]. Wannier [Com95, Voi97]. waste [Sch00, Ler00]. water [FDS94]. watts [BBC^{+}95a]. Wave [GCT^{+}97, Blo85, Bou02, BDT85, Bri85, Jac07, Bou02]. Wave-Mixing [GCT^{+}97]. Wave-packet [Bou02]. wavelength [Koh85]. waves [And84, Bri98, Cla84a, Cla84b, Fre85, JBTV07, Lar06, ZB90]. Weak [PT88]. well [AB84]. wells [GBC^{+}07]. Wetting [DSS87, QdR98, Pat96]. Wheeler [Jac07, JWG^{+}07]. whispering [NBCB^{+}07]. white [Mey91]. Winding [Mon95]. wire [Saf97]. wires [Gog94, Pie01, Tor02]. within [Per05]. world [BDLM09].

X [AB94, ABB94, Ben92, Boc92, BH92, BDL^{+}92, CCJ^{+}94, CRKP94, CGST92, CPDMR94, Dav92, DMFR94, Dhe90, Fou92, Gau92, GRA^{+}94, GABG97, GGP84, GARG94, HG94, IS94, JJC84, Joy92, JP99, KCK^{+}92, KVC^{+}97, KSG^{+}90, Kli94, KLC^{+}97, LS94, LBS^{+}92, LRRVT92, Las90, MRG^{+}88, NLNA94, NG94, PP92b, Pou97, RMA94, RCK^{+}94, RKCP97, Rou92, SBE92, Sch92, SLI^{+}94, SPJP92, Tro94, Tro02, VKCP97, ZAC^{+}97, vB89]. X-mous [Joy92]. X-ray [GGP84, JJC84, JP99, KSG^{+}90, Las90, MRG^{+}88, vB89]. X-U.V. [BDL^{+}92]. X-UV [AB94, Ben92, Dhe90, KLC^{+}97, Sch92, Tro94, ZAC^{+}97]. Xe [JLS^{+}95]. XeCl [HTBL94, PBL^{+}94]. xénon [BSMS92]. xMxCuO [Kop88]. XUV [Lem92, Men94, Nen92, NG94, Sau92].

YAG [MVCF95, BBC^{+}95a, CB95, NVJ^{+}95]. YAG/Nd [MVCF95, CB95]. YAlO [GJJR95]. Yb [HJK^{+}07]. YBa [MRG^{+}88, Ver88]. YBaCuO [CDMC92]. YBCO [DY88]. years [Str85].

Z [AFH94, CFHF94]. Z-pinch [AFH94, CFHF94]. ZGP [KEH07]. zones [TPD^{+}97].

References

Averbuch:1988:SNM

[AA88] P. Averbuch and M. Avignon. Superconductivity: new models, new applications. Annales de Physique (Paris, France), 13(5):325–??, ???? 1988. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1988/05/anphys_1988__13_5_325_
0/anphys_1988__13_5_325_0.html;   http://www.annphys.org/
articles/anphys/pdf/1988/05/anphys_1988__13_5_325_0.pdf.

Alamanos:1991:RGN

[AA91] N. Alamanos and F. Auger. Résonances géantes dans les noyaux chauds. (French) []. Annales de Physique (Paris, France), 16(1):41–??, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1991/01/anphys_1991__16_1_41_0/anphys_
1991__16_1_41_0.html;   http://www.annphys.org/articles/
anphys/pdf/1991/01/anphys_1991__16_1_41_0.pdf;   http:
//www.annphys.org/articles/anphys/ref/1991/01/anphys_1991_
_16_1_41_0/anphys_1991__16_1_41_0.html.

Allegrini:1985:CCE

[AAB^{+}85] M. Allegrini, G. Alzetta, P. Bicchi, S. Gozzini, and L. Moi. Control of chemical equilibrium through optical pumping. Annales de Physique (Paris, France), 10(6):883–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/06/anphys_1985__10_6_883_0/anphys_1985__10_6_883_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
06/anphys_1985__10_6_883_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1985/06/anphys_1985__10_6_883_0/anphys_
1985__10_6_883_0.html.

Angleraud:1997:EED

[AAC97] B. Angleraud, J. Aubreton, and A. Catherinot. Étude de l’expansion du panache plasma formé lors de l’interaction entre un laser ultraviolet et une cible de nitrure de bore ; application à la réalisation de couches minces. (French) []. Annales de Physique (Paris, France), 22(C1):C1–121, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972222/

19972222.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972222.pdf.

Allard:1984:ARG

[AB84] N. F. Allard and Y. G. Biraud. Alkali-rare gas interaction potential modelisation by square-well potentials: physical interpretation. (French) []. Annales de Physique (Paris, France), 9(4):585–??, ???? 1984. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1984/04/anphys_1984__9_4_585_0/anphys_1984__9_4_585_0.
html;   http://www.annphys.org/articles/anphys/pdf/1984/
04/anphys_1984__9_4_585_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1984/04/anphys_1984__9_4_585_0/anphys_
1984__9_4_585_0.html.

Andre:1994:OBF

[AB94] J.-M. André and R. Barchewitz. Optique de Bragg–Fresnel multicouche pour le rayonnement x-uv. (French) []. Annales de Physique (Paris, France), 19(C1):C1–17, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941903/
19941903.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941903.pdf.

Andre:1994:RXM

[ABB94] J.-M. André, R. Barchewitz, and C. Bonnelle. Rayonnement X mou produit à l’aide d’électrons relativistes traversant des multicouches. (French) []. Annales de Physique (Paris, France), 19(C1):C1–33, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941905/
19941905.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941905.pdf.

Abid:1994:EIT

[Abi94] Chérifa Abid. Études des instabilités thermiques dans un écoulement ouvert horizontal, soumis à un phénomène de convection mixte. (French) []. Annales de Physique (Paris, France), 19(6):739–??, ???? 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1994/06/anphys_1994__19_6_739_
0/anphys_1994__19_6_739_0.html;   http://www.annphys.org/
articles/anphys/pdf/1994/06/anphys_1994__19_6_739_0.pdf;

 http://www.annphys.org/articles/anphys/ref/1994/06/anphys_
1994__19_6_739_0/anphys_1994__19_6_739_0.html.

Astruc:1984:ETR

[ABS84] J. P. Astruc, R. Barbe, and J. P. Schermann. Electron transfer reactions between excited atoms and molecules. Annales de Physique (Paris, France), 9(4):637–??, ???? 1984. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1984/
04/anphys_1984__9_4_637_0/anphys_1984__9_4_637_0.html;
 http://www.annphys.org/articles/anphys/pdf/1984/04/anphys_
1984__9_4_637_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1984/04/anphys_1984__9_4_637_0/anphys_1984_
_9_4_637_0.html.

Antoine:1997:GHO

[ACG^{+}97] P. Antoine, B. Carré, M. Gaarde, M. Lewenstein, A. L’Huillier, A. Persson, and P. Salières. Génération d’harmoniques d’ordre élevé dans les gaz : résultats récents. (French) []. Annales de Physique (Paris, France), 22(C1):C1–45, ???? 1997. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1997/
07/19972207/19972207.html;   http://www.annphys.org/articles/
anphys/pdf/1997/07/19972207.pdf.

Avrillier:1994:APL

[AEA^{+}94] S. Avrillier, D. Ettori, G. Allegre, F. Hor, and M. Desgeorges. Applications potentielles des lasers à excimères en neurochirurgie. (French) []. Annales de Physique (Paris, France), 19(C1):C1–195, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941949/
19941949.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941949.pdf.

Antoni:1994:CMS

[AFF^{+}94] F. Antoni, E. Fogarassy, C. Fuchs, B. Prévot, and J. P. Stoquert. Couches minces de Si_{1−x}Ge_{x} déposées par la technique d’ablation laser. (French) []. Annales de Physique (Paris, France), 19(C1):C1–39, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941908/

19941908.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941908.pdf.

Andre:1994:CHD

[AFH94] V. André, C. Fleurier, and D. Hong. Correction haute densité par effet des microchamps électriques dans un plasma Zpinch de deutérium. (French) []. Annales de Physique (Paris, France), 19(C1):C1–35, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941906/
19941906.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941906.pdf.

Akamatsu:1994:OEM

[AGF94]  S. Akamatsu, M. Ginibre, and G. Faivre.  Observation

expérimentale de la morphologie ‘en algue’ et d’états asymétriques appariés en solidification directionnelle de films minces d’alliages CBr_{4}-C_{2}Cl_{6}. (French) []. Annales de Physique (Paris, France), 19(6):645–??, ???? 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1994/06/anphys_1994__19_6_645_
0/anphys_1994__19_6_645_0.html;   http://www.annphys.org/
articles/anphys/pdf/1994/06/anphys_1994__19_6_645_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1994/06/anphys_
1994__19_6_645_0/anphys_1994__19_6_645_0.html.

Angleraud:1994:IDP

[AGG^{+}94] B. Angleraud, C. Germain, C. Girault, C. Champeaux, J. Aubreton, and A. Catherinot. Imagerie du panache de photoablation par laser U.V. de différents matériaux cibles. (French) []. Annales de Physique (Paris, France), 19(C1):C1–37, October 1994. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1994/
07/19941907/19941907.html;   http://www.annphys.org/articles/
anphys/pdf/1994/07/19941907.pdf.

Acker:1991:OPP

[AGTS91] A. Acker, F. Gleizes, R. Tylenda, and B. Stenholm. The observed position of planetary nebulae nuclei on the luminosity — temperature diagram. Annales de Physique (Paris, France), 16(4):361–??, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1991/04/anphys_1991__16_4_361_0/anphys_1991__16_4_

361_0.html;   http://www.annphys.org/articles/anphys/pdf/
1991/04/anphys_1991__16_4_361_0.pdf;   http://www.annphys.
org/articles/anphys/ref/1991/04/anphys_1991__16_4_361_
0/anphys_1991__16_4_361_0.html.

Abdel-Hafez:1992:EDC

[AHO92] A. M. Abdel-Hafez and A.-S. F. Obada. Effect of detuning and coherence phase on normal squeezing. Annales de Physique (Paris, France), 17(4):337–??, ???? 1992. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1992/04/anphys_1992__17_4_337_0/anphys_1992__17_4_337_
0.html;   http://www.annphys.org/articles/anphys/pdf/1992/
04/anphys_1992__17_4_337_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1992/04/anphys_1992__17_4_337_0/anphys_
1992__17_4_337_0.html.

Abdel-Hafez:1995:PPM

[AHOE95] A. M. Abdel-Hafez, A.-S. F. Obada, and A. H. Essawy. Phase properties of multiphoton Jaynes–Cummings model. Annales de Physique (Paris, France), 20(1):47–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
01/ann01021995/ann01021995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/01/ann01021995.pdf.

Ashkenazi:1988:CEB

[AK88] J. Ashkenazi and C. G. Kuper. Correlated-electron bands and interlayer pairing in high-temperature superconductors. Annales de Physique (Paris, France), 13(5):407–??, ???? 1988. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1988/05/anphys_1988__13_5_407_0/anphys_1988__13_5_407_
0.html;   http://www.annphys.org/articles/anphys/pdf/1988/
05/anphys_1988__13_5_407_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1988/05/anphys_1988__13_5_407_0/anphys_
1988__13_5_407_0.html.

Alastuey:1986:PED

[Ala86] A. Alastuey. Propriétés d’équilibre du plasma classique à une composante en trois et deux dimensions. (French) []. Annales de Physique (Paris, France), 11(6):653–??, ???? 1986. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1986/06/anphys_1986__11_6_653_0/anphys_1986__11_6_653_
0.html;   http://www.annphys.org/articles/anphys/pdf/1986/
06/anphys_1986__11_6_653_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1986/06/anphys_1986__11_6_653_0/anphys_
1986__11_6_653_0.html.

Alexander:1987:ADE

[Ale87] J. M. Alexander. Angular distributions of evaporated particles, fission and intermediate-mass fragments: on the search for consistent models. Annales de Physique (Paris, France), 12(6):603–??, ???? 1987. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1987/06/anphys_1987__12_6_603_0/anphys_1987__12_6_603_
0.html;   http://www.annphys.org/articles/anphys/pdf/1987/
06/anphys_1987__12_6_603_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1987/06/anphys_1987__12_6_603_0/anphys_
1987__12_6_603_0.html.

Albert:1995:MLV

[ALE95] O. Albert, J. C. Loulergue, and J. Etchepare. Mesure linéaire de la variation de phase associé à une non-linéarité Kerr. (French) []. Annales de Physique (Paris, France), 20(5–6):589–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05251995/ann05251995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05251995.pdf.

Aminoff:1985:VSO

[Ami85] C. G. Aminoff. Velocity-selective optical pumping and collision effects. Annales de Physique (Paris, France), 10(6):995–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/06/anphys_1985__10_6_995_0/anphys_1985__10_6_995_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
06/anphys_1985__10_6_995_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1985/06/anphys_1985__10_6_995_0/anphys_
1985__10_6_995_0.html.

Andrews:1984:LWG

[And84] D. G. Andrews. Lectures on waves in geophysical fluids. Annales de Physique (Paris, France), 9(2):307–??, ???? 1984. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1984/
02/anphys_1984__9_2_307_0/anphys_1984__9_2_307_0.html;
 http://www.annphys.org/articles/anphys/pdf/1984/02/anphys_
1984__9_2_307_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1984/02/anphys_1984__9_2_307_0/anphys_1984_
_9_2_307_0.html.

Anonymous:1984:PHL

[Ano84] Anonymous. Problèmes d’hydrodynamique liés à la physique des atmosphères stellaires — avant-propos. (French) []. Annales de Physique (Paris, France), 9(2):R1, ???? 1984. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1984/
02/anphys_1984__9_2_R3_0/anphys_1984__9_2_R3_0.html;
http://www.annphys.org/articles/anphys/pdf/1984/02/anphys_
1984__9_2_R3_0.pdf.

Anonymous:1990:ECL

[Ano90] Anonymous. Erratum — colloque sur les lasers et l’optique quantique C1-90 des Annales de Physique. (French) []. Annales de Physique (Paris, France), 15(3):285–??, ???? 1990. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1990/03/anphys_1990__15_3_285_0/anphys_1990__15_3_285_
0.html;   http://www.annphys.org/articles/anphys/pdf/1990/
03/anphys_1990__15_3_285_0.pdf.

Anonymous:1991:Ea

[Ano91a] Anonymous. Errata. Annales de Physique (Paris, France), 16(6):707–??, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1991/06/anphys_1991__16_6_707_
0/anphys_1991__16_6_707_0.html;   http://www.annphys.org/
articles/anphys/pdf/1991/06/anphys_1991__16_6_707_0.pdf.

Anonymous:1991:Eb

[Ano91b] Anonymous. Errata. Annales de Physique (Paris, France), 16(6):707–??, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1991/06/anphys_1991__16_6_707_
1/anphys_1991__16_6_707_1.html;   http://www.annphys.org/
articles/anphys/pdf/1991/06/anphys_1991__16_6_707_1.pdf.

Anonymous:1992:VNC

[Ano92] Anonymous. Vol. 17, No. C1 (juin 1992). Annales de Physique (Paris, France), ??(??):??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

Anonymous:1994:VNC

[Ano94] Anonymous. Vol. 19, No. C1 (octobre 1994). Annales de Physique (Paris, France), 19(C1):??, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

Anonymous:1995:VNa

[Ano95a] Anonymous. Vol. 20, No. 5–6 (1995). Annales de Physique (Paris, France), 20(5–6):??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

Anonymous:1995:VNb

[Ano95b] Anonymous. Vol. 29, No. 2–3 (2004). Annales de Physique (Paris, France), 29(2–3):??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

Anonymous:1997:VN

[Ano97a] Anonymous. Vol. 22, No. 1–2 (1997). Annales de Physique (Paris, France), 22(1–2):??, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

Anonymous:1997:VNC

[Ano97b] Anonymous. Vol. 22, No. C1 (février–avril 1997). (French) []. Annales de Physique (Paris, France), 22(C1):??, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

Anonymous:2005:VN

[Ano05] Anonymous. Vol. 30, No. 4–5 (2005). Annales de Physique (Paris, France), 30(4–5):??, ???? 2005. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

Anonymous:2006:VN

[Ano06] Anonymous. Vol. 31, No. 4–5 (2006). Annales de Physique (Paris, France), 31(4–5):??, ???? 2006. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

Anonymous:2007:VNa

[Ano07a] Anonymous. Vol. 32, No. 2–3 (2007). Annales de Physique (Paris, France), 32(2–3):??, ???? 2007. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

Anonymous:2007:VNb

[Ano07b] Anonymous. Vol. 32, No. 4–5 (2007). Annales de Physique (Paris, France), 32(2–3):??, ???? 2007. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

Anonymous:2008:VNa

[Ano08a] Anonymous. Vol. 33, No. 1–2 (2008). Annales de Physique (Paris, France), 33(1–2):??, ???? 2008. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

Anonymous:2008:VNb

[Ano08b] Anonymous. Vol. 33, No. 4–5 (2008). Annales de Physique (Paris, France), 33(1–2):??, ???? 2008. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

Anonymous:2009:VNa

[Ano09a] Anonymous. Vol. 34, No. 2–3 (2009). Annales de Physique (Paris, France), 34(2–3):??, ???? 2009. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

Anonymous:2009:VNb

[Ano09b] Anonymous. Vol. 34, No. 4–5 (2009). Annales de Physique (Paris, France), 34(2–3):??, ???? 2009. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

Arnould:1990:NRA

[AR90] M. Arnould and M. Rayet. Nuclear reactions in astrophysics. Annales de Physique (Paris, France), 15(3):183–??, ???? 1990. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1990/03/anphys_1990__15_3_183_0/anphys_1990__15_3_183_
0.html;   http://www.annphys.org/articles/anphys/pdf/1990/
03/anphys_1990__15_3_183_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1990/03/anphys_1990__15_3_183_0/anphys_
1990__15_3_183_0.html.

Alamanos:1996:RRE

[ARC96] N. Alamanos and P. Roussel-Chomaz. Recent results on elastic and inelastic scattering. Annales de Physique (Paris, France), 21(6):601–??, ???? 1996. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1996/06/ann06021996/ann06021996.html;
 http://www.annphys.org/articles/anphys/pdf/1996/06/ann06021996.
pdf.

Arimondo:1985:ARO

[AS85] E. Arimondo and G. W. Series. Acoustic resonance in an optically pumped Rb vapour. Annales de Physique (Paris, France), 10(6):1045–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1985/06/anphys_1985__10_6_1045_0/anphys_
1985__10_6_1045_0.html;   http://www.annphys.org/articles/
anphys/pdf/1985/06/anphys_1985__10_6_1045_0.pdf;   http:
//www.annphys.org/articles/anphys/ref/1985/06/anphys_1985_
_10_6_1045_0/anphys_1985__10_6_1045_0.html.

Arnould:2001:PNA

[AS01] M. Arnould and M. Samyn. La physique nucléaire en astrophysique. (French) [Nuclear physics in astrophysics]. Annales de Physique (Paris, France), 26(5):1–??, ???? 2001. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic).  URL http://www.annphys.org/articles/anphys/
abs/2001/05/annales052001/annales052001.html;   http:/
/www.annphys.org/articles/anphys/pdf/2001/05/annales052001.
pdf.

Akjouj:1993:ITS

[ASD93] A. Akjouj, B. Sylla, and L. Dobrzynski. Introduction à une théorie des systèmes composites : exemples simples de matériaux lamellaires. (French) []. Annales de Physique (Paris, France), 18(5):363–??, ???? 1993. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1993/05/anphys_1993__18_5_363_
0/anphys_1993__18_5_363_0.html;   http://www.annphys.org/
articles/anphys/pdf/1993/05/anphys_1993__18_5_363_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1993/05/anphys_
1993__18_5_363_0/anphys_1993__18_5_363_0.html.

Autric:1994:TMM

[ASP94] M. Autric, T. Sarnet, and C. Prat. Traitements de matériaux métalliques et céramiques par laser à excimères. (French) []. Annales de Physique (Paris, France), 19(C1):C1–285, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1994/07/19941961/19941961.html;   http://www.annphys.org/
articles/anphys/pdf/1994/07/19941961.pdf.

Avrillier:1992:EAU

[ATE92] S. Avrillier, E. Tinet, and D. Ettori. État actuel de l’utilisation des lasers à excimères en médecine. (French) []. Annales de Physique (Paris, France), ??(??):13–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921703/
19921703.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921703.pdf.

Andre:1992:QSS

[AVC92] J. C. Andre, M. L. Viriot, and S. Corbel. De quelles sources spécifiques a-t-on besoin en transformation de la matière ?. (French) []. Annales de Physique (Paris, France), ??(??):117–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1992/07/19921724/19921724.html;   http://www.annphys.
org/articles/anphys/pdf/1992/07/19921724.pdf.

Averbeck:1997:CIU

[Ave97a] D. Averbeck. Carcinogenesis induced by UV and ionizing radiation. Annales de Physique (Paris, France), 22(C1):C1–157, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1997/07/19972235/19972235.html;   http://www.annphys.org/
articles/anphys/pdf/1997/07/19972235.pdf.

Averbuch:1997:NES

[Ave97b] P. Averbuch. Note des Éditeurs scientifiques. (French) []. Annales de Physique (Paris, France), 22(1–2):1–??, ???? 1997. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1997/
01/ann01011997/ann01011997.html;   http://www.annphys.
org/articles/anphys/pdf/1997/01/ann01011997.pdf.

Ayral:1995:ORN

[Ayr95] J.-L. Ayral. OPO en régime nanoseconde ; état de l’art et applications. (French) []. Annales de Physique (Paris, France), 20 (5–6):501–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1995/05/ann05041995/ann05041995.html;
 http://www.annphys.org/articles/anphys/pdf/1995/05/ann05041995.
pdf.

Baudinaud:1992:IRL

[BA92] V. Baudinaud and M. Autric. Interaction rayonnement lasermatière et applications potentielles des lasers à excimères. (French) []. Annales de Physique (Paris, France), ??(??):1–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1992/07/19921701/19921701.html;   http://www.annphys.org/
articles/anphys/pdf/1992/07/19921701.pdf.

Babinet:1986:CEI

[Bab86] R. Babinet. Collisions entre ions lourds à haute énergie : approche expérimentale. (French) []. Annales de Physique (Paris, France), 11(2):113–??, ???? 1986. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1986/02/anphys_1986__11_2_113_
0/anphys_1986__11_2_113_0.html;   http://www.annphys.org/
articles/anphys/pdf/1986/02/anphys_1986__11_2_113_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1986/02/anphys_
1986__11_2_113_0/anphys_1986__11_2_113_0.html.

Balseiro:1988:SCT

[BAGR88] C. A. Balseiro, B. Alascio, E. Gagliano, and A. Rojo. Superconductivity and charge transfer excitations in high T_{c} superconductors. Annales de Physique (Paris, France), 13(5):415–??, ???? 1988. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1988/05/anphys_1988__13_5_415_0/anphys_1988__13_5_415_
0.html;   http://www.annphys.org/articles/anphys/pdf/1988/
05/anphys_1988__13_5_415_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1988/05/anphys_1988__13_5_415_0/anphys_
1988__13_5_415_0.html.

Balian:1996:MSH

[Bal96] Roger Balian. De la mécanique statistique hors équilibre aux équations de transport. (French) [Nuclei in collision from nonequilibrium statistical mechanics to transport equations]. Annales de Physique (Paris, France), 21(5):437–??, ???? 1996. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1996/
05/ann05011996/ann05011996.html;   http://www.annphys.
org/articles/anphys/pdf/1996/05/ann05011996.pdf.

Barreau:2000:FNB

[Bar00] G. Barreau. Fission nucléaire de basse énergie et réactions induites par neutrons. (French) [Low-energy nuclear fission and neutron-induced reactions]. Annales de Physique (Paris, France), 25(2):11–??, ???? 2000. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/2000/02/chap1/chap1.html;   http:/
/www.annphys.org/articles/anphys/pdf/2000/02/chap1.pdf.

Basdevant:1985:SRQ

[BB85] J. L. Basdevant and S. Boukraa. Semi-relativistic quark-antiquark potential models. Annales de Physique (Paris, France), 10 (5):475–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1985/05/anphys_1985__10_5_475_
0/anphys_1985__10_5_475_0.html;   http://www.annphys.org/
articles/anphys/pdf/1985/05/anphys_1985__10_5_475_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1985/05/anphys_
1985__10_5_475_0/anphys_1985__10_5_475_0.html.

Barillet:1995:ISF

[BBC^{+}95a] R. Barillet, A. Brillet, R. Chiche, F. Cleva, L. Latrach, and C. N. Man. Injection et stabilisation en fréquence d’un laser ND:YAG de 2 watts. (French) []. Annales de Physique (Paris, France), 20(5–6):585–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05231995/ann05231995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05231995.pdf.

Belin:1995:MPO

[BBC^{+}95b] C. Belin, D. Bonvallot, C. Carre, C. Croutxé, C. Garel, C. Ley, L. Simonin, C. Turck, and D. J. Lougnot. Les matériaux photopolymères en optique. (French) []. Annales de Physique (Paris, France), 20(5–6):717–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05661995/ann05661995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05661995.pdf.

Briaudeau:1995:RSD

[BBD95] S. Briaudeau, D. Bloch, and M. Ducoy. Résonance sub-Doppler spécifique des cellules minces. (French) []. Annales de Physique (Paris, France), 20(5–6):647–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05541995/ann05541995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05541995.pdf.

Boucher:2007:SEL

[BBDC07] Y. G. Boucher, V. Bardinal, E. Daran, and T. Camps. Spontaneous emission and lateral photo-detection in vertical microcavities. Annales de Physique (Paris, France), 32(2–3):111–??, ???? 2007. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/2007/02/ann02200720/ann02200720.html;   http://www.
annphys.org/articles/anphys/pdf/2007/02/ann02200720.pdf;
 http://www.annphys.org/articles/anphys/ref/2007/02/ann02200720/
ann02200720.html.

Breuil:1995:COE

[BBF^{+}95] N. Breuil, A. Barthélémy, C. Froehly, M. Ariaudo, A. Reineix, and B. Jecko. Caractérisation opto-électronique de CdTe polycristalin. (French) []. Annales de Physique (Paris, France), 20(5–6):629–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1995/05/ann05451995/ann05451995.html;   http://www.
annphys.org/articles/anphys/pdf/1995/05/ann05451995.pdf.

Billy:1985:AIM

[BBG^{+}85] N. Billy, M. Broyer, B. Girard, G. Gouedard, J. C. Lehmann, and J. Vigué. Application des idées et des méthodes du pompage optique à l’étude des molécules. (French) []. Annales de Physique (Paris, France), 10(6):1101–??, ???? 1985. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1985/
06/anphys_1985__10_6_1101_0/anphys_1985__10_6_1101_0.html;
 http://www.annphys.org/articles/anphys/pdf/1985/06/anphys_
1985__10_6_1101_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1985/06/anphys_1985__10_6_1101_0/anphys_1985_
_10_6_1101_0.html.

Basrour:1995:ADP

[BBH95] S. Basrour, S. Ballandras, and D. Hauden. Application du procédé LIGA en micro-optique. (French) []. Annales de Physique (Paris, France), 20(5–6):693–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05631995/ann05631995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05631995.pdf.

Boucher:2007:CPM

[BBL07] Y. G. Boucher, F. F. L. Bentivegna, and I. L. Lyubchanskii. Couplonic parameters of a 1D magneto-photonic crystal. Annales de Physique (Paris, France), 32(2–3):107–??, ???? 2007. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2007/
02/ann02200719/ann02200719.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200719.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200719/
ann02200719.html.

Beck:1994:ECO

[BBS94] L. Beck, C. Bizeuil, and P. Stemmler. Étalonnage de cristaux organiques auprès du rayonnement synchrotron de Super ACO. (French) []. Annales de Physique (Paris, France), 19(C1):C1–43, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/1994/07/19941910/19941910.html;   http://www.
annphys.org/articles/anphys/pdf/1994/07/19941910.pdf.

Boilot:1995:GHO

[BCB^{+}95] J.-P. Boilot, F. Chaput, J. Biteau, A.-V. Véret-Lemarinier, J.- P. Galaud, D. Riehl, and Y. Lévy. Gels hybrides ‘organiquesinorganiques’. Elaboration et applications en optique. (French) []. Annales de Physique (Paris, France), 20(5–6):727–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/

1995/05/ann05671995/ann05671995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05671995.pdf.

Belouet:1997:DFM

[BCKV97] C. Belouet, D. Chambonnet, D. Keller, and C. Villard. Dépôt de films minces par ablation laser, le potentiel industriel. (French) []. Annales de Physique (Paris, France), 22(C1):C1–253, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1997/07/19972249/19972249.html;   http://www.annphys.org/
articles/anphys/pdf/1997/07/19972249.pdf.

Beaufils:2007:TMC

[BCP^{+}07] Q. Beaufils, R. Chicireanu, A. Pouderous, B. Laburthe-Tolra, E. Maréchal, L. Vernac, J.-C. Keller, and O. Gorceix. Trapping metastable chromium atoms in a crossed optical dipole trap. Annales de Physique (Paris, France), 32(2–3):171–??, ???? 2007. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2007/
02/ann02200735/ann02200735.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200735.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200735/
ann02200735.html.

Bourguignon:1997:RGI

[BDB^{+}97] B. Bourguignon, B. Dragnea, B. Boulmer, J.-P. Budin, and D. Débarre. Recuit et gravure induits par laser sur silicium : expériences et simulations numériques. (French) []. Annales de Physique (Paris, France), 22(C1):C1–229, ???? 1997. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1997/
07/19972245/19972245.html;   http://www.annphys.org/articles/
anphys/pdf/1997/07/19972245.pdf.

Bourgade:1992:DSA

[BDL^{+}92] J. L. Bourgade, D. Desenne, J. P. Lebreton, M. Louis-Jacquet, C. Reverdin, R. Sauneuf, D. Schirmann, and P. Troussel. Les diagnostics spécifiques de l’analyse des lasers X-U.V.. (French) []. Annales de Physique (Paris, France), ??(??):173–??, ???? 1992. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1992/
07/19921731/19921731.html;   http://www.annphys.org/articles/
anphys/pdf/1992/07/19921731.pdf.

Bouneau:2009:CME

[BDLM09] S. Bouneau, S. David, J.-M. Loiseaux, and O. Méplan. Construction d’un monde énergétique en 2050. (French) [Building an energy world in 2050]. Annales de Physique (Paris, France), 34(1):1–??, ???? 2009. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/2009/01/anphys012009/anphys012009.
html;   http://www.annphys.org/articles/anphys/pdf/2009/
01/anphys012009.pdf;   http://www.annphys.org/articles/
anphys/ref/2009/01/anphys012009/anphys012009.html.

Badoz:1991:SQS

[BdR91] P. A. Badoz, F. Arnaud d’Avitaya, and E. Rosencher. Search for quantum size effects in ultrathin epitaxial metallic films. Annales de Physique (Paris, France), 16(6):623–??, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1991/06/anphys_1991__16_6_623_0/anphys_1991__16_6_623_
0.html;   http://www.annphys.org/articles/anphys/pdf/1991/
06/anphys_1991__16_6_623_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1991/06/anphys_1991__16_6_623_0/anphys_
1991__16_6_623_0.html.

Brillet:1985:IGWa

[BDT85] A. Brillet, T. Damour, and Ph. Tourrenc. Introduction to gravitational wave research. Annales de Physique (Paris, France), 10(3):201–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1985/03/anphys_1985__10_3_201_
0/anphys_1985__10_3_201_0.html;   http://www.annphys.org/
articles/anphys/pdf/1985/03/anphys_1985__10_3_201_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1985/03/anphys_
1985__10_3_201_0/anphys_1985__10_3_201_0.html.

Boulanger:1995:AEE

[BDT^{+}95] P. Le Boulanger, J.-L. Douadan, M. Thuau, S. Girard, J. Margerie, B. Teisseire, and B. Viana. Absorption dans l’état excité 4I_{13/2} de cristaux dopés par l’ion Er^{+3}. (French) []. Annales de Physique (Paris, France), 20(5–6):587–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05241995/ann05241995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05241995.pdf.

Bretagne:1992:MLE

[BE92] J. Bretagne and E. Estocq. Modélisation des lasers à excimères excités par décharge. (French) []. Annales de Physique (Paris, France), ??(??):29–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921705/
19921705.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921705.pdf.

Bedogni:1991:HSS

[Bed91] R. Bedogni. The hydrodynamics of a supernova shock in the interstellar medium. Annales de Physique (Paris, France), 16 (4):375–??, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1991/04/anphys_1991__16_4_375_
0/anphys_1991__16_4_375_0.html;   http://www.annphys.org/
articles/anphys/pdf/1991/04/anphys_1991__16_4_375_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1991/04/anphys_
1991__16_4_375_0/anphys_1991__16_4_375_0.html.

Bene:1985:AKD

[Bén85a] G. J. Béné. Alfred Kastler et le développement du Groupement Ampère. (French) []. Annales de Physique (Paris, France), 10(6):597–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1985/06/anphys_1985__10_6_597_
0/anphys_1985__10_6_597_0.html;   http://www.annphys.org/
articles/anphys/pdf/1985/06/anphys_1985__10_6_597_0.pdf.

Bene:1985:ISR

[Bén85b] G. J. Béné. Introduction à la spectroscopie de la relaxation : application à l’eau biologique. (French) []. Annales de Physique (Paris, France), 10(6):1139–??, ???? 1985. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1985/
06/anphys_1985__10_6_1139_0/anphys_1985__10_6_1139_0.html;
 http://www.annphys.org/articles/anphys/pdf/1985/06/anphys_
1985__10_6_1139_0.pdf.

Benattar:1992:PCP

[Ben92] R. Benattar. Les plasmas créés par laser dans le domaine X-UV: application à la ‘calibration’. (French) []. Annales de Physique (Paris, France), ??(??):39–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921706/
19921706.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921706.pdf.

Berthier:1984:PII

[Ber84] G. Berthier. Protonated interstellar ions of diatomic molecules: quantum-mechanical predictions. Annales de Physique (Paris, France), 9(4):617–??, ???? 1984. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1984/04/anphys_1984__9_4_617_0/
anphys_1984__9_4_617_0.html;   http://www.annphys.org/
articles/anphys/pdf/1984/04/anphys_1984__9_4_617_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1984/04/anphys_
1984__9_4_617_0/anphys_1984__9_4_617_0.html.

Berman:1985:CRO

[Ber85] P. R. Berman. Current research in optical pumping. Annales de Physique (Paris, France), 10(6):985–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/06/anphys_1985__10_6_985_0/anphys_1985__10_6_985_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
06/anphys_1985__10_6_985_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1985/06/anphys_1985__10_6_985_0/anphys_
1985__10_6_985_0.html.

Berger:1987:NDS

[Ber87] E. L. Berger. Nuclear dependence of structure functions: the EMC effect. Annales de Physique (Paris, France), 12(1):1–??, ???? 1987. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1987/01/anphys_1987__12_1_1_0/anphys_1987__12_1_1_
0.html;   http://www.annphys.org/articles/anphys/pdf/1987/
01/anphys_1987__12_1_1_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1987/01/anphys_1987__12_1_1_0/anphys_
1987__12_1_1_0.html.

Berthier:1988:FDM

[Ber88] Serge Berthier. Fonction diélectrique des milieux inhomogènes. (French) []. Annales de Physique (Paris, France), 13(6):503–??, ???? 1988. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1988/06/anphys_1988__13_6_503_0/anphys_1988__13_6_
503_0.html;   http://www.annphys.org/articles/anphys/pdf/
1988/06/anphys_1988__13_6_503_0.pdf;   http://www.annphys.
org/articles/anphys/ref/1988/06/anphys_1988__13_6_503_
0/anphys_1988__13_6_503_0.html.

Boccara:1995:IOM

[BG95] A. C. Boccara and P. Gleyzes. Imagerie optique en milieu turbide. Annales de Physique (Paris, France), 20(5–6):701–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1995/05/ann05641995/ann05641995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05641995.pdf.

Brimont:2007:ESR

[BGC^{+}07] C. Brimont, M. Gallart, O. Crégut, B. Hönerlage, and P. Gilliot. Excitonic spin relaxation in GaN. Annales de Physique (Paris, France), 32(2–3):115–??, ???? 2007. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.

annphys.org/articles/anphys/abs/2007/02/ann02200721/ann02200721.
html;   http://www.annphys.org/articles/anphys/pdf/2007/
02/ann02200721.pdf;   http://www.annphys.org/articles/
anphys/ref/2007/02/ann02200721/ann02200721.html.

Bottinelli:1984:LDE

[BGP84] L. Bottinelli, L. Gouguenheim, and G. Paturel. L’échelle des distances extragalactiques. (French) []. Annales de Physique (Paris, France), 9(3):503–??, ???? 1984. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1984/03/anphys_1984__9_3_503_0/
anphys_1984__9_3_503_0.html;   http://www.annphys.org/
articles/anphys/pdf/1984/03/anphys_1984__9_3_503_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1984/03/anphys_
1984__9_3_503_0/anphys_1984__9_3_503_0.html.

Boher:1992:MNP

[BH92] P. Boher and P. Houdy. Multicouches nanométriques pour l’optique X : quelques exemples applicables aux lasers X. (French) []. Annales de Physique (Paris, France), ??(??):141–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/

1992/07/19921727/19921727.html;   http://www.annphys.org/
articles/anphys/pdf/1992/07/19921727.pdf.

Ballandras:1994:AMP

[BH94] S. Ballandras and D. Hauden. Applications aux microtechniques de la photolithographie profonde par UV et par rayonnement synchrotron. (French) []. Annales de Physique (Paris, France), 19(C1):C1–73, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941922/
19941922.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941922.pdf.

Boufendi:1994:DNS

[BHS^{+}94] L. Boufendi, J. Hermann, E. Stoffels, W. Stoffels, and A. Bouchoule. Detection in the nanometer scale of particles generated in an Ar-SiH_{4} radiofrequency low pressure discharge. Annales de Physique (Paris, France), 19(C1):C1–185, October 1994. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1994/
07/19941947/19941947.html;   http://www.annphys.org/articles/
anphys/pdf/1994/07/19941947.pdf.

Best:1991:MAS

[BHWS91] Monika Best, Peter H. Hauschildt, Rainer Wehrse, and Giora Shaviv. Model atmospheres for SNe II during early stages. Annales de Physique (Paris, France), 16(4):397–??, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1991/04/anphys_1991__16_4_397_0/anphys_1991__16_4_397_
0.html;   http://www.annphys.org/articles/anphys/pdf/1991/
04/anphys_1991__16_4_397_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1991/04/anphys_1991__16_4_397_0/anphys_
1991__16_4_397_0.html.

Botet:1988:TAS

[BJ88] R. Botet and R. Jullien. A theory of aggregating systems of particles: the clustering of clusters process. Annales de Physique (Paris, France), 13(3):153–??, ???? 1988. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1988/03/anphys_1988__13_3_153_0/anphys_1988__13_3_153_
0.html;   http://www.annphys.org/articles/anphys/pdf/1988/

03/anphys_1988__13_3_153_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1988/03/anphys_1988__13_3_153_0/anphys_
1988__13_3_153_0.html.

Billy:2007:GAL

[BJZ^{+}07] J. Billy, V. Josse, Z. Zuo, W. Guerin, A. Aspect, and P. Bouyer. Guided atom laser: a new tool for guided atom optics. Annales de Physique (Paris, France), 32(2–3):17–??, ???? 2007. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2007/
02/ann02200701/ann02200701.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200701.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200701/
ann02200701.html.

Betts:1986:LPM

[BL86] D. S. Betts and M. Leduc. L’hélium polarisé : méthodes et applications. (French) []. Annales de Physique (Paris, France), 11(3):267–??, ???? 1986. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1986/03/anphys_1986__11_3_267_
0/anphys_1986__11_3_267_0.html;   http://www.annphys.org/
articles/anphys/pdf/1986/03/anphys_1986__11_3_267_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1986/03/anphys_
1986__11_3_267_0/anphys_1986__11_3_267_0.html.

Bolle:1992:SSG

[BL92] M. Bolle and S. Lazare. Structuration submicronique à grande échelle de surface de polymères à l’aide du rayonnement du laser à excimère. (French) []. Annales de Physique (Paris, France), ??(??):51–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921708/
19921708.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921708.pdf.

Blamont:1985:RSA

[Bla85] J. Blamont. Recherches sur le sodium atmosphérique. (French) []. Annales de Physique (Paris, France), 10(6):705–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/06/anphys_1985__10_6_705_0/anphys_1985__10_6_705_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
06/anphys_1985__10_6_705_0.pdf.

Blanc:1994:PRC

[Bla94] Catherine Le Blanc. Principes et réalisation d’une chaîne laser femtoseconde haute intensité basée sur le saphir dopé au titane. (French) []. Annales de Physique (Paris, France), 19(1):1–??, ???? 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1994/01/anphys_1994__19_1_1_0/anphys_1994__19_1_1_
0.html;   http://www.annphys.org/articles/anphys/pdf/1994/
01/anphys_1994__19_1_1_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1994/01/anphys_1994__19_1_1_0/anphys_
1994__19_1_1_0.html.

Blais:2003:AAP

[Bla03] A. Blais. Algorithmes et architectures pour ordinateurs quantiques supraconducteurs. (French) [Algorithms and architectures for superconducting quantum computers]. Annales de Physique (Paris, France), 28(5):1–??, ???? 2003. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2003/
05/ann052003/ann052003.html;   http://www.annphys.org/
articles/anphys/pdf/2003/05/ann052003.pdf;   http://www.
annphys.org/articles/anphys/ref/2003/05/ann052003/ann052003.
html.

Buil:1995:DTP

[BLD^{+}95] S. Buil, C. Lalaude, J. P. Delville, E. Freysz, and A. Ducasse. Dynamique d’une transition de phase du premier ordre induite par laser observée en mélange à quatres ondes. (French) []. Annales de Physique (Paris, France), 20(5–6):563–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05121995/ann05121995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05121995.pdf.

Bloembergen:1985:CIH

[Blo85] N. Bloembergen. Collision-induced Hanle resonances in four-wave light mixing. Annales de Physique (Paris, France), 10(6):681–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1985/06/anphys_1985__10_6_681_0/anphys_1985__10_6_
681_0.html;   http://www.annphys.org/articles/anphys/pdf/
1985/06/anphys_1985__10_6_681_0.pdf;   http://www.annphys.

org/articles/anphys/ref/1985/06/anphys_1985__10_6_681_
0/anphys_1985__10_6_681_0.html.

Boulmer-Leborgne:1994:DST

[BLTH94] C. Boulmer-Leborgne, A. L. Thomann, and J. Hermann. Direct synthesis of titanium nitride on Ti sample surface by laser-plasma. Annales de Physique (Paris, France), 19(C1):C1–277, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1994/07/19941960/19941960.html;   http://www.annphys.org/
articles/anphys/pdf/1994/07/19941960.pdf.

Boiron:1994:CIP

[BLW94] Marc Boiron, Maurice Lombardi, and Laurent Wiesenfeld. Chaos induit par transition non-adiabatique dans les molécules. (French) []. Annales de Physique (Paris, France), 19(6):759–??, ???? 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1994/06/anphys_1994__19_6_759_0/anphys_1994__19_6_759_
0.html;   http://www.annphys.org/articles/anphys/pdf/1994/
06/anphys_1994__19_6_759_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1994/06/anphys_1994__19_6_759_0/anphys_
1994__19_6_759_0.html.

Bockelee-Morvan:1989:CCN

[BM89] D. Bockelée-Morvan. La composition chimique des noyaux cométaires : bilan de plus d’un siècle d’observations. (French) []. Annales de Physique (Paris, France), 14(2):89–??, ???? 1989. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1989/
02/anphys_1989__14_2_89_0/anphys_1989__14_2_89_0.html;
 http://www.annphys.org/articles/anphys/pdf/1989/02/anphys_
1989__14_2_89_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1989/02/anphys_1989__14_2_89_0/anphys_1989_
_14_2_89_0.html.

Barsella:2007:SDL

[BMV^{+}07] A. Barsella, F. May, J.-P. Vola, A. Fort, S. Sanaur, A.-J. Attias, M. Lefebvre, and E. Rosencher. Self-doubled laser emission in an organic system. Annales de Physique (Paris, France), 32(2–3):55– ??, ???? 2007. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/2007/02/ann02200706/ann02200706.html;   http:

//www.annphys.org/articles/anphys/pdf/2007/02/ann02200706.
pdf;   http://www.annphys.org/articles/anphys/ref/2007/
02/ann02200706/ann02200706.html.

Bobin:1986:ILP

[Bob86] J. L. Bobin. Interaction laser plasma et accélération des particules. (French) []. Annales de Physique (Paris, France), 11(6):593–??, ???? 1986. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1986/06/anphys_1986__11_6_593_0/anphys_1986__11_6_
593_0.html;   http://www.annphys.org/articles/anphys/pdf/
1986/06/anphys_1986__11_6_593_0.pdf;   http://www.annphys.
org/articles/anphys/ref/1986/06/anphys_1986__11_6_593_
0/anphys_1986__11_6_593_0.html.

Bobroff:2005:ISC

[Bob05] J. Bobroff. Impuretés et systèmes corrélés. Des chaînes aux cuprates supraconducteurs. (French) [Impurities and correlated systems. Chains of cuprate superconductors]. Annales de Physique (Paris, France), 30(1):1–??, ???? 2005. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2005/
01/ann012005/ann012005.html;   http://www.annphys.org/
articles/anphys/pdf/2005/01/ann012005.pdf;   http://www.
annphys.org/articles/anphys/ref/2005/01/ann012005/ann012005.
html.

Bocher:1992:CXI

[Boc92] J.-L. Bocher. Conversion X dans l’interaction laser matière. (French) []. Annales de Physique (Paris, France), ??(??):169–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1992/07/19921730/19921730.html;   http://www.annphys.
org/articles/anphys/pdf/1992/07/19921730.pdf.

Bommier:1991:DRT

[Bom91] Véronique Bommier. Derivation of the radiative transfer equation for line polarization studies in the presence of magnetic field in astrophysics. Annales de Physique (Paris, France), 16(5):599–??, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1991/05/anphys_1991__16_5_599_0/anphys_1991__16_5_
599_0.html;   http://www.annphys.org/articles/anphys/pdf/

1991/05/anphys_1991__16_5_599_0.pdf;   http://www.annphys.
org/articles/anphys/ref/1991/05/anphys_1991__16_5_599_
0/anphys_1991__16_5_599_0.html.

Bonnet:1984:PES

[Bon84] J. J. Bonnet. Photon emission spectroscopy of highly charged ions following low energy charge exchange collisions. Annales de Physique (Paris, France), 9(4):629–??, ???? 1984. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1984/
04/anphys_1984__9_4_629_0/anphys_1984__9_4_629_0.html;
 http://www.annphys.org/articles/anphys/pdf/1984/04/anphys_
1984__9_4_629_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1984/04/anphys_1984__9_4_629_0/anphys_1984_
_9_4_629_0.html.

Borde:1985:MOD

[Bor85] Ch. J. Bordé. Méthodes optiques de détection des ondes gravitationnelles — préface. (French) []. Annales de Physique (Paris, France), 10(3):R1, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1985/
03/anphys_1985__10_3_R5_0/anphys_1985__10_3_R5_0.html;
 http://www.annphys.org/articles/anphys/pdf/1985/03/anphys_
1985__10_3_R5_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1985/03/anphys_1985__10_3_R5_0/anphys_1985_
_10_3_R5_0.html.

Borderie:1992:TSH

[Bor92] B. Borderie. Time-scales in heavy-ion collisions around the Fermi energy. Limits for thermalization of hot nuclei and multifragmentation. Annales de Physique (Paris, France), 17(4):349–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1992/04/anphys_1992__17_4_349_0/anphys_1992__17_4_349_
0.html;   http://www.annphys.org/articles/anphys/pdf/1992/
04/anphys_1992__17_4_349_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1992/04/anphys_1992__17_4_349_0/anphys_
1992__17_4_349_0.html.

Borde:1995:ACA

[Bor95] Ch. J. Bordé. Amplifications de champs atomiques par émission stimulée d’atomes. (French) []. Annales de Physique (Paris, France), 20(5–6):477–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.

annphys.org/articles/anphys/abs/1995/05/ann05011995/ann05011995.
html;   http://www.annphys.org/articles/anphys/pdf/1995/
05/ann05011995.pdf.

Bouchiat:1985:POV

[Bou85] M.-A. Bouchiat. Pompage optique et violation de parité dans l’atome. (French) []. Annales de Physique (Paris, France), 10 (6):923–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1985/06/anphys_1985__10_6_923_
0/anphys_1985__10_6_923_0.html;   http://www.annphys.org/
articles/anphys/pdf/1985/06/anphys_1985__10_6_923_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1985/06/anphys_
1985__10_6_923_0/anphys_1985__10_6_923_0.html.

Bouyer:1993:SSI

[Bou93] Jean-Philippe Bouyer. Spectral stabilization of an InGaAsP semiconductor laser by injection-locking. Annales de Physique (Paris, France), 18(2):89–??, ???? 1993. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1993/02/anphys_1993__18_2_89_0/
anphys_1993__18_2_89_0.html;   http://www.annphys.org/
articles/anphys/pdf/1993/02/anphys_1993__18_2_89_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1993/02/anphys_
1993__18_2_89_0/anphys_1993__18_2_89_0.html.

Bouchareine:1996:CFM

[Bou96] P. Bouchareine. Charles Fabry métrologue. (French) [Charles Fabry metrologist]. Annales de Physique (Paris, France), 21 (6):589–??, ???? 1996. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1996/06/ann06011996/ann06011996.html;
 http://www.annphys.org/articles/anphys/pdf/1996/06/ann06011996.
pdf.

Bournel:2000:MED

[Bou00] A. Bournel. Magnéto-électronique dans des dispositifs à semiconducteurs. (French) [Magneto-electronics in semiconductor devices]. Annales de Physique (Paris, France), 25(1):1–??, ???? 2000. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/

anphys/abs/2000/01/bournel/bournel.html;   http://www.
annphys.org/articles/anphys/pdf/2000/01/bournel.pdf.

Bouchene:2002:DIP

[Bou02] M. A. Bouchene. Dynamique et interférence de paquets d’ondes dans les atomes et dimères d’alcalins. (French) [Wave-packet dynamics and wave interference in atoms and alkali dimers]. Annales de Physique (Paris, France), 27(6):1–??, ???? 2002. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2002/
06/ann062002/ann062002.html;   http://www.annphys.org/
articles/anphys/pdf/2002/06/ann062002.pdf.

Barone:1987:EE

[BP87] V. Barone and E. Predazzi. The EMC effect. Annales de Physique (Paris, France), 12(6):525–??, ???? 1987. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1987/06/anphys_1987__12_6_525_0/anphys_1987__12_6_525_
0.html;   http://www.annphys.org/articles/anphys/pdf/1987/
06/anphys_1987__12_6_525_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1987/06/anphys_1987__12_6_525_0/anphys_
1987__12_6_525_0.html.

Boutin:2007:DSI

[BPJ^{+}07] C. Boutin, J. Plain, R. Jaffiol, R. Deturche, and P. Royer. Dyesubstrate interactions evidenced by fluorescence correlation spectroscopy. Annales de Physique (Paris, France), 32(2–3):143–??, ???? 2007. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/2007/02/ann02200728/ann02200728.html;   http://www.
annphys.org/articles/anphys/pdf/2007/02/ann02200728.pdf;
 http://www.annphys.org/articles/anphys/ref/2007/02/ann02200728/
ann02200728.html.

Brunet:1988:EPS

[BR88] Y. Brunet and M. Renard. Electrotechnical prospects for superconducting applications. Annales de Physique (Paris, France), 13(5):365–??, ???? 1988. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1988/05/anphys_1988__13_5_365_
0/anphys_1988__13_5_365_0.html;   http://www.annphys.org/
articles/anphys/pdf/1988/05/anphys_1988__13_5_365_0.pdf;

 http://www.annphys.org/articles/anphys/ref/1988/05/anphys_
1988__13_5_365_0/anphys_1988__13_5_365_0.html.

Braat:1999:CEI

[Bra99] J. Braat. Calcul efficace de l’intensité image en microscopie confocale appliqué à la lecture d’un disque optique. (French) [Efficient computation of the image intensity in confocal microscopy applied to the reading of an optical disk]. Annales de Physique (Paris, France), 24(3):31–??, ???? 1999. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1999/03/abschap3/
abschap3.html;   http://www.annphys.org/articles/anphys/
pdf/1999/03/abschap3.pdf.

Brunel:1995:EPL

[BRB^{+}95] M. Brunel, M. Ropars, F. Bretenaker, A. Le Floch, J. F. Kerdraon, G. Proux, and A. Junchat. États propres des lasers multiaxes. (French) []. Annales de Physique (Paris, France), 20(5–6):609–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1995/05/ann05351995/ann05351995.html;   http://www.
annphys.org/articles/anphys/pdf/1995/05/ann05351995.pdf.

Bretenaker:1993:TES

[Bre93] Fabien Bretenaker. Theoretical and experimental study of eigenmodes and eigenstates in ring lasers. Applications to gyrometry and to the detection of small effects. Annales de Physique (Paris, France), 18(6):449–??, ???? 1993. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1993/06/anphys_1993__18_6_449_0/anphys_1993__18_6_449_
0.html;   http://www.annphys.org/articles/anphys/pdf/1993/
06/anphys_1993__18_6_449_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1993/06/anphys_1993__18_6_449_0/anphys_
1993__18_6_449_0.html.

Brezin:1994:PRF

[Bré94] Edouard Brézin. Préface ‘Réunion Franco-Russe au col de Porte, mars 94’. (French) []. Annales de Physique (Paris, France), 19(4): R1, ???? 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/1994/04/anphys_1994__19_4_R3_0/anphys_1994_

_19_4_R3_0.html;   http://www.annphys.org/articles/anphys/
pdf/1994/04/anphys_1994__19_4_R3_0.pdf.

Breton:1997:SSL

[Bre97] V. Breton. La structure en spin longitudinal du nucléon. (French) [The structure of the nucleon longitudinal spin]. Annales de Physique (Paris, France), 22(3):283–??, ???? 1997. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1997/
03/ann0302/ann0302.html;   http://www.annphys.org/articles/
anphys/pdf/1997/03/ann0302.pdf.

Bourgeois:1997:DVP

[BRG^{+}97] F. Bourgeois, R. Reininger, P. Gürtler, J. L. Subtil, C. Jonin, and P. Laporte. Durées de vie, processus endoénergétiques et potentiels de l’exciplexe Kr^{∗} Ar. (French) []. Annales de Physique (Paris, France), 22(C1):C1–105, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972218/
19972218.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972218.pdf.

Brillet:1985:IGWb

[Bri85] A. Brillet. Interferometric gravitational wave antennae. Annales de Physique (Paris, France), 10(3):219–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/03/anphys_1985__10_3_219_0/anphys_1985__10_3_219_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
03/anphys_1985__10_3_219_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1985/03/anphys_1985__10_3_219_0/anphys_
1985__10_3_219_0.html.

Brignon:1998:CPP

[Bri98] A. Brignon. Conjugaison de phase par mélange d’ondes dans les milieux lasers solides. (French) [Phase conjugation by mixing waves in solid laser environments]. Annales de Physique (Paris, France), 23(5):1–??, ???? 1998. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1998/
05/ann051998/ann051998.html;   http://www.annphys.org/
articles/anphys/pdf/1998/05/ann051998.pdf.

Boutin:2007:EDA

[BRJ^{+}07] C. Boutin, Y. Roche, R. Jaffiol, J.-M. Millot, C. Millot, J. Plain, R. Deturche, P. Jeannesson, M. Manfait, and P. Royer. Effect of different agents onto multidrug resistant cells revealed by fluorescence correlation spectroscopy. Annales de Physique (Paris, France), 32(2–3):139–??, ???? 2007. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2007/
02/ann02200727/ann02200727.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200727.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200727/
ann02200727.html.

Brossel:1985:PDC

[Bro85] J. Brossel. Présentation du Congrès. (French) [Presentation of the Congress]. Annales de Physique (Paris, France), 10(6):529–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1985/06/anphys_1985__10_6_529_0/anphys_1985__10_6_
529_0.html;   http://www.annphys.org/articles/anphys/pdf/
1985/06/anphys_1985__10_6_529_0.pdf.

Borderie:1990:HIP

[BRTG90] B. Borderie, M. F. Rivet, and L. Tassan-Got. Heavyion peripheral collisions in the Fermi energy domain: Fragmentation processes or dissipative collisions? Annales de Physique (Paris, France), 15(4):287–??, ???? 1990. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1990/04/anphys_1990__15_4_287_0/anphys_1990__15_4_287_
0.html;   http://www.annphys.org/articles/anphys/pdf/1990/
04/anphys_1990__15_4_287_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1990/04/anphys_1990__15_4_287_0/anphys_
1990__15_4_287_0.html.

Bommier:1991:DME

[BSB91] Véronique Bommier and Sylvie Sahal-Bréchot. Derivation of the master equation for the atomic density matrix for line polarization studies in the presence of magnetic field and depolarizing collisions in astrophysics. Annales de Physique (Paris, France), 16(5):555–??, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1991/05/anphys_1991__16_5_555_

0/anphys_1991__16_5_555_0.html;   http://www.annphys.org/
articles/anphys/pdf/1991/05/anphys_1991__16_5_555_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1991/05/anphys_
1991__16_5_555_0/anphys_1991__16_5_555_0.html.

Bac:1997:URL

[BSM^{+}97] S. Bac, G. Soullié, A. Mirone, M. Idir, P. Guerin, F.-R. Ladan, and P. Troussel. Utilisation d’un réseau lamellaire multicouche comme monochromateur à bande passante étroite et comme polychromateur. (French) []. Annales de Physique (Paris, France), 22(C1):C1–147, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972231/
19972231.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972231.pdf.

Berejny:1992:CDX

[BSMS92] P. Berejny, M. Saissac, P. Millet, and Y. Salamero. Cinétique du xénon excité sur l’état 5d(5/2)_{J=3}^{0} à partir de l’analyse temporelle du deuxième continuum. (French) []. Annales de Physique (Paris, France), ??(??):47–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921707/
19921707.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921707.pdf.

Bernabeu:1985:MOO

[BTP85] E. Bernabeu, J. Tornos, and J. Pelayo. Multipole observables in an optically pumped alkali vapour. Annales de Physique (Paris, France), 10(6):1031–??, ???? 1985. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1985/
06/anphys_1985__10_6_1031_0/anphys_1985__10_6_1031_0.html;
 http://www.annphys.org/articles/anphys/pdf/1985/06/anphys_
1985__10_6_1031_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1985/06/anphys_1985__10_6_1031_0/anphys_1985_
_10_6_1031_0.html.

Bucka:1985:ATR

[Buc85] H. Bucka. Aspects of time in radiation scattering processes. Annales de Physique (Paris, France), 10(6):1079–??, ???? 1985. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1985/
06/anphys_1985__10_6_1079_0/anphys_1985__10_6_1079_0.html;
 http://www.annphys.org/articles/anphys/pdf/1985/06/anphys_
1985__10_6_1079_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1985/06/anphys_1985__10_6_1079_0/anphys_1985_
_10_6_1079_0.html.

Barille:1995:REM

[BWDR95] R. Barille, D. Wang, K. D. Dorkenoo, and G. Rivoire. Relation entre la modulation de fréquance et la diffusion de Rayleigh stimulée à large bande spectrale dans CS_{2}. (French) []. Annales de Physique (Paris, France), 20(5–6):619–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05401995/ann05401995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05401995.pdf.

Bychkov:1994:PHF

[Byc94] Yu. A. Bychkov. The problem of the half-filled Landau level in 2D electron systems. Annales de Physique (Paris, France), 19(4):399–??, ???? 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1994/04/anphys_1994__19_4_399_
0/anphys_1994__19_4_399_0.html;   http://www.annphys.org/
articles/anphys/pdf/1994/04/anphys_1994__19_4_399_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1994/04/anphys_
1994__19_4_399_0/anphys_1994__19_4_399_0.html.

Catherinot:1994:DCM

[CAA^{+}94] A. Catherinot, B. Angleraud, J. Aubreton, C. Champeaux, C. Germain, and C. Girault. Dépôt de couches minces par photoablation. (French) []. Annales de Physique (Paris, France), 19(C1):C1–237, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941955/
19941955.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941955.pdf.

Cagnac:1985:TMR

[Cag85] B. Cagnac. Les transitions multiples résonnantes. (French) []. Annales de Physique (Paris, France), 10(6):619–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/

1985/06/anphys_1985__10_6_619_0/anphys_1985__10_6_619_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
06/anphys_1985__10_6_619_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1985/06/anphys_1985__10_6_619_0/anphys_
1985__10_6_619_0.html.

Chaleard:1997:ASP

[CAG^{+}97] C. Chaleard, N. André, C. Geertsen, J. L. Lacour, C. Nouvellon, and P. Mauchien. Analyse spectroscopique d’un plasma laser produit à pression atmosphérique. (French) []. Annales de Physique (Paris, France), 22(C1):C1–113, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972221/
19972221.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972221.pdf.

Calvayrac:1998:DNL

[Cal98] F. Calvayrac. Dynamique non-linéaire des électrons de valence dans les agrégats métalliques. (French) [Nonlinear dynamics of valence electrons in metal clusters]. Annales de Physique (Paris, France), 23(3):1–??, ???? 1998. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1998/
03/ann031998/ann031998.html;   http://www.annphys.org/
articles/anphys/pdf/1998/03/ann031998.pdf.

Calvo:1999:MTS

[Cal99] F. Calvo. Manifestations de la transition solide-liquide dans les agrégats. (French) [Demonstration of the solid–liquid transition in aggregates]. Annales de Physique (Paris, France), 24 (4):1–??, ???? 1999. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1999/04/anphys041999/anphys041999.
html;   http://www.annphys.org/articles/anphys/pdf/1999/
04/anphys041999.pdf.

Camescasse:1999:RFE

[Cam99] F.-X. Camescasse. Relaxation femtoseconde des électrons dans les semiconducteurs en régime non-markovien. (French) [Femtosecond relaxation of electrons in non-Markovian semiconductors]. Annales de Physique (Paris, France), 24(5):1–??, ???? 1999. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/

anphys/abs/1999/05/ann051999/ann051999.html;   http://
www.annphys.org/articles/anphys/pdf/1999/05/ann051999.
pdf.

Canet:2004:PRD

[Can04] L. Canet. Processus de réaction-diffusion et groupe de renormalisation non perturbatif. (French) [Reaction-diffusion process and non-perturbative renormalization group]. Annales de Physique (Paris, France), 29(4):33–??, ???? 2004. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.

annphys.org/articles/anphys/abs/2004/04/ann2004052003/
ann2004052003.html;   http://www.annphys.org/articles/
anphys/pdf/2004/04/ann2004052003.pdf.

Castex:1994:LSI

[Cas94] M.-C. Castex. Des lasers sans inversion de population. (French) []. Annales de Physique (Paris, France), 19(C1):C1–137, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1994/07/19941929/19941929.html;   http://www.annphys.org/
articles/anphys/pdf/1994/07/19941929.pdf.

Cassagne:1998:MBI

[Cas98] D. Cassagne. Matériaux à bandes interdites photoniques. (French) [Photonic bandgap materials]. Annales de Physique (Paris, France), 23(4):1–??, ???? 1998. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://

www.annphys.org/articles/anphys/abs/1998/04/ann041998/
ann041998.html;   http://www.annphys.org/articles/anphys/
pdf/1998/04/ann041998.pdf.

Cavedon:2000:SIF

[Cav00a] J.-M. Cavedon. La séparation isotopique. (French) [Isotopic separation]. Annales de Physique (Paris, France), 25(2):327–??, ???? 2000. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
2000/02/chap12/chap12.html;   http://www.annphys.org/articles/
anphys/pdf/2000/02/chap12.pdf.

Cavedon:2000:RNP

[Cav00b] J.-M. Cavedon. Réacteurs nucléaires: le point de vue du physicien. (French) [Nuclear reactors: the view of the physicist]. Annales de Physique (Paris, France), 25(2):211–??, ???? 2000. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2000/
02/chap7/chap7.html;   http://www.annphys.org/articles/
anphys/pdf/2000/02/chap7.pdf.

Couderc:1995:APL

[CB95]  V. Couderc and A. Barthélémy.  Amélioration des per-

formances d’un laser YAG/Nd pompé par diode à blocage de modes à miroir non linéaire. (French) []. Annales de Physique (Paris, France), 20(5–6):573–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05171995/ann05171995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05171995.pdf.

Cardinal:1995:SIA

[CBOf95] T. Cardinal, A. Berthereau, R. Olazcuaga, and G. Le flem. Spectroscopie interférométrique appliquée à l’ingénierie de matériaux non linéaires. (French) []. Annales de Physique (Paris, France), 20(5–6):717–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1995/05/ann05651995/ann05651995.html;
 http://www.annphys.org/articles/anphys/pdf/1995/05/ann05651995.
pdf.

Cagnoux:1987:LAM

[CCH^{+}87] J. Cagnoux, P. Chartagnac, P. Hereil, M. Perez, and L. Seaman. Lagrangian analysis. Modern tool of the dynamics of solids. Annales de Physique (Paris, France), 12(5):451–??, ???? 1987. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1987/05/anphys_1987__12_5_451_0/anphys_1987__12_5_451_
0.html;   http://www.annphys.org/articles/anphys/pdf/1987/
05/anphys_1987__12_5_451_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1987/05/anphys_1987__12_5_451_0/anphys_
1987__12_5_451_0.html.

Cachoncinlle:1994:AIL

[CCJ^{+}94] C. Cachoncinlle, I. Challine, F. Jobard, E. Robert, and J. M. Pouvesle. Approche de l’interaction laser UV — cible liquide par ombroscopie X. (French) []. Annales de Physique (Paris, France), 19(C1):C1–47, October 1994. CODEN ANPHAJ.

ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941912/
19941912.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941912.pdf.

Chavel:1999:OSC

[CdBB^{+}99] P. Chavel, N. de Beaucoudrey, A. Bellemain, L. Rudolph, I. Exman, Y. Houzelle, and J. M. Vanzo. Optical switchcubes for communications in parallel processors. Annales de Physique (Paris, France), 24(3):197–??, ???? 1999. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://

www.annphys.org/articles/anphys/abs/1999/03/abschap10/
abschap10.html;   http://www.annphys.org/articles/anphys/
pdf/1999/03/abschap10.pdf.

Champeaux:1992:ILU

[CDMC92] C. Champeaux, D. Damiani, P. Marchet, and A. Catherinot. Interaction laser U.V (KrF 248 nm). Cible supraconductrice YBaCuO : — étude du panache par spectrométrie de masse — application à la réalisation de couches minces. (French) []. Annales de Physique (Paris, France), ??(??):53–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921709/
19921709.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921709.pdf.

Ciobanu:1994:ULU

[CFHF94] S. S. Ciobanu, C. Fleurier, D. Hong, and C. Fievet. Utilisation de la lumière UV d’un Z-pinch pour le diagnostic d’un arc de disjoncteur industriel. (French) []. Annales de Physique (Paris, France), 19(C1):C1–49, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941913/
19941913.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941913.pdf.

Chesnoy:1984:VER

[CG84] J. Chesnoy and G. M. Gale. Vibrational energy relaxation in liquids. Annales de Physique (Paris, France), 9(6):893–??, ???? 1984. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1984/06/anphys_1984__9_6_893_0/anphys_1984__9_6_893_0.
html;   http://www.annphys.org/articles/anphys/pdf/1984/

06/anphys_1984__9_6_893_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1984/06/anphys_1984__9_6_893_0/anphys_
1984__9_6_893_0.html.

Chen:1987:EPS

[CGM87] Y. Chen, B. Gil, and H. Mathieu. Exciton et polariton dans les semiconducteurs cubiques : étude de la réflectivité. (French) []. Annales de Physique (Paris, France), 12(3):109–??, ???? 1987. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1987/03/anphys_1987__12_3_109_0/anphys_1987__12_3_109_
0.html;   http://www.annphys.org/articles/anphys/pdf/1987/
03/anphys_1987__12_3_109_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1987/03/anphys_1987__12_3_109_0/anphys_
1987__12_3_109_0.html.

Calisti:1992:GPR

[CGST92] A. Calisti, L. Godbert, R. Stamm, and B. Talin. Génération de profils de raies X et UV dans les plasmas chauds. (French) []. Annales de Physique (Paris, France), ??(??):175–??, ???? 1992. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1992/
07/19921732/19921732.html;   http://www.annphys.org/articles/
anphys/pdf/1992/07/19921732.pdf.

Chavel:1999:I

[Cha99] Pierre Chavel. Introduction. Annales de Physique (Paris, France), 24(3):1–??, ???? 1999. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1999/03/intro/intro.html;   http:/
/www.annphys.org/articles/anphys/pdf/1999/03/intro.pdf.

Chave:2000:DST

[Cha00] J. Chave. Dynamique spatio-temporelle de la forêt tropicale. (French) [Spatiotemporal dynamics of the rainforest]. Annales de Physique (Paris, France), 25(6):1–??, ???? 2000. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2000/
06/ann06/ann06.html;   http://www.annphys.org/articles/
anphys/pdf/2000/06/ann06.pdf.

Champenois:2007:OTF

[CHH^{+}07] C. Champenois, G. Hagel, M. Houssin, M. Knoop, M. Vedel, C. Zumsteg, and F. Vedel. From optical to terahertz, frequency metrology with a single ion. Annales de Physique (Paris, France), 32(2–3):25–??, ???? 2007. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/2007/02/ann02200702/ann02200702.html;
 http://www.annphys.org/articles/anphys/pdf/2007/02/ann02200702.
pdf;   http://www.annphys.org/articles/anphys/ref/2007/
02/ann02200702/ann02200702.html.

Chiappetta:1992:PED

[Chi92] P. Chiappetta. La physique à l’échelle du TeV. (French) []. Annales de Physique (Paris, France), 17(1):1–??, ???? 1992. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1992/
01/anphys_1992__17_1_1_0/anphys_1992__17_1_1_0.html;
http://www.annphys.org/articles/anphys/pdf/1992/01/anphys_
1992__17_1_1_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1992/01/anphys_1992__17_1_1_0/anphys_1992__
17_1_1_0.html.

Chomaz:1996:SAC

[Cho96] Philippe Chomaz. Selected aspects of collective motions in nuclei: Order and disorder. Annales de Physique (Paris, France), 21(6):669–??, ???? 1996. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1996/06/ann06031996/ann06031996.html;
 http://www.annphys.org/articles/anphys/pdf/1996/06/ann06031996.
pdf.

Ciliberto:1994:SAT

[Cil94] S. Ciliberto. Several aspects of the transition to spatiotemporal chaos. Annales de Physique (Paris, France), 19(6):715–??, ???? 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1994/06/anphys_1994__19_6_715_0/anphys_1994__19_6_715_
0.html;   http://www.annphys.org/articles/anphys/pdf/1994/
06/anphys_1994__19_6_715_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1994/06/anphys_1994__19_6_715_0/anphys_
1994__19_6_715_0.html.

Cindro:1988:RNM

[Cin88] N. Cindro. Resonances and nuclear molecular configurations in heavy-ion reactions. Annales de Physique (Paris, France), 13(4):289–??, ???? 1988. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1988/04/anphys_1988__13_4_289_
0/anphys_1988__13_4_289_0.html;   http://www.annphys.org/
articles/anphys/pdf/1988/04/anphys_1988__13_4_289_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1988/04/anphys_
1988__13_4_289_0/anphys_1988__13_4_289_0.html.

Clarke:1984:ELP

[Cla84a] J. F. Clarke. Erratum — lectures on plane waves in reacting gases. Annales de Physique (Paris, France), 9(4):713–??, ???? 1984. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1984/
04/anphys_1984__9_4_713_0/anphys_1984__9_4_713_0.html;
 http://www.annphys.org/articles/anphys/pdf/1984/04/anphys_
1984__9_4_713_0.pdf. See [Cla84b].

Clarke:1984:LPW

[Cla84b] J. F. Clarke. Lectures on plane waves in reacting gases. Annales de Physique (Paris, France), 9(2):211–??, ???? 1984. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1984/
02/anphys_1984__9_2_211_0/anphys_1984__9_2_211_0.html;
 http://www.annphys.org/articles/anphys/pdf/1984/02/anphys_
1984__9_2_211_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1984/02/anphys_1984__9_2_211_0/anphys_1984_

_9_2_211_0.html. See erratum [Cla84a].

Claudin:1999:PTS

[Cla99] Ph. Claudin. La physique des tas de sable. — description phénoménologique de la propagation des contraintes dans les matériaux granulaires. (French) [physics of piles of sand — phenomenological description of stress propagation in granular materials]. Annales de Physique (Paris, France), 24(2):1–??, ???? 1999. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/1999/02/ann0299/ann0299.html;   http://www.
annphys.org/articles/anphys/pdf/1999/02/ann0299.pdf.

Clement:2008:PSD

[Clé08] D. Clément. Propriétés statiques et dynamiques d’un condensat de Bose–Einstein dans un potentiel aléatoire. (French) [Static and dynamic properties of a Bose–Einstein condensate in a random potential]. Annales de Physique (Paris, France), 33(1–2):1–??, ???? 2008. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/2008/03/ann452008/ann452008.html;   http://
www.annphys.org/articles/anphys/pdf/2008/03/ann452008.
pdf;   http://www.annphys.org/articles/anphys/ref/2008/
03/ann452008/ann452008.html.

Charon:1997:QNI

[CLM^{+}97] Y. Charon, Ph. Lanièce, R. Mastrippolito, R. Siebert, H. Tricoire, and L. Valentin. Quantitative nuclear imaging in biology. Annales de Physique (Paris, France), 22(6):707–??, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/1997/06/ann0602/ann0602.html;   http://www.
annphys.org/articles/anphys/pdf/1997/06/ann0602.pdf.

Calvez:1995:AFP

[CMF^{+}95] A. Le Calvez, S. Montant, E. Freysz, R. Maleck-Rassoul, and A. Ducasse. Analyse femtoseconde et picoseconde de la rotation des molécules de cristaux liquides en phase smectique A autour de leur grand axe. (French) []. Annales de Physique (Paris, France), 20(5–6):613–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05371995/ann05371995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05371995.pdf.

Coles:1985:HIS

[Col85] B. R. Coles. An historical introduction on spin glasses. Annales de Physique (Paris, France), 10(1):63–??, ???? 1985. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1985/
01/anphys_1985__10_1_63_0/anphys_1985__10_1_63_0.html;
 http://www.annphys.org/articles/anphys/pdf/1985/01/anphys_
1985__10_1_63_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1985/01/anphys_1985__10_1_63_0/anphys_1985_
_10_1_63_0.html.

Comet:1992:ITR

[Com92] M. Comet. Intérêt des traceurs radioactifs en cardiologie. (French) []. Annales de Physique (Paris, France), 17(2):77–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1992/02/anphys_1992__17_2_77_0/anphys_1992__17_2_77_0.
html;   http://www.annphys.org/articles/anphys/pdf/1992/
02/anphys_1992__17_2_77_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1992/02/anphys_1992__17_2_77_0/anphys_
1992__17_2_77_0.html.

Combescot:1995:WE

[Com95] Monique Combescot. On Wannier excitons. Annales de Physique (Paris, France), 20(4):457–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
04/ann04031995/ann04031995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/04/ann04031995.pdf.

Coulomb:1985:OSF

[Cou85] J. Coulomb. <<souvenirs >>. Annales de Physique (Paris, France), 10(6):535–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1985/06/anphys_1985__10_6_535_
0/anphys_1985__10_6_535_0.html;   http://www.annphys.org/
articles/anphys/pdf/1985/06/anphys_1985__10_6_535_0.pdf.

Couprie:1991:LEL

[Cou91] M. E. Couprie. Laser à électrons libres sur anneau de stockage. (French) []. Annales de Physique (Paris, France), 16(1):73–??, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1991/01/anphys_1991__16_1_73_0/anphys_1991__16_1_73_0.
html;   http://www.annphys.org/articles/anphys/pdf/1991/
01/anphys_1991__16_1_73_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1991/01/anphys_1991__16_1_73_0/anphys_
1991__16_1_73_0.html.

Couprie:1994:FEL

[Cou94] M. E. Couprie. Free electron laser in the UV. Annales de Physique (Paris, France), 19(C1):C1–127, October 1994. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1994/
07/19941928/19941928.html;   http://www.annphys.org/articles/
anphys/pdf/1994/07/19941928.pdf.

Courtois:1996:SRRa

[Cou96a] Jean-Yves Courtois. Spectroscopie Raman et Rayleigh stimulée des mélasses optiques unidimensionnelles (partie I). (French) [Stimulated Raman and Rayleigh spectroscopy of one-dimensional optical molasses (part I)]. Annales de Physique (Paris, France), 21(1):1–??, ???? 1996. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://

www.annphys.org/articles/anphys/abs/1996/01/ann011996/
ann011996.html;   http://www.annphys.org/articles/anphys/
pdf/1996/01/ann011996.pdf.

Courtois:1996:SRRb

[Cou96b] Jean-Yves Courtois. Spectroscopie Raman et Rayleigh stimulée des mélasses optiques unidimensionnelles (partie II). (French) [Stimulated Raman and Rayleigh spectroscopy of one-dimensional optical molasses (part II)]. Annales de Physique (Paris, France), 21(2):93–??, ???? 1996. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://

www.annphys.org/articles/anphys/abs/1996/02/ann021996/
ann021996.html;   http://www.annphys.org/articles/anphys/
pdf/1996/02/ann021996.pdf.

Chenais-Popovics:1994:CSX

[CPDMR94] C. Chenais-Popovics, S. Delysse, M. Merdji, and O. Rancu. Caractérisation d’une source X créée par laser. (French) []. Annales de Physique (Paris, France), 19(C1):C1–181, October 1994. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1994/
07/19941945/19941945.html;   http://www.annphys.org/articles/
anphys/pdf/1994/07/19941945.pdf.

Crampton:1985:HAA

[Cra85] S. B. Crampton. Hydrogen atom adsorption on molecular solid surfaces. Annales de Physique (Paris, France), 10(6):893–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/06/anphys_1985__10_6_893_0/anphys_1985__10_6_893_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
06/anphys_1985__10_6_893_0.pdf;   http://www.annphys.org/

articles/anphys/ref/1985/06/anphys_1985__10_6_893_0/anphys_
1985__10_6_893_0.html.

Cachoncinlle:1994:FGR

[CRKP94] C. Cachoncinlle, E. Robert, A. Khacef, and J. M. Pouvesle. Fluorescence des gaz rares à haute pression excités par flash X rapide. (French) []. Annales de Physique (Paris, France), 19(C1):C1–9, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941902/
19941902.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941902.pdf.

Cohen-Tannoudji:1985:AKA

[CT85a] C. Cohen-Tannoudji. Alfred Kastler and the Atomic Physics Group at École Normale in the 1950–1960’s. Personal recollections. Annales de Physique (Paris, France), 10(6):545–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/06/anphys_1985__10_6_545_0/anphys_1985__10_6_545_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
06/anphys_1985__10_6_545_0.pdf.

Cohen-Tannoudji:1985:AFL

[CT85b] C. Cohen-Tannoudji. Atomes dans un faisceau lumineux ou dans le vide de photons. (French) []. Annales de Physique (Paris, France), 10(6):797–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1985/06/anphys_1985__10_6_797_
0/anphys_1985__10_6_797_0.html;   http://www.annphys.org/
articles/anphys/pdf/1985/06/anphys_1985__10_6_797_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1985/06/anphys_
1985__10_6_797_0/anphys_1985__10_6_797_0.html.

Cuby:1988:DCP

[Cub88] J. G. Cuby. Les dispositifs à comptage de photons. (French) []. Annales de Physique (Paris, France), 13(1):1–??, ???? 1988. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1988/
01/anphys_1988__13_1_1_0/anphys_1988__13_1_1_0.html;
http://www.annphys.org/articles/anphys/pdf/1988/01/anphys_
1988__13_1_1_0.pdf;   http://www.annphys.org/articles/

anphys/ref/1988/01/anphys_1988__13_1_1_0/anphys_1988__
13_1_1_0.html.

Cugnon:1986:CIL

[Cug86] J. Cugnon. Collisions d’ions lourds à haute énergie, aspects théoriques. (French) []. Annales de Physique (Paris, France), 11(2):201–??, ???? 1986. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1986/02/anphys_1986__11_2_201_
0/anphys_1986__11_2_201_0.html;   http://www.annphys.org/
articles/anphys/pdf/1986/02/anphys_1986__11_2_201_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1986/02/anphys_
1986__11_2_201_0/anphys_1986__11_2_201_0.html.

Cugnon:1996:RUH

[Cug96] Joseph Cugnon. Relativistic and ultrarelativistic heavy-ion collisions. Theoretical aspects. Annales de Physique (Paris, France), 21(5):537–??, ???? 1996. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1996/05/ann05041996/ann05041996.html;
 http://www.annphys.org/articles/anphys/pdf/1996/05/ann05041996.
pdf.

Cugnon:2000:SR

[Cug00] J. Cugnon. Spallation reactions. Annales de Physique (Paris, France), 25(2):93–??, ???? 2000. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/2000/02/chap3/chap3.html;   http:
//www.annphys.org/articles/anphys/pdf/2000/02/chap3.pdf.

Curien:1985:AKH

[Cur85] Hubert Curien. Alfred Kastler, un homme et un savant. (French) [Alfred Kastler, a man and a scientist]. Annales de Physique (Paris, France), 10(6):527–??, ???? 1985. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/06/anphys_1985__10_6_527_0/anphys_1985__10_6_527_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
06/anphys_1985__10_6_527_0.pdf.

Cugnon:1989:ANI

[CV89] J. Cugnon and J. Vandermeulen. Antiproton–nucleus interaction. Annales de Physique (Paris, France), 14(1):49–??, ???? 1989. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1989/
01/anphys_1989__14_1_49_0/anphys_1989__14_1_49_0.html;
 http://www.annphys.org/articles/anphys/pdf/1989/01/anphys_
1989__14_1_49_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1989/01/anphys_1989__14_1_49_0/anphys_1989_
_14_1_49_0.html.

Cibert:1995:HSI

[CVR^{+}95] J. Cibert, E. Vanelle, R. Romestain, D. Block, R. André, Le Si Dang, J. P. Likforman, A. Alexandrou, and D. Hulin. Hétérostructures de semiconducteurs II–VI à champ piézoélectrique. (French) []. Annales de Physique (Paris, France), 20(5–6):557–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1995/05/ann05111995/ann05111995.html;   http://www.
annphys.org/articles/anphys/pdf/1995/05/ann05111995.pdf.

Debarre:1994:AGL

[DAB^{+}94] D. Débarre, A. Aliouchouche, J. Boulmer, B. Bourguignon, J. P. Budin, and S. Carrez. Ablation et gravure laser. (French) []. Annales de Physique (Paris, France), 19(C1):C1–245, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1994/07/19941956/19941956.html;   http://www.annphys.org/
articles/anphys/pdf/1994/07/19941956.pdf.

Damiani:1994:MUPb

[Dam94] D. Damiani. Micro-usinage par laser à excimères. (French) []. Annales de Physique (Paris, France), 19(C1):C1–261, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1994/07/19941958/19941958.html;   http://www.annphys.org/
articles/anphys/pdf/1994/07/19941958.pdf.

Daniel:1985:IHI

[Dan85] E. Daniel. Introduction historique sur les impuretés magnétiques dans les métaux. (French) []. Annales de Physique (Paris, France), 10(1):5–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1985/01/anphys_1985__10_1_5_0/
anphys_1985__10_1_5_0.html;   http://www.annphys.org/articles/
anphys/pdf/1985/01/anphys_1985__10_1_5_0.pdf;   http:/

/www.annphys.org/articles/anphys/ref/1985/01/anphys_1985_
_10_1_5_0/anphys_1985__10_1_5_0.html.

Darrigol:1984:GDC

[Dar84] O. Darrigol. La genèse du concept de champ quantique. (French) []. Annales de Physique (Paris, France), 9(3):433–??, ???? 1984. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1984/
03/anphys_1984__9_3_433_0/anphys_1984__9_3_433_0.html;
 http://www.annphys.org/articles/anphys/pdf/1984/03/anphys_
1984__9_3_433_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1984/03/anphys_1984__9_3_433_0/anphys_1984_
_9_3_433_0.html.

Davi:1992:RIM

[Dav92] P. Davi. Réalisation industrielle de miroirs multicouches pour optique X. (French) []. Annales de Physique (Paris, France), ??(??):177–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921733/
19921733.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921733.pdf.

David:2006:HEG

[Dav06] A. David. High efficiency GaN-based LEDs: light extraction by photonic crystals. Annales de Physique (Paris, France), 31(6):1– ??, ???? 2006. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/2006/06/ann062006/ann062006.html;   http://
www.annphys.org/articles/anphys/pdf/2006/06/ann062006.
pdf;   http://www.annphys.org/articles/anphys/ref/2006/
06/ann062006/ann062006.html.

deBergh:1984:MSP

[dB84] C. de Bergh. Molecular spectroscopy and planetary atmospheres. Annales de Physique (Paris, France), 9(4):575–??, ???? 1984. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1984/
04/anphys_1984__9_4_575_0/anphys_1984__9_4_575_0.html;
 http://www.annphys.org/articles/anphys/pdf/1984/04/anphys_
1984__9_4_575_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1984/04/anphys_1984__9_4_575_0/anphys_1984_
_9_4_575_0.html.

Djabi:2007:OBL

[DBD07] S. Djabi, H. Boudoukha, and M. Djabi. Optical bistability in a laser containing a saturable absorber. Annales de Physique (Paris, France), 32(2–3):63–??, ???? 2007. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2007/
02/ann02200708/ann02200708.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200708.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200708/
ann02200708.html.

Desbordes:1989:TRS

[DBG89] D. Desbordes, D. Brisot, and C. Guerraud. Taille et régularité de la structure du front de la détonation dans les gaz. Corrélation avec le diamètre critique de transmission de la détonation. (French) []. Annales de Physique (Paris, France), 14(6):629–??, ???? 1989. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1989/06/anphys_1989__14_6_629_0/anphys_1989__14_6_629_
0.html;   http://www.annphys.org/articles/anphys/pdf/1989/
06/anphys_1989__14_6_629_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1989/06/anphys_1989__14_6_629_0/anphys_
1989__14_6_629_0.html.

duCros:1985:EEO

[dC85] F. Teissier du Cros. Electromagnetic and elasto-optical systems for the reception or generation of gravitational radiation. Annales de Physique (Paris, France), 10(3):263–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/03/anphys_1985__10_3_263_0/anphys_1985__10_3_263_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
03/anphys_1985__10_3_263_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1985/03/anphys_1985__10_3_263_0/anphys_
1985__10_3_263_0.html.

Debarre:1987:RCP

[DC87] A. Débarre and Ph. Cahuzac. On the radiative-collision processes. Annales de Physique (Paris, France), 12(4):313–??, ???? 1987. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1987/04/anphys_1987__12_4_313_0/anphys_1987__12_4_313_
0.html;   http://www.annphys.org/articles/anphys/pdf/1987/

04/anphys_1987__12_4_313_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1987/04/anphys_1987__12_4_313_0/anphys_
1987__12_4_313_0.html.

Damiani:1992:MUPa

[DC92] D. Damiani and A. Catherinot. Micro-usinage par laser à excimères. (French) []. Annales de Physique (Paris, France), ?? (??):57–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1992/07/19921710/19921710.html;   http:
//www.annphys.org/articles/anphys/pdf/1992/07/19921710.
pdf.

deChatellus:2007:EPP

[dCP07] H. Guillet de Chatellus and J.-P. Pique. Efficiency of 2-photon polychromatic laser guide stars. Annales de Physique (Paris, France), 32(2–3):75–??, ???? 2007. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/2007/02/ann02200711/ann02200711.
html;   http://www.annphys.org/articles/anphys/pdf/2007/
02/ann02200711.pdf;   http://www.annphys.org/articles/
anphys/ref/2007/02/ann02200711/ann02200711.html.

Djerroud:2007:TOM

[DDL^{+}07] K. Djerroud, C. Daussy, O. Lopez, A. Amy-Klein, S. Briaudeau, Y. Hermier, and C. Chardonnet. Towards an optical measurement of the Boltzmann constant at the 10^{−5} level. Annales de Physique (Paris, France), 32(2–3):175–??, ???? 2007. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2007/
02/ann02200736/ann02200736.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200736.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200736/
ann02200736.html.

Dehmelt:1985:SAP

[Deh85] H. Dehmelt. Single atomic particle at rest in free space: new value for electron radius. Annales de Physique (Paris, France), 10(6):777–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1985/06/anphys_1985__10_6_777_
0/anphys_1985__10_6_777_0.html;   http://www.annphys.org/
articles/anphys/pdf/1985/06/anphys_1985__10_6_777_0.pdf;

 http://www.annphys.org/articles/anphys/ref/1985/06/anphys_
1985__10_6_777_0/anphys_1985__10_6_777_0.html.

Deiss:1997:SLB

[Dei97] J. L. Deiss. Les sources de lumière dans le bleu et le proche UV basées sur les diodes à semiconducteurs : état actuel de l’art et perspectives. (French) []. Annales de Physique (Paris, France), 22(C1):C1–21, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972203/
19972203.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972203.pdf.

Delpech:2000:SEA

[Del00] M. Delpech. Les solutions étudiées: améliorations et options innovantes. Les améliorations du cycle actuel. (French) [Studied solutions: improvements and innovative options. Improvement of the current cycle]. Annales de Physique (Paris, France), 25 (2):227–??, ???? 2000. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/2000/02/chap8/chap8.html;   http:/
/www.annphys.org/articles/anphys/pdf/2000/02/chap8.pdf.

Depommier:1987:GPS

[Dep87] P. Depommier. Grands principes de symétrie à l’épreuve de l’expérience. (French) []. Annales de Physique (Paris, France), 12(4):347–??, ???? 1987. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1987/04/anphys_1987__12_4_347_
0/anphys_1987__12_4_347_0.html;   http://www.annphys.org/
articles/anphys/pdf/1987/04/anphys_1987__12_4_347_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1987/04/anphys_
1987__12_4_347_0/anphys_1987__12_4_347_0.html.

Desesquelles:1995:MAN

[Dés95] P. Désesquelles. Multivariate analysis in nuclear physics. Annales de Physique (Paris, France), 20(1):1–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
01/ann01011995/ann01011995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/01/ann01011995.pdf.

Descouvemont:2001:DTR

[Des01] P. Descouvemont. Développements théoriques récents en astrophysique nucléaire. (French) [Recent theoretical developments in nuclear astrophysics]. Annales de Physique (Paris, France), 26(3):1–??, ???? 2001. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.

annphys.org/articles/anphys/abs/2001/03/annales032001/
annales032001.html;   http://www.annphys.org/articles/
anphys/pdf/2001/03/annales032001.pdf.

Dettwiller:1999:TCP

[Det99] L. Dettwiller. Théorie de la cohérence de polarisation avec les matrices de Jones et la sphère de Poincaré — applications. (French) [Theory of polarization coherence with Jones matrices and the Poincaré sphere — applications]. Annales de Physique (Paris, France), 24(3):51–??, ???? 1999. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1999/03/abschap5/
abschap5.html;   http://www.annphys.org/articles/anphys/
pdf/1999/03/abschap5.pdf.

Deutsch:1986:ICF

[Deu86] C. Deutsch. Inertial confinement fusion driven by intense ion beams. Annales de Physique (Paris, France), 11(1):1–??, ???? 1986. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1986/01/anphys_1986__11_1_1_0/anphys_1986__11_1_1_
0.html;   http://www.annphys.org/articles/anphys/pdf/1986/
01/anphys_1986__11_1_1_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1986/01/anphys_1986__11_1_1_0/anphys_
1986__11_1_1_0.html.

Dohnalik:1985:AKI

[DG85] T. Dohnalik and W. Gawlik. Alfred Kastler’s inspiration of the optical pumping studies in Poland. Annales de Physique (Paris, France), 10(6):605–??, ???? 1985. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/06/anphys_1985__10_6_605_0/anphys_1985__10_6_605_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
06/anphys_1985__10_6_605_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1985/06/anphys_1985__10_6_605_0/anphys_
1985__10_6_605_0.html.

Debarre:2007:C

[DG07] A. Débarre and J.-M. Gérard. COLOQ’10. Annales de Physique (Paris, France), 32(2–3):3–??, ???? 2007. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2007/
02/ann02200700/ann02200700.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200700.pdf.

Diu:1990:LTD

[DGLR90] B. Diu, C. Guthmann, D. Lederer, and B. Roulet. The laws of thermodynamics deduced from statistical mechanics. Annales de Physique (Paris, France), 15(2):163–??, ???? 1990. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1990/02/anphys_1990__15_2_163_0/anphys_1990__15_2_163_
0.html;   http://www.annphys.org/articles/anphys/pdf/1990/
02/anphys_1990__15_2_163_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1990/02/anphys_1990__15_2_163_0/anphys_
1990__15_2_163_0.html.

Diu:1991:PIL

[DGLR91] Bernard Diu, Claudine Guthmann, Danielle Lederer, and Bernard Roulet. On the physical importance of the ‘logarithmic transformation’ in statistical mechanics. Annales de Physique (Paris, France), 16(3):345–??, ???? 1991. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1991/03/anphys_1991__16_3_345_0/anphys_1991__16_3_345_
0.html;   http://www.annphys.org/articles/anphys/pdf/1991/
03/anphys_1991__16_3_345_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1991/03/anphys_1991__16_3_345_0/anphys_
1991__16_3_345_0.html.

Dhez:1990:XXU

[Dhe90] P. Dhez. X and X-UV multilayered optics: principles, fabrication methods, tests and applications. Annales de Physique (Paris, France), 15(6):493–??, ???? 1990. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1990/06/anphys_1990__15_6_493_0/anphys_1990__15_6_493_
0.html;   http://www.annphys.org/articles/anphys/pdf/1990/
06/anphys_1990__15_6_493_0.pdf;   http://www.annphys.org/

articles/anphys/ref/1990/06/anphys_1990__15_6_493_0/anphys_
1990__15_6_493_0.html.

Diu:1990:FPC

[Diu90] B. Diu. Fundamental postulate and canonical distribution in statistical mechanics. Annales de Physique (Paris, France), 15 (5):433–??, ???? 1990. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1990/05/anphys_1990__15_5_433_
0/anphys_1990__15_5_433_0.html;   http://www.annphys.org/
articles/anphys/pdf/1990/05/anphys_1990__15_5_433_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1990/05/anphys_
1990__15_5_433_0/anphys_1990__15_5_433_0.html.

Depollier:1986:LAO

[DKL86] C. Dépollier, J. Kergomard, and F. Laloe. Localisation d’anderson des ondes dans les réseaux acoustiques unidimensionnels aléatoires. (French) []. Annales de Physique (Paris, France), 11(5):457–??, ???? 1986. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1986/05/anphys_1986__11_5_457_0/anphys_1986__11_5_
457_0.html;   http://www.annphys.org/articles/anphys/pdf/
1986/05/anphys_1986__11_5_457_0.pdf;   http://www.annphys.
org/articles/anphys/ref/1986/05/anphys_1986__11_5_457_
0/anphys_1986__11_5_457_0.html.

Dupre:1989:SDD

[DLKP89] G. Dupré, J. H. Lee, R. Knystautas, and O. Péraldi. Sensibilité à la détonation de divers mélanges gazeux explosifs : hydrazine, hydrogène et hydrocarbures/air ou oxygène. (French) []. Annales de Physique (Paris, France), 14(6):607–??, ???? 1989. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1989/06/anphys_1989__14_6_607_0/anphys_1989__14_6_607_
0.html;   http://www.annphys.org/articles/anphys/pdf/1989/
06/anphys_1989__14_6_607_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1989/06/anphys_1989__14_6_607_0/anphys_
1989__14_6_607_0.html.

deLesegno:2007:AOS

[dLMDP07] B. Viaris de Lesegno, M. Mestre, F. Diry, and L. Pruvost. Atom optics with a spatial light modulator. Annales de Physique (Paris, France), 32(2–3):179–??, ???? 2007. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2007/
02/ann02200737/ann02200737.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200737.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200737/
ann02200737.html.

Dezoret:1994:CAA

[DM94] D. Dézoret and R. Marmoret. Cristal adaptatif : aspect technologique. (French) []. Annales de Physique (Paris, France), 19(C1):C1–51, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941914/
19941914.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941914.pdf.

deMicheli:1995:OPO

[dM95] M. P. de Micheli. Oscillateurs paramétriques optiques en configuration de quasi-accord de phase. (French) []. Annales de Physique (Paris, France), 20(5–6):517–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05061995/ann05061995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05061995.pdf.

Dezoret:1994:CAR

[DMFR94] D. Dézoret, R. Marmoret, A. Freund, and R. Ravelet. Cristal adaptatif refroidi pour rayonnement X. Annales de Physique (Paris, France), 19(C1):C1–25, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941904/
19941904.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941904.pdf.

Durand:1995:SML

[DNB^{+}95] P. E. Durand, G. Nogues, V. Bernard, A. Amy, and C. Chardonnet. Sélection des molécules lentes en spectroscopie sans effet Doppler à deux photons. (French) []. Annales de Physique (Paris, France), 20(5–6):601–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05311995/ann05311995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05311995.pdf.

Deguchi:1991:MES

[DNQR91] Shuji Deguchi and Nguyen-Quang-Rieu. Molecules in evolved stars. Annales de Physique (Paris, France), 16(4):413–??, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1991/04/anphys_1991__16_4_413_0/anphys_1991__16_4_413_
0.html;   http://www.annphys.org/articles/anphys/pdf/1991/
04/anphys_1991__16_4_413_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1991/04/anphys_1991__16_4_413_0/anphys_
1991__16_4_413_0.html.

Dorthe:1984:DVE

[Dor84] G. Dorthe. Dynamics of very exoergic four atom reactions. Annales de Physique (Paris, France), 9(4):621–??, ???? 1984. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1984/
04/anphys_1984__9_4_621_0/anphys_1984__9_4_621_0.html;
 http://www.annphys.org/articles/anphys/pdf/1984/04/anphys_
1984__9_4_621_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1984/04/anphys_1984__9_4_621_0/anphys_1984_
_9_4_621_0.html.

deSilans:2007:SRS

[dSLR^{+}07] T. Passerat de Silans, A. Laliotis, M. Romanelli, P. Chaves de Souza Segundo, I. Maurin, D. Bloch, M. Ducloy, A. Sarkisyan, and D. Sarkisyan. Selective reflection spectroscopy of a vapour at a calcium fluoride interface. Annales de Physique (Paris, France), 32(2–3):191–??, ???? 2007. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2007/
02/ann02200740/ann02200740.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200740.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200740/
ann02200740.html.

Dash:1987:WPC

[DSS87] J. G. Dash, M. Schick, and J. Suzanne. Wetting phenomena and crystal growth. Annales de Physique (Paris, France), 12(4):299–??, ???? 1987. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1987/04/anphys_1987__12_4_299_
0/anphys_1987__12_4_299_0.html;   http://www.annphys.org/
articles/anphys/pdf/1987/04/anphys_1987__12_4_299_0.pdf;

 http://www.annphys.org/articles/anphys/ref/1987/04/anphys_
1987__12_4_299_0/anphys_1987__12_4_299_0.html.

deUnamuno:1997:STF

[dUF97] S. de Unamuno and E. Fogarassy. Simulation thermique de la fusion et de la recristallisation par lasers à excimères du silicium amorphe pour les applications transistors en couches minces. (French) []. Annales de Physique (Paris, France), 22(C1):C1–225, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1997/07/19972243/19972243.html;   http://www.annphys.
org/articles/anphys/pdf/1997/07/19972243.pdf.

Dumont:1985:DPO

[Dum85] M. Dumont. Du pompage optique laser à l’absorption saturée résolue en temps. (French) []. Annales de Physique (Paris, France), 10(6):1015–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1985/06/anphys_1985__10_6_1015_0/anphys_
1985__10_6_1015_0.html;   http://www.annphys.org/articles/
anphys/pdf/1985/06/anphys_1985__10_6_1015_0.pdf;   http:
//www.annphys.org/articles/anphys/ref/1985/06/anphys_1985_
_10_6_1015_0/anphys_1985__10_6_1015_0.html.

Delaporte:1997:EIT

[DVT^{+}97] Ph. Delaporte, M. Voitik, Ch. Tarras, M. Sentis, O. Uteza, B. Fontaine, and B. Forestier. Excimères ioniques triatomiques d’alcalin gaz rare. (French) []. Annales de Physique (Paris, France), 22(C1):C1–107, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972219/
19972219.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972219.pdf.

Dobiasch:1985:QEE

[DW85] P. Dobiasch and H. Walther. Quantum electrodynamic effects in finite space. Annales de Physique (Paris, France), 10(6):825–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1985/06/anphys_1985__10_6_825_0/anphys_1985__10_6_
825_0.html;   http://www.annphys.org/articles/anphys/pdf/
1985/06/anphys_1985__10_6_825_0.pdf;   http://www.annphys.

org/articles/anphys/ref/1985/06/anphys_1985__10_6_825_
0/anphys_1985__10_6_825_0.html.

Dabora:1989:TEO

[DWD89] E. K. Dabora, H. Gg. Wagner, and D. Desbordes. A technique for establishing oblique detonations at high velocities. Annales de Physique (Paris, France), 14(6):637–??, ???? 1989. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1989/06/anphys_1989__14_6_637_0/anphys_1989__14_6_637_
0.html;   http://www.annphys.org/articles/anphys/pdf/1989/
06/anphys_1989__14_6_637_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1989/06/anphys_1989__14_6_637_0/anphys_
1989__14_6_637_0.html.

Ding:1988:FPO

[DY88] S. Y. Ding and Z. Yu. Flux penetration for oxide superconductor YBCO. Annales de Physique (Paris, France), 13(5):437–??, ???? 1988. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1988/05/anphys_1988__13_5_437_0/anphys_1988__13_5_437_
0.html;   http://www.annphys.org/articles/anphys/pdf/1988/
05/anphys_1988__13_5_437_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1988/05/anphys_1988__13_5_437_0/anphys_
1988__13_5_437_0.html.

Dyugaev:1994:TQN

[Dyu94] A. M. Dyugaev. Theory of quantum nondegenerated liquids. Annales de Physique (Paris, France), 19(4):433–??, ???? 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1994/04/anphys_1994__19_4_433_0/anphys_1994__19_4_433_
0.html;   http://www.annphys.org/articles/anphys/pdf/1994/
04/anphys_1994__19_4_433_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1994/04/anphys_1994__19_4_433_0/anphys_
1994__19_4_433_0.html.

Devdariani:1989:ING

[DZB89] A. Z. Devdariani, A. L. Zagrebin, and K. B. Blagoev. Interactions of noble gas atoms. Processes due to elastic scattering. Annales de Physique (Paris, France), 14(5):467–??, ???? 1989. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/

1989/05/anphys_1989__14_5_467_0/anphys_1989__14_5_467_
0.html;   http://www.annphys.org/articles/anphys/pdf/1989/
05/anphys_1989__14_5_467_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1989/05/anphys_1989__14_5_467_0/anphys_
1989__14_5_467_0.html.

Devdariani:1992:ETI

[DZB92] A. Z. Devdariani, A. L. Zagrebin, and K. Blagoev. Excitation transfer and intermultiplet transitions in collisions of He and Ne atoms at thermal energies. Annales de Physique (Paris, France), 17(5):365–??, ???? 1992. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1992/05/anphys_1992__17_5_365_0/anphys_1992__17_5_365_
0.html;   http://www.annphys.org/articles/anphys/pdf/1992/
05/anphys_1992__17_5_365_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1992/05/anphys_1992__17_5_365_0/anphys_
1992__17_5_365_0.html.

Esmiller:1994:AIL

[EB94] B. Esmiller and V. Baudinaud. Applications industrielles des lasers U.V. au Centre Commun de Recherches Louis-Blériot d’AEROSPATIALE. (French) []. Annales de Physique (Paris, France), 19(C1):C1–41, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941909/
19941909.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941909.pdf.

Esmiller:1997:RSM

[EB97] B. Esmiller and V. Baudinaud. Reformation de surfaces de matériaux métalliques par laser à excimères ou laser à vapeur de cuivre. (French) []. Annales de Physique (Paris, France), 22(C1):C1–205, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972241/
19972241.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972241.pdf.

Eliashberg:1994:MTS

[Eli94] G. M. Eliashberg. Mott transition and superconductivity. Annales de Physique (Paris, France), 19(4):353–??, ???? 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1994/04/anphys_1994__19_4_353_0/anphys_1994__19_4_353_
0.html;   http://www.annphys.org/articles/anphys/pdf/1994/
04/anphys_1994__19_4_353_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1994/04/anphys_1994__19_4_353_0/anphys_
1994__19_4_353_0.html.

Encrenaz:1988:PCA

[Enc88] Thérèse Encrenaz. La physique cométaire après l’exploration de la comète de Halley. (French) []. Annales de Physique (Paris, France), 13(6):465–??, ???? 1988. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1988/06/anphys_1988__13_6_465_
0/anphys_1988__13_6_465_0.html;   http://www.annphys.org/
articles/anphys/pdf/1988/06/anphys_1988__13_6_465_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1988/06/anphys_
1988__13_6_465_0/anphys_1988__13_6_465_0.html.

Epherre:1984:NLV

[Eph84] M. Epherre. Noyaux loin de la vallée de stabilité. production et étude expérimentale des états fondamentaux. (French) []. Annales de Physique (Paris, France), 9(5):717–??, ???? 1984. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1984/
05/anphys_1984__9_5_717_0/anphys_1984__9_5_717_0.html;
 http://www.annphys.org/articles/anphys/pdf/1984/05/anphys_
1984__9_5_717_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1984/05/anphys_1984__9_5_717_0/anphys_1984_
_9_5_717_0.html.

Elbel:1985:AMD

[EQ85] M. Elbel and R. Quad. An alternative method of determining the quadrupole moments of the chlorine nuclides. Annales de Physique (Paris, France), 10(6):675–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/06/anphys_1985__10_6_675_0/anphys_1985__10_6_675_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
06/anphys_1985__10_6_675_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1985/06/anphys_1985__10_6_675_0/anphys_
1985__10_6_675_0.html.

Esaulov:1986:EDA

[Esa86] V. A. Esaulov. Electron detachment from atomic negative ions. Annales de Physique (Paris, France), 11(5):493–??, ???? 1986. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1986/05/anphys_1986__11_5_493_0/anphys_1986__11_5_493_
0.html;   http://www.annphys.org/articles/anphys/pdf/1986/
05/anphys_1986__11_5_493_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1986/05/anphys_1986__11_5_493_0/anphys_
1986__11_5_493_0.html.

Feinberg:1988:OHP

[FABC88] D. Feinberg, M. Avignon, M. Boiron, and B. K. Chakraverty. Oxygen hole pairing in CuO_{2} planes: a quasi-one dimensional case. Annales de Physique (Paris, France), 13(5):447–??, ???? 1988. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1988/05/anphys_1988__13_5_447_0/anphys_1988__13_5_447_
0.html;   http://www.annphys.org/articles/anphys/pdf/1988/
05/anphys_1988__13_5_447_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1988/05/anphys_1988__13_5_447_0/anphys_
1988__13_5_447_0.html.

Fauve:1994:NDC

[Fau94] Stéphan Fauve. Nonlinear dynamics of coupled oscillators. Annales de Physique (Paris, France), 19(6):691–??, ???? 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1994/06/anphys_1994__19_6_691_0/anphys_1994__19_6_691_
0.html;   http://www.annphys.org/articles/anphys/pdf/1994/
06/anphys_1994__19_6_691_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1994/06/anphys_1994__19_6_691_0/anphys_
1994__19_6_691_0.html.

Freysz:1994:SFV

[FDS94] E. Freysz, Q. Du, and Y. R. Shen. Sum frequency vibrationnal spectroscopy of water molecules at interfaces. Annales de Physique (Paris, France), 19(C1):C1–95, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941924/
19941924.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941924.pdf.

Fontaine:1992:PRD

[FDSF92] B. Fontaine, Ph. Delaporte, M. Sentis, and B. Forestier. Progrès récents du groupe nouveaux lasers de l’IMFM–UM34 CNRS dans le domaine des sources UV–VUV cohérentes et incohérentes pulsées de puissance. (French) []. Annales de Physique (Paris, France), ??(??):179–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921734/
19921734.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921734.pdf.

Ferrand:2002:SNM

[Fer02]  P. Ferrand.  Le silicium nanoporeux: microstructuration

diélectrique et application aux structures photoniques avancées. (French) [Nanoporous silicon: dielectric microstructuring and application to advanced photonic structures]. Annales de Physique (Paris, France), 27(2):1–??, ???? 2002. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2002/
02/ann022002/ann022002.html;   http://www.annphys.org/
articles/anphys/pdf/2002/02/ann022002.pdf.

Frumker:1999:UPF

[FF99] E. Frumker and A. A. Friesem. Updatable phase filters for correlation with incoherent light. Annales de Physique (Paris, France), 24(3):87–??, ???? 1999. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1999/03/abschap6/
abschap6.html;   http://www.annphys.org/articles/anphys/
pdf/1999/03/abschap6.pdf.

Fatome:2007:AFH

[FFPM07] J. Fatome, C. Finot, S. Pitois, and G. Millot. All-fibered high-quality low duty-cycle picosecond high repetition rate pulse sources. Annales de Physique (Paris, France), 32(2–3):67–??, ???? 2007. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
2007/02/ann02200709/ann02200709.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200709.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200709/
ann02200709.html.

Finot:2007:SCO

[FGD07] C. Finot, A. Guenot, and P. Dupriez. Spectral compression of optical parabolic similaritons. Annales de Physique (Paris, France), 32(2–3):71–??, ???? 2007. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/2007/02/ann02200710/ann02200710.html;
 http://www.annphys.org/articles/anphys/pdf/2007/02/ann02200710.
pdf;   http://www.annphys.org/articles/anphys/ref/2007/
02/ann02200710/ann02200710.html.

Fauchais:1989:PPP

[FGVV89] P. Fauchais, A. Grimaud, A. Vardelle, and M. Vardelle. La projection par plasma : une revue. (French) []. Annales de Physique (Paris, France), 14(3):261–??, ???? 1989. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1989/03/anphys_1989__14_3_261_0/anphys_1989__14_3_261_
0.html;   http://www.annphys.org/articles/anphys/pdf/1989/
03/anphys_1989__14_3_261_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1989/03/anphys_1989__14_3_261_0/anphys_
1989__14_3_261_0.html.

Fuchs:1992:RVD

[FHF92] C. Fuchs, R. Henck, and E. Fogarassy. Réacteur VUV à décharge interne pour le dépôt en couches minces de silicium amorphe, d’oxyde et de nitrure de silicium. (French) []. Annales de Physique (Paris, France), ??(??):59–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921711/
19921711.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921711.pdf.

Fleury:2009:EIC

[Fle09] G. Fleury. Effet de l’interaction coulombienne sur la localisation d’Anderson dans le gaz bidimensionnel d’électrons. (French) [Effect of Coulomb interaction on Anderson localization in the two-dimensional electron gas]. Annales de Physique (Paris, France), 34(2–3):1–??, ???? 2009. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2009/
03/anphys20103445/anphys20103445.html;   http://www.annphys.
org/articles/anphys/pdf/2009/03/anphys20103445.pdf;   http:

//www.annphys.org/articles/anphys/ref/2009/03/anphys20103445/
anphys20103445.html.

Flicstein:1997:ABT

[FMOS97] J. Flicstein, J. Mba, P. Ossart, and J. M. Le Solliec. Apport bénéfique des traitements par photons pour la réalisation de composants III–V de bonne qualité. (French) []. Annales de Physique (Paris, France), 22(C1):C1–269, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972251/
19972251.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972251.pdf.

Fettouhi:1995:EDB

[FMSZ95] N. E. Fettouhi, B. Macke, B. Segard, and J. Zemmouri. Effets dynamiques en bistabilité optique : loi d’échelle. (French) []. Annales de Physique (Paris, France), 20(5–6):615–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05381995/ann05381995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05381995.pdf.

Fargin:1995:COM

[FOF95] E. Fargin, R. Olazcuaga, and G. Le Flem. Caractérisation et optimisation de matériaux non-linéaires. Application à la technologie des télécommunications. (French) []. Annales de Physique (Paris, France), 20(5–6):617–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05391995/ann05391995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05391995.pdf.

Fourme:1992:SRX

[Fou92] R. Fourme. Les sources de rayons X pour la bio-cristallographie. (French) []. Annales de Physique (Paris, France), ??(??):247–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1992/07/19921746/19921746.html;   http://www.annphys.
org/articles/anphys/pdf/1992/07/19921746.pdf.

Frascaria:1991:CEV

[Fra91] N. Frascaria. De ces états de vibrations que l’on appelle résonances géantes et multiphonons. (French) []. Annales de Physique (Paris, France), 16(1):1–??, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1991/
01/anphys_1991__16_1_1_0/anphys_1991__16_1_1_0.html;
http://www.annphys.org/articles/anphys/pdf/1991/01/anphys_
1991__16_1_1_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1991/01/anphys_1991__16_1_1_0/anphys_1991__
16_1_1_0.html.

Freed:1985:SWS

[Fre85] J. H. Freed. Spin waves in spin-polarized H. Annales de Physique (Paris, France), 10(6):901–??, ???? 1985. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/06/anphys_1985__10_6_901_0/anphys_1985__10_6_901_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
06/anphys_1985__10_6_901_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1985/06/anphys_1985__10_6_901_0/anphys_
1985__10_6_901_0.html.

Frehaut:2000:PCS

[Fré00] J. Fréhaut. Physique des cibles de spallation: les codes et leur validation. (French) [Physics of spallation targets: codes and their validation]. Annales de Physique (Paris, France), 25 (2):115–??, ???? 2000. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/2000/02/chap4/chap4.html;   http:/
/www.annphys.org/articles/anphys/pdf/2000/02/chap4.pdf.

Friedel:1985:ABF

[Fri85] J. Friedel. André Blandin 1933–1983. Annales de Physique (Paris, France), 10(1):1–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1985/01/anphys_1985__10_1_1_0/
anphys_1985__10_1_1_0.html;   http://www.annphys.org/articles/
anphys/pdf/1985/01/anphys_1985__10_1_1_0.pdf.

Friedjung:1991:SDS

[Fri91] Michael Friedjung. Stages of development of symbiotic stars. Annales de Physique (Paris, France), 16(4):423–??, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1991/04/anphys_1991__16_4_423_0/anphys_1991__16_4_423_

0.html;   http://www.annphys.org/articles/anphys/pdf/1991/
04/anphys_1991__16_4_423_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1991/04/anphys_1991__16_4_423_0/anphys_
1991__16_4_423_0.html.

Friedel:1992:NNI

[Fri92] Paul Friedel. Neural networks in industry; a case study. Annales de Physique (Paris, France), 17(6):553–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1992/06/anphys_1992__17_6_553_0/anphys_1992__17_6_553_
0.html;   http://www.annphys.org/articles/anphys/pdf/1992/
06/anphys_1992__17_6_553_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1992/06/anphys_1992__17_6_553_0/anphys_
1992__17_6_553_0.html.

Fontaine:1992:NSP

[FS92] B. Fontaine and M. Sentis. Nouvelles sources photoniques VUV à pompage collisionnel. (French) []. Annales de Physique (Paris, France), ??(??):219–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921742/
19921742.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921742.pdf.

Fichet:1995:IVW

[FSBD95] M. Fichet, F. Schuller, D. Bloch, and M. Ducloy. Interaction de van der Waals résonnante entre un atome dans un état excité et une surface diélectrique : spectres de réflexion sélective. (French) []. Annales de Physique (Paris, France), 20(5–6):649–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1995/05/ann05551995/ann05551995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05551995.pdf.

Faklaris:2007:DNP

[FSC^{+}07] O. Faklaris, Y. Sonnefraud, A. Cuche, T. Sauvage, V. Joshi, J. P. Boudou, P. A. Curmi, J. F. Roch, S. Huant, and F. Treussart. Diamond nanoparticles as photoluminescent nanoprobes. Annales de Physique (Paris, France), 32(2–3):155–??, ???? 2007. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2007/
02/ann02200731/ann02200731.html;   http://www.annphys.

org/articles/anphys/pdf/2007/02/ann02200731.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200731/
ann02200731.html.

Faurie:1995:LBE

[FTGB95] J.-P. Faurie, E. Tournié, P. Gibart, and B. Beaumont. Le laser bleu : état de l’art. (French) []. Annales de Physique (Paris, France), 20(5–6):743–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1995/05/ann05691995/ann05691995.html;
 http://www.annphys.org/articles/anphys/pdf/1995/05/ann05691995.
pdf.

Gauthier:1997:PRP

[GABG97] J. C. Gauthier, P. Audebert, S. Bastiani, and J. P. Geindre. Progrès récents dans la production et l’utilisation de rayonnement X au-dessus du keV par plasma femtoseconde. (French) []. Annales de Physique (Paris, France), 22(C1):C1–61, ???? 1997. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1997/
07/19972209/19972209.html;   http://www.annphys.org/articles/
anphys/pdf/1997/07/19972209.pdf.

Gaetan:2009:IDA

[Gaë09] A. Gaëtan. Intrication de deux atomes en utilisant le blocage de Rydberg. (French) [Entanglement of two atoms using Rydberg locking]. Annales de Physique (Paris, France), 34(6):1–??, ???? 2009. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/2009/04/ann062009/ann062009.html;   http://
www.annphys.org/articles/anphys/pdf/2009/04/ann062009.
pdf;   http://www.annphys.org/articles/anphys/ref/2009/
04/ann062009/ann062009.html.

Galle:1985:MIA

[Gal85] P. Galle. La microscopie ionique analytique des tissus biologiques. (French) []. Annales de Physique (Paris, France), 10(4):287–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1985/04/anphys_1985__10_4_287_0/anphys_1985__10_4_
287_0.html;   http://www.annphys.org/articles/anphys/pdf/
1985/04/anphys_1985__10_4_287_0.pdf;   http://www.annphys.
org/articles/anphys/ref/1985/04/anphys_1985__10_4_287_
0/anphys_1985__10_4_287_0.html.

Gall:1999:SHD

[Gal99] B. J. P. Gall. Super et hyper déformations du noyau atomique: du discret au continu. (French) [super and hyper deformation of the atomic nucleus: from the discreet to the continuous]. Annales de Physique (Paris, France), 24(6):1–??, ???? 1999. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/1999/06/ann061999/ann061999.html;   http://
www.annphys.org/articles/anphys/pdf/1999/06/ann061999.
pdf.

Geindre:1994:PIU

[GARG94] J. P. Geindre, P. Audebert, A. Rousse, and J. C. Gauthier. Production d’impulsions ultracourtes de rayons X par interaction laser matière. (French) []. Annales de Physique (Paris, France), 19(C1):C1–103, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941925/
19941925.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941925.pdf.

Gauthier:1992:DIE

[Gau92] J. C. Gauthier. Dynamique d’ionisation et émission X des plasmas créés par laser. (French) []. Annales de Physique (Paris, France), ??(??):83–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921720/
19921720.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921720.pdf.

Gayral:2001:CSE

[Gay01] B. Gayral. Controlling spontaneous emission dynamics in semiconductor microcavities. Annales de Physique (Paris, France), 26(2):1–??, ???? 2001. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.

annphys.org/articles/anphys/abs/2001/02/annales022001/
annales022001.html;   http://www.annphys.org/articles/
anphys/pdf/2001/02/annales022001.pdf.

Guy:1995:LVN

[GB95] O. Guy and A. Barthélémy. Laser verre : Nd picoseconde pompé par diode. (French) []. Annales de Physique (Paris, France), 20 (5–6):575–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1995/05/ann05181995/ann05181995.html;
 http://www.annphys.org/articles/anphys/pdf/1995/05/ann05181995.
pdf.

Gilliot:2007:CCQ

[GBC^{+}07] P. Gilliot, C. Brimont, S. Cronenberger, O. Crégut, M. Gallart, B. Hönerlage, K. Kheng, and H. Mariette. Coherent control and quantum kinetics of LO-phonon emission in CdZnTe quantum wells. Annales de Physique (Paris, France), 32(2–3):119–??, ???? 2007. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
2007/02/ann02200722/ann02200722.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200722.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200722/
ann02200722.html.

Giammanco:1997:HGA

[GCT^{+}97] F. Giammanco, P. Ceccherini, C. Tagliavini, M. Malvezzi, P. Villoresi, and G. Tondello. Harmonic generation in N_{2} and Ar by wave-mixing. Annales de Physique (Paris, France), 22(C1):C1–69, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1997/07/19972210/19972210.html;   http://www.annphys.
org/articles/anphys/pdf/1997/07/19972210.pdf.

Garet:1995:SDT

[GDA^{+}95] F. Garet, L. Duvillaret, G. Angénieux, J.-L. Coutaz, and J. F. Whitaker. Spectroscopie dans le domaine térahertz à l’aide d’impulsions optiques ultracourtes. (French) []. Annales de Physique (Paris, France), 20(5–6):565–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05131995/ann05131995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05131995.pdf.

Geneux:1985:RMA

[Gen85] E. Geneux. Résonance magnétique et aimantation transversale. (French) [magnetic resonance and magnetization transversal]. Annales de Physique (Paris, France), 10(6):1053–??, ???? 1985. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1985/
06/anphys_1985__10_6_1053_0/anphys_1985__10_6_1053_0.html;
 http://www.annphys.org/articles/anphys/pdf/1985/06/anphys_
1985__10_6_1053_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1985/06/anphys_1985__10_6_1053_0/anphys_1985_
_10_6_1053_0.html.

Genova:1987:ERO

[Gen87] F. Genova. Les émissions radio <<aurorales >>des planètes. (French) []. Annales de Physique (Paris, France), 12(2):57–??, ???? 1987. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1987/02/anphys_1987__12_2_57_0/anphys_1987__12_2_57_0.
html;   http://www.annphys.org/articles/anphys/pdf/1987/
02/anphys_1987__12_2_57_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1987/02/anphys_1987__12_2_57_0/anphys_
1987__12_2_57_0.html.

Gerbier:2004:CBE

[Ger04] F. Gerbier. Condensat de Bose–Einstein dans un piège anisotrope. Cohérence en phase et propriétés thermodynamiques. (French) [Bose–Einstein condensate in an anisotropic trap. Phase coherence and thermodynamic properties]. Annales de Physique (Paris, France), 29(1):1–??, ???? 2004. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2004/
01/ann012004/ann012004.html;   http://www.annphys.org/
articles/anphys/pdf/2004/01/ann012004.pdf;   http://www.
annphys.org/articles/anphys/ref/2004/01/ann012004/ann012004.
html.

Giacobino:1995:RDB

[GF95] E. Giacobino and C. Fabre. Réduction du bruit quantique dans l’oscillateur paramétrique et dans les lasers. (French) []. Annales de Physique (Paris, France), 20(5–6):509–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05051995/ann05051995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05051995.pdf.

Giovannella:1988:LFS

[GFC88] C. Giovannella, L. Fruchter, and C. Chappert. Low field scaling properties of high T_{c} superconductor glasses. Annales de Physique (Paris, France), 13(5):393–??, ???? 1988.

CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1988/05/anphys_1988__13_5_393_0/anphys_1988__13_5_393_
0.html;   http://www.annphys.org/articles/anphys/pdf/1988/
05/anphys_1988__13_5_393_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1988/05/anphys_1988__13_5_393_0/anphys_
1988__13_5_393_0.html.

Garnero:1999:PIT

[GFCB99] L. Garnero, J. Fonroget, F. Coutand, and P. Bailly. Problèmes inverses en tomographie d’émission — nouvelles tendances. (French) [Inverse problems in emission tomography — new trends]. Annales de Physique (Paris, France), 24(3):155–??, ???? 1999. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1999/
03/abschap9/abschap9.html;   http://www.annphys.org/articles/
anphys/pdf/1999/03/abschap9.pdf.

Gales:1987:SIS

[GG87] S. Galès and Nguyen Van Giai. Symétrie d’isospin et structure nucléaire. (French) []. Annales de Physique (Paris, France), 12(3):183–??, ???? 1987. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1987/03/anphys_1987__12_3_183_
0/anphys_1987__12_3_183_0.html;   http://www.annphys.org/
articles/anphys/pdf/1987/03/anphys_1987__12_3_183_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1987/03/anphys_
1987__12_3_183_0/anphys_1987__12_3_183_0.html.

Germain:1994:EAC

[GGA^{+}94] C. Germain, C. Girault, J. Aubreton, A. Catherinot, S. Bec, and A. Tonck. Étude de l’ablation d’une cible graphite par un laser U.V., application à la réalisation de couches de carbone dur. (French) []. Annales de Physique (Paris, France), 19(C1):C1–183, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941946/
19941946.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941946.pdf.

Garrelie:1997:ALF

[GGA^{+}97] F. Garrelie, C. Girault, B. Angleraud, C. Champeaux, and A. Catherinot. Ablation laser en fond de cratère. (French) []. Annales de Physique (Paris, France), 22(C1):C1–123, ???? 1997. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1997/
07/19972223/19972223.html;   http://www.annphys.org/articles/
anphys/pdf/1997/07/19972223.pdf.

Garrelie:1997:EAC

[GGCC97] F. Garrelie, C. Girault, C. Champeaux, and A. Catherinot. Étude de l’ablation d’une cible de cuivre par un laser à excimères KrF. (French) []. Annales de Physique (Paris, France), 22(C1):C1–125, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972224/
19972224.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972224.pdf.

Grob:1997:SDC

[GGM^{+}97] A. Grob, J. J. Grob, D. Muller, B. Prévot, and E. Fogarassy. Solubilité du carbone et relaxation des contraintes dans des films de Si_{1−x−y}Ge_{x}C_{y} préparés par implantation ionique et recristallisation par laser à excimères. (French) []. Annales de Physique (Paris, France), 22(C1):C1–227, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972244/
19972244.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972244.pdf.

Geindre:1984:XRS

[GGP84] J.-P. Geindre, J.-C. Gauthier, and C. Popovics. X-ray spectroscopy of dense plasmas produced by laser irradiation at the GRECO ILM. Annales de Physique (Paris, France), 9(4):665–??, ???? 1984. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1984/04/anphys_1984__9_4_665_0/anphys_1984__9_4_665_0.
html;   http://www.annphys.org/articles/anphys/pdf/1984/
04/anphys_1984__9_4_665_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1984/04/anphys_1984__9_4_665_0/anphys_
1984__9_4_665_0.html.

Grangier:1995:MQN

[GGPR95] Ph. Grangier, Ph. Grelu, J.-Ph. Poizat, and J.-F. Roch. Les mesures quantiques non-destructives en optique. Annales de Physique (Paris, France), 20(5–6):675–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05601995/ann05601995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05601995.pdf.

Gallot:1995:CEN

[GHZ^{+}95] G. Gallot, F. Hache, A. Zéboulon, T. J. Driscoll, and G. M. Gale. Cascades d’effets non linéaires du seconde-ordre en régime femtoseconde dans le BBO : autocompression dans un oscillateur paramétrique optique. (French) []. Annales de Physique (Paris, France), 20(5–6):639–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05501995/ann05501995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05501995.pdf.

Giannoni:1984:MCG

[Gia84] M.-J. Giannoni. Mouvements collectifs de grandes amplitudes dans les noyaux : une approche microscopiquex. (French) []. Annales de Physique (Paris, France), 9(1):39–??, ???? 1984. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1984/
01/anphys_1984__9_1_39_0/anphys_1984__9_1_39_0.html;
http://www.annphys.org/articles/anphys/pdf/1984/01/anphys_
1984__9_1_39_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1984/01/anphys_1984__9_1_39_0/anphys_1984__
9_1_39_0.html.

Gilabert:1990:IBL

[Gil90] Alain Gilabert. Interaction between light and superconductors. Annales de Physique (Paris, France), 15(3):255–??, ???? 1990. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1990/03/anphys_1990__15_3_255_0/anphys_1990__15_3_255_
0.html;   http://www.annphys.org/articles/anphys/pdf/1990/
03/anphys_1990__15_3_255_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1990/03/anphys_1990__15_3_255_0/anphys_
1990__15_3_255_0.html.

Giovannini:1995:MCP

[Gio95] H. Giovannini. Le multiplexage de cohérence : principes et applications. (French) []. Annales de Physique (Paris, France), 20 (5–6):655–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1995/05/ann05581995/ann05581995.html;
 http://www.annphys.org/articles/anphys/pdf/1995/05/ann05581995.
pdf.

Guy:1995:CRB

[GJJR95] S. Guy, M. F. Joubert, B. Jacquier, and S. C. Rand. Conversion rouge/bleu par avalanche de photons dans un cristal de YAlO_{3} dopé thulium. (French) []. Annales de Physique (Paris, France), 20(5–6):577–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05191995/ann05191995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05191995.pdf.

Godard:1994:LEK

[GMS^{+}94] B. Godard, P. Murer, M. Stehle, P. Laborde, J. Bonnet, and D. Pigache. Laser excimère 1 kW (10J × 100 Hz)

— applications. (French) []. Annales de Physique (Paris, France), 19(C1):C1–203, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941950/
19941950.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941950.pdf.

Gardou:1997:EER

[GMS^{+}97] J. P. Gardou, F. Marchal, N. Sewraj, S. Tusseau, P. Millet, and Y. Salamero. Étude de l’emprisonnement des raies de résonance dans les mélanges de gaz rares. (French) []. Annales de Physique (Paris, France), 22(C1):C1–109, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972220/
19972220.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972220.pdf.

Gorris-Neveux:1995:SRS

[GNMS^{+}95] M. Gorris-Neveux, P. Monnot, S. Saltiel, M. Ducloy, R. Barbé, and J.-C. Keller. Spectroscopie de réflexion sélective à deux photons. (French) []. Annales de Physique (Paris, France), 20(5–6):579–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/

abs/1995/05/ann05201995/ann05201995.html;   http://www.
annphys.org/articles/anphys/pdf/1995/05/ann05201995.pdf.

Gogolin:1994:STT

[Gog94] A. O. Gogolin. Selected topics in the theory of 1D quantum wires. Annales de Physique (Paris, France), 19(4):411–??, ???? 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1994/04/anphys_1994__19_4_411_0/anphys_1994__19_4_411_
0.html;   http://www.annphys.org/articles/anphys/pdf/1994/
04/anphys_1994__19_4_411_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1994/04/anphys_1994__19_4_411_0/anphys_
1994__19_4_411_0.html.

Gourgoulhon:1993:EEG

[Gou93] Eric Gourgoulhon. Equilibre et effondrement gravitationnel des étoiles en relativité générale et en symétrie sphérique. (French) []. Annales de Physique (Paris, France), 18(1):1–??, ???? 1993. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1993/
01/anphys_1993__18_1_1_0/anphys_1993__18_1_1_0.html;
http://www.annphys.org/articles/anphys/pdf/1993/01/anphys_
1993__18_1_1_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1993/01/anphys_1993__18_1_1_0/anphys_1993__
18_1_1_0.html.

Girard:1995:AAS

[GPL95] G. Girard, P. Pringent, and A. Lavenant. Accord et affinement spectrale de l’émission optique de lasers/amplificateurs à semiconducteurs à large ouverture. (French) []. Annales de Physique (Paris, France), 20(5–6):635–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05481995/ann05481995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05481995.pdf.

Galeano:1994:EEC

[GR94] Javier Galeano and Miguel A. Rubio. Experiments on earthquakes in a continuous elastic medium. Annales de Physique (Paris, France), 19(6):735–??, ???? 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1994/06/anphys_1994__19_6_735_
0/anphys_1994__19_6_735_0.html;   http://www.annphys.org/

articles/anphys/pdf/1994/06/anphys_1994__19_6_735_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1994/06/anphys_
1994__19_6_735_0/anphys_1994__19_6_735_0.html.

Grandati:1992:EII

[Gra92] Y. Grandati. Élements d’introduction à l’invariance conforme. (French) []. Annales de Physique (Paris, France), 17(3):159–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1992/03/anphys_1992__17_3_159_0/anphys_1992__17_3_
159_0.html;   http://www.annphys.org/articles/anphys/pdf/
1992/03/anphys_1992__17_3_159_0.pdf;   http://www.annphys.
org/articles/anphys/ref/1992/03/anphys_1992__17_3_159_
0/anphys_1992__17_3_159_0.html.

Gauthier:1994:SXS

[GRA^{+}94] J. C. Gauthier, A. Rousse, P. Audebert, J. P. Geindre, F. Falliès, A. Mysyrowicz, G. Grillon, and A. Antonetti. Source X subpicoseconde par un plasma laser femtoseconde. (French) []. Annales de Physique (Paris, France), 19(C1):C1–53, October 1994. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1994/
07/19941915/19941915.html;   http://www.annphys.org/articles/
anphys/pdf/1994/07/19941915.pdf.

Grynberg:1985:RIP

[Gry85] G. Grynberg. Les résonances induites par collision en optique non linéaire : une illustration des concepts du pompage optique. (French) []. Annales de Physique (Paris, France), 10(6):695–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1985/06/anphys_1985__10_6_695_0/anphys_1985__10_6_
695_0.html;   http://www.annphys.org/articles/anphys/pdf/
1985/06/anphys_1985__10_6_695_0.pdf;   http://www.annphys.
org/articles/anphys/ref/1985/06/anphys_1985__10_6_695_
0/anphys_1985__10_6_695_0.html.

Gounand:1984:BOL

[GSH^{+}84] F. Gounand, J. Szudy, M. Hugon, B. Sayer, and P. R. Fournier. Broadening of optical lines originating from Rydberg states: a simple model. (French) []. Annales de Physique (Paris, France), 9(4):597–??, ???? 1984. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1984/
04/anphys_1984__9_4_597_0/anphys_1984__9_4_597_0.html;
 http://www.annphys.org/articles/anphys/pdf/1984/04/anphys_
1984__9_4_597_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1984/04/anphys_1984__9_4_597_0/anphys_1984_
_9_4_597_0.html.

Gregoire:1986:IEH

[GT86] C. Grégoire and B. Tamain. Intermediate energy heavy ion reactions. Annales de Physique (Paris, France), 11(4):323–??, ???? 1986. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1986/04/anphys_1986__11_4_323_0/anphys_1986__11_4_323_
0.html;   http://www.annphys.org/articles/anphys/pdf/1986/
04/anphys_1986__11_4_323_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1986/04/anphys_1986__11_4_323_0/anphys_
1986__11_4_323_0.html.

Gulminelli:2004:PCN

[Gul04] F. Gulminelli. Phase coexistence in nuclei. Annales de Physique (Paris, France), 29(6):1–??, ???? 2004. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2004/
06/ann062004/ann062004.html;   http://www.annphys.org/
articles/anphys/pdf/2004/06/ann062004.pdf;   http://www.
annphys.org/articles/anphys/ref/2004/06/ann062004/ann062004.
html.

Greenberg:1989:CEI

[GZH89] J. Mayo Greenberg, Nansheng Zhao, and Joniek Hage. Chemical evolution of interstellar dust, comets and the origins of life. Annales de Physique (Paris, France), 14(2):103–??, ???? 1989. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1989/02/anphys_1989__14_2_103_0/anphys_1989__14_2_103_
0.html;   http://www.annphys.org/articles/anphys/pdf/1989/
02/anphys_1989__14_2_103_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1989/02/anphys_1989__14_2_103_0/anphys_
1989__14_2_103_0.html.

Hache:1995:OPO

[Hac95a] F. Hache. Oscillateurs paramétriques optiques. Généralités — régimes picoseconde et femtoseconde. (French) []. Annales de Physique (Paris, France), 20(5–6):493–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05031995/ann05031995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05031995.pdf.

Henkel:1995:PCR

[HAC^{+}95b] C. Henkel, A. Aspect, J.-Y. Courtois, R. Kaiser, K. Molmer, and C. I. Westbrook. Perte de cohérence dans la réflexion d’atomes sur une onde évanscente. (French) []. Annales de Physique (Paris, France), 20(5–6):643–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05521995/ann05521995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05521995.pdf.

Heidmann:2007:QOM

[HAC^{+}07] A. Heidmann, O. Arcizet, T. Caniard, C. Molinelli, P. Verlot, T. Briant, and P.-F. Cohadon. Quantum optics with micromirrors. Annales de Physique (Paris, France), 32(2–3):33–??, ???? 2007. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
2007/02/ann02200703/ann02200703.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200703.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200703/
ann02200703.html.

Hameury:1985:SG

[Ham85] J. M. Hameury. Les sursauts gamma. Annales de Physique (Paris, France), 10(5):369–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1985/05/anphys_1985__10_5_369_
0/anphys_1985__10_5_369_0.html;   http://www.annphys.org/
articles/anphys/pdf/1985/05/anphys_1985__10_5_369_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1985/05/anphys_
1985__10_5_369_0/anphys_1985__10_5_369_0.html.

Happer:1985:SEP

[Hap85] W. Happer. Spin exchange, past, present and future. Annales de Physique (Paris, France), 10(6):645–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/06/anphys_1985__10_6_645_0/anphys_1985__10_6_645_

0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
06/anphys_1985__10_6_645_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1985/06/anphys_1985__10_6_645_0/anphys_
1985__10_6_645_0.html.

Haroche:1985:EAR

[Har85] S. Haroche. Electrodynamique des atomes de Rydberg dans une cavité résonnante. (French) []. Annales de Physique (Paris, France), 10(6):811–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1985/06/anphys_1985__10_6_811_
0/anphys_1985__10_6_811_0.html;   http://www.annphys.org/
articles/anphys/pdf/1985/06/anphys_1985__10_6_811_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1985/06/anphys_
1985__10_6_811_0/anphys_1985__10_6_811_0.html.

Heidmann:1989:DAR

[Hei89] Jean Heidmann. Dix années de recherches futures de signaux extraterrestres. (French) []. Annales de Physique (Paris, France), 14(2):133–??, ???? 1989. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1989/02/anphys_1989__14_2_133_
0/anphys_1989__14_2_133_0.html;   http://www.annphys.org/
articles/anphys/pdf/1989/02/anphys_1989__14_2_133_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1989/02/anphys_
1989__14_2_133_0/anphys_1989__14_2_133_0.html.

Hennequin:1994:MOF

[Hen94] Daniel Hennequin. Morphogénèse optique. (French) []. Annales de Physique (Paris, France), 19(6):743–??, ???? 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1994/06/anphys_1994__19_6_743_0/anphys_1994__19_6_743_
0.html;   http://www.annphys.org/articles/anphys/pdf/1994/
06/anphys_1994__19_6_743_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1994/06/anphys_1994__19_6_743_0/anphys_
1994__19_6_743_0.html.

Henck:1997:ASP

[HFF97] R. Henck, C. Fuchs, and E. Fogarassy. Analyse spectrale de plasmas créés par ablation par laser à excimères de cibles Ge, Si et SiGe. (French) []. Annales de Physique (Paris, France), 22(C1):C1–127, ???? 1997. CODEN ANPHAJ.

ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972225/
19972225.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972225.pdf.

Haghiri-Gosnet:1994:LPR

[HG94] A. M. Haghiri-Gosnet. La lithographie par rayons X à très haute résolution — ses limites ultimes. (French) []. Annales de Physique (Paris, France), 19(C1):C1–65, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941921/
19941921.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941921.pdf.

Hourani:1989:RLF

[HHP89] E. Hourani, M. Hussonnois, and D. N. Poenaru. Radioactivities by light fragment (C, Ne, Mg) emission. Annales de Physique (Paris, France), 14(3):311–??, ???? 1989. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1989/03/anphys_1989__14_3_311_0/anphys_1989__14_3_311_
0.html;   http://www.annphys.org/articles/anphys/pdf/1989/
03/anphys_1989__14_3_311_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1989/03/anphys_1989__14_3_311_0/anphys_
1989__14_3_311_0.html.

Hraiech:2007:BYP

[HJK^{+}07] S. Hraiech, A. Jouini, K. Jin Kim, Y. Guyot, C. Goutaudier, A. Yoshikawa, M. Trabelsi-Ayadi, and G. Boulon. Breakage of Yb^{3+} pairs by Na^{+} in Yb^{3+}-doped CaF_{2} laser host. Annales de Physique (Paris, France), 32(2–3):59–??, ???? 2007. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2007/
02/ann02200707/ann02200707.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200707.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200707/
ann02200707.html.

Hermann:1985:OPS

[HLS85] C. Hermann, G. Lampel, and V. I. Safarov. Optical pumping in semiconductors. Annales de Physique (Paris, France), 10(6):1117– ??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/

anphys/abs/1985/06/anphys_1985__10_6_1117_0/anphys_1985_
_10_6_1117_0.html;   http://www.annphys.org/articles/anphys/
pdf/1985/06/anphys_1985__10_6_1117_0.pdf;   http://www.
annphys.org/articles/anphys/ref/1985/06/anphys_1985__10_
6_1117_0/anphys_1985__10_6_1117_0.html.

Hilscher:1992:DNF

[HR92] D. Hilscher and H. Rossner. Dynamics of nuclear fission. Annales de Physique (Paris, France), 17(6):471–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1992/06/anphys_1992__17_6_471_0/anphys_1992__17_6_471_
0.html;   http://www.annphys.org/articles/anphys/pdf/1992/
06/anphys_1992__17_6_471_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1992/06/anphys_1992__17_6_471_0/anphys_
1992__17_6_471_0.html.

Hemici:1994:TFF

[HSD^{+}94] M. Hemici, R. Saoudi, E. Descroix, E. Audouard, and P. Laporte. Techniques de frustration de fluorescence dans le VUV : Effet du couplage spin-orbite dans l’excimère de krypton. (French) []. Annales de Physique (Paris, France), 19(C1):C1–55, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1994/07/19941916/19941916.html;   http://www.annphys.org/
articles/anphys/pdf/1994/07/19941916.pdf.

Henry:1997:PML

[HST97] D. Henry, P. Schiavonne, and O. Toublan. 0,1μ m en production de masse en lithographie optique ?. (French) []. Annales de Physique (Paris, France), 22(C1):C1–185, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972239/
19972239.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972239.pdf.

Hermann:1994:ARC

[HTBL94] J. Hermann, A. L. Thomann, and C. Boulmer-Leborgne. Ablation réactive d’une cible de titane par un laser XeCl en présence d’azote. (French) []. Annales de Physique (Paris, France), 19(C1):C1–57, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941917/

19941917.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941917.pdf.

Huard:2006:IBE

[Hua06] B. Huard. Interactions between electrons, mesoscopic Josephson effect and asymmetric current fluctuations. Annales de Physique (Paris, France), 31(4–5):1–??, ???? 2006. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2006/
04/ann0452006/ann0452006.html;   http://www.annphys.org/
articles/anphys/pdf/2006/04/ann0452006.pdf;   http://www.
annphys.org/articles/anphys/ref/2006/04/ann0452006/ann0452006.
html.

Hughes:1985:ME

[Hug85] V. W. Hughes. The muon and the electron. Annales de Physique (Paris, France), 10(6):955–??, ???? 1985. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/06/anphys_1985__10_6_955_0/anphys_1985__10_6_955_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
06/anphys_1985__10_6_955_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1985/06/anphys_1985__10_6_955_0/anphys_
1985__10_6_955_0.html.

Hulin:1995:DCE

[Hul95] D. Hulin. Du contrôle des électrons à celui des photons. (French) []. Annales de Physique (Paris, France), 20(5–6):535–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1995/05/ann05081995/ann05081995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05081995.pdf.

Idier:1994:MIN

[IFRS94] D. Idier, M. Farine, B. Remaud, and F. Sébille. Modelling an infinite nucleonic system. Static and dynamical properties. Study of density fluctuations. Annales de Physique (Paris, France), 19(2):159–??, ???? 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1994/02/anphys_1994__19_2_159_
0/anphys_1994__19_2_159_0.html;   http://www.annphys.org/
articles/anphys/pdf/1994/02/anphys_1994__19_2_159_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1994/02/anphys_
1994__19_2_159_0/anphys_1994__19_2_159_0.html.

Idir:1997:OBF

[IMCD97] M. Idir, A. Mirone, G. Cauchon, and P. Dhez. Optiques de Bragg– Fresnel. Principes et applications. (French) []. Annales de Physique (Paris, France), 22(C1):C1–139, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972230/
19972230.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972230.pdf.

Ikhlef:1994:IXI

[IS94] A. Ikhlef and M. Skowronek. Imagerie X instantanée (ns) utilisant une source X de faible énergie (mJ), des fibres scintillantes et une caméra intensifiée. (French) []. Annales de Physique (Paris, France), 19(C1):C1–61, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941919/
19941919.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941919.pdf.

Jacques:2007:SPU

[Jac07] V. Jacques. Sources de photons uniques et expérience à choix retardé de Wheeler: la dualité onde corpuscule à l’épreuve de l’expérience. (French) [single photon sources and Wheeler’s experiment of delayed choice: wave–corpuscle duality in a test experiment]. Annales de Physique (Paris, France), 32(6):1–??, ???? 2007. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/2007/04/ann062007/ann062007.html;   http://
www.annphys.org/articles/anphys/pdf/2007/04/ann062007.
pdf;   http://www.annphys.org/articles/anphys/ref/2007/
04/ann062007/ann062007.html.

Jacquey:2007:MRI

[JBTV07] M. Jacquey, M. Büchner, G. Trénec, and J. Vigué. Measurement of the refraction index for atom waves. Annales de Physique (Paris, France), 32(2–3):183–??, ???? 2007. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2007/
02/ann02200738/ann02200738.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200738.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200738/
ann02200738.html.

Jacquot:2007:ODD

[JCL07] M. Jacquot, Y. Chembo, and L. Larger. Optoelectronic delay dynamics: from optical chaos communication to high purity microwave oscillators. Annales de Physique (Paris, France), 32 (2–3):39–??, ???? 2007. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/2007/02/ann02200704/ann02200704.html;
 http://www.annphys.org/articles/anphys/pdf/2007/02/ann02200704.
pdf;   http://www.annphys.org/articles/anphys/ref/2007/
02/ann02200704/ann02200704.html.

Jurczak:1995:EPC

[JDS^{+}95] C. Jurczak, B. Desruelle, K. Sengstock, J. Y. Courtois, R. Kaiser, N. Vansteenkiste, C. I. Westbrook, and A. Aspect. Étude par corrélations d’intensité sélectives en polarisation d’un réseau d’atomes froids. (French) []. Annales de Physique (Paris, France), 20(5–6):597–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05291995/ann05291995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05291995.pdf.

Jefferson:1988:ECC

[Jef88] J. H. Jefferson. Electron correlations in copper oxides. Annales de Physique (Paris, France), 13(5):425–??, ???? 1988. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1988/05/anphys_1988__13_5_425_0/anphys_1988__13_5_425_
0.html;   http://www.annphys.org/articles/anphys/pdf/1988/
05/anphys_1988__13_5_425_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1988/05/anphys_1988__13_5_425_0/anphys_
1988__13_5_425_0.html.

Jamelot:1984:PIA

[JJC84] G. Jamelot, P. Jaeglé, and A. Carillon. Population inversions for amplification of spontaneous emission in the soft X-ray range. A step towards X-ray laser. Annales de Physique (Paris, France), 9(4):657–??, ???? 1984. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1984/
04/anphys_1984__9_4_657_0/anphys_1984__9_4_657_0.html;
 http://www.annphys.org/articles/anphys/pdf/1984/04/anphys_
1984__9_4_657_0.pdf;   http://www.annphys.org/articles/

anphys/ref/1984/04/anphys_1984__9_4_657_0/anphys_1984_
_9_4_657_0.html.

Jonin:1995:FEX

[JLS^{+}95] C. Jonin, P. Laporte, R. Saoudi, M. Hemici, and M. Bon. Fluorescence de l’excimère Xe^{2∗} induite par excitation à deux photons : effet des hautes températures. (French) []. Annales de Physique (Paris, France), 20(5–6):621–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05411995/ann05411995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05411995.pdf.

Joly:1993:UPS

[Jol93] Monique Joly. An unsolved problem in the spectrum of quasars: the Fe II emission lines. Annales de Physique (Paris, France), 18(3):241–??, ???? 1993. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1993/03/anphys_1993__18_3_241_
0/anphys_1993__18_3_241_0.html;   http://www.annphys.org/
articles/anphys/pdf/1993/03/anphys_1993__18_3_241_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1993/03/anphys_
1993__18_3_241_0/anphys_1993__18_3_241_0.html.

Joyeux:1992:MRX

[Joy92] D. Joyeux. Microscopie en rayons X-mous : techniques et performances. (French) []. Annales de Physique (Paris, France), ??(??):231–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921744/
19921744.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921744.pdf.

Joyeux:1999:HRS

[JP99] D. Joyeux and F. Polack. High resolution, stigmatic optical reconstruction of X-ray Gabor holograms. Annales de Physique (Paris, France), 24(3):5–??, ???? 1999. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1999/03/abschap1/
abschap1.html;   http://www.annphys.org/articles/anphys/
pdf/1999/03/abschap1.pdf.

Julia:1989:SP

[Jul89] B. Julia. From supernovae to phosphorites. I. Annales de Physique (Paris, France), 14(2):147–??, ???? 1989. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1989/02/anphys_1989__14_2_147_0/anphys_1989__14_2_147_
0.html;   http://www.annphys.org/articles/anphys/pdf/1989/
02/anphys_1989__14_2_147_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1989/02/anphys_1989__14_2_147_0/anphys_
1989__14_2_147_0.html.

Jusserand:1988:SDE

[Jus88] Bernard Jusserand. Superlattice and disorder effects on vibrations in III–V compounds. Annales de Physique (Paris, France), 13(6):597–??, ???? 1988. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1988/06/anphys_1988__13_6_597_
0/anphys_1988__13_6_597_0.html;   http://www.annphys.org/
articles/anphys/pdf/1988/06/anphys_1988__13_6_597_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1988/06/anphys_
1988__13_6_597_0/anphys_1988__13_6_597_0.html.

Jendral:1999:OER

[JWBB99] A. Jendral, S. Wentzel, R. Bräuer, and O. Bryngdahl. Obstruction effect in reconstruction of synthetic near field holograms. Annales de Physique (Paris, France), 24(3):21–??, ???? 1999. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1999/
03/abschap2/abschap2.html;   http://www.annphys.org/articles/
anphys/pdf/1999/03/abschap2.pdf.

Jacques:2007:WDC

[JWG^{+}07] V. Jacques, E. Wu, F. Grosshans, F. Treussart, A. Aspect, Ph. Grangier, and J.-F. Roch. Wheeler’s delayed-choice thought experiment: Experimental realization and theoretical analysis. Annales de Physique (Paris, France), 32(2–3):195–??, ???? 2007. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2007/
02/ann02200741/ann02200741.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200741.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200741/
ann02200741.html.

Kaminskii:1991:AMC

[Kam91] Alexander A. Kaminskii. Achievements of modern crystal-laser physics. Annales de Physique (Paris, France), 16(6):639–??, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1991/06/anphys_1991__16_6_639_0/anphys_1991__16_6_639_
0.html;   http://www.annphys.org/articles/anphys/pdf/1991/
06/anphys_1991__16_6_639_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1991/06/anphys_1991__16_6_639_0/anphys_
1991__16_6_639_0.html.

Kanamori:1985:OS

[Kan85] J. Kanamori. Ordering on surfaces. Annales de Physique (Paris, France), 10(1):43–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1985/01/anphys_1985__10_1_43_0/
anphys_1985__10_1_43_0.html;   http://www.annphys.org/
articles/anphys/pdf/1985/01/anphys_1985__10_1_43_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1985/01/anphys_
1985__10_1_43_0/anphys_1985__10_1_43_0.html.

Kante:2009:MMO

[Kan09] B. Kante. Les métamatériaux, des micro-ondes à l optique: théorie et applications. (French) [Metamaterials, microwaves in optics: theory and applications]. Annales de Physique (Paris, France), 34(2–3):1–??, ???? 2009. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/2009/02/anphys022009/anphys022009.
html;   http://www.annphys.org/articles/anphys/pdf/2009/
02/anphys022009.pdf;   http://www.annphys.org/articles/
anphys/ref/2009/02/anphys022009/anphys022009.html.

Kats:1994:LLC

[Kat94] E. I. Kats. Line liquid crystals. Annales de Physique (Paris, France), 19(4):447–??, ???? 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1994/04/anphys_1994__19_4_447_
0/anphys_1994__19_4_447_0.html;   http://www.annphys.org/
articles/anphys/pdf/1994/04/anphys_1994__19_4_447_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1994/04/anphys_
1994__19_4_447_0/anphys_1994__19_4_447_0.html.

Kobylko:2007:TCR

[KBW^{+}07] M. Kobylko, A.-M. Bonnot, A. R. Hight Walker, A. Richard, M. Kociak, and A. Débarre. Towards correlating Raman excitation profile and electron diffraction of the same single carbon nanotube. Annales de Physique (Paris, France), 32(2–3):131–??, ???? 2007. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
2007/02/ann02200725/ann02200725.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200725.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200725/
ann02200725.html.

Key:1992:PRL

[KCK^{+}92] M. H. Key, H. Z. Chen, R. Kodama, P. Norreys, A. Kidd, J. Zhang, G. J. Tallents, L. Dwivedi, J. Krishnan, S. A. Ramsden, and G. J. Pert. Progrès des recherches sur le laser X : une émission saturée et cohérente à 236 å. (French) []. Annales de Physique (Paris, France), ??(??):239–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921745/
19921745.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921745.pdf.

Kermene:1995:CIO

[KDC^{+}95] V. Kermene, A. Desfarges, B. Colombeau, C. Froehly, and M. Vampouille. Couplage intracavité d’ondes lasers gaussiennes en régimes continu et pulsé. (French) []. Annales de Physique (Paris, France), 20(5–6):631–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05461995/ann05461995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05461995.pdf.

Kieleck:2007:IZO

[KEH07] C. Kieleck, M. Eichhorn, and A. Hirth. Investigation of ZGP OPOs setups at high repetition rates (20–50 kHz). Annales de Physique (Paris, France), 32(2–3):79–??, ???? 2007. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2007/
02/ann02200712/ann02200712.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200712.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200712/
ann02200712.html.

Krastev:1997:DPR

[KLC^{+}97] K. Krastev, F. Leguern, K. Coat, R. Barchewitz, J.-M. André, M.- F. Ravet, E. Cambril, F. Rousseaux, and P. Davi. Diffraction par des réseaux lamellaires multicouches : application à la réalisation d’un polychromateur X-UV. (French) []. Annales de Physique (Paris, France), 22(C1):C1–149, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972232/
19972232.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972232.pdf.

Kleppner:1985:AST

[Kle85] D. Kleppner. Atomic stabilization through electronic and nuclear polarization. Annales de Physique (Paris, France), 10(6):877–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1985/06/anphys_1985__10_6_877_0/anphys_1985__10_6_
877_0.html;   http://www.annphys.org/articles/anphys/pdf/
1985/06/anphys_1985__10_6_877_0.pdf;   http://www.annphys.
org/articles/anphys/ref/1985/06/anphys_1985__10_6_877_
0/anphys_1985__10_6_877_0.html.

Klisnick:1994:SLX

[Kli94] A. Klisnick. Sources laser X de laboratoire et applications. (French) []. Annales de Physique (Paris, France), 19(C1):C1–119, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/1994/07/19941927/19941927.html;   http://www.
annphys.org/articles/anphys/pdf/1994/07/19941927.pdf.

Kohn:1985:LWP

[Koh85] W. Kohn. Long wavelength perturbations and excitations at surfaces. Annales de Physique (Paris, France), 10(1):37–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/01/anphys_1985__10_1_37_0/anphys_1985__10_1_37_0.
html;   http://www.annphys.org/articles/anphys/pdf/1985/
01/anphys_1985__10_1_37_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1985/01/anphys_1985__10_1_37_0/anphys_
1985__10_1_37_0.html.

Kope:1988:CBC

[Kop88] T. K. Kope. Competition between charge ordering and superconductivity in La_{2}-xMxCuO_{4}. Annales de Physique (Paris, France), 13(5):431–??, ???? 1988. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1988/05/anphys_1988__13_5_431_
0/anphys_1988__13_5_431_0.html;   http://www.annphys.org/
articles/anphys/pdf/1988/05/anphys_1988__13_5_431_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1988/05/anphys_
1988__13_5_431_0/anphys_1988__13_5_431_0.html.

Kachanov:1995:PMS

[KRL^{+}95] A. Kachanov, D. Romanini, E. Lacot, F. Stoeckel, and D. Barbier. Perte de la mémoire spectrale dans un laser solide très multimode. (French) []. Annales de Physique (Paris, France), 20(5–6):569–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1995/05/ann05151995/ann05151995.html;   http://www.
annphys.org/articles/anphys/pdf/1995/05/ann05151995.pdf.

Kroto:1989:RLS

[Kro89] H. W. Kroto. The role of linear and spheroïdal carbon molecules in interstellar grain formation. Annales de Physique (Paris, France), 14(2):169–??, ???? 1989. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1989/02/anphys_1989__14_2_169_
0/anphys_1989__14_2_169_0.html;   http://www.annphys.org/
articles/anphys/pdf/1989/02/anphys_1989__14_2_169_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1989/02/anphys_
1989__14_2_169_0/anphys_1989__14_2_169_0.html.

Knoop:1995:RLN

[KRVV95] M. Knoop, J. Rink, M. Vedel, and F. Vedel. Refroidissement laser d’un nuage d’ions Ca^{+} en piège de Paul. (French) []. Annales de Physique (Paris, France), 20(5–6):645–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05531995/ann05531995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05531995.pdf.

Karapanagioti:1994:EDC

[KSF^{+}94] N. Karapanagioti, Y. L. Shao, O. Faucher, D. Charalambidis, H. Bachau, and E. Cormier. Expérience ‘deux couleurs’ dans le magnésium. Excitation et couplage d’états autoionisants par des sources UV. (French) []. Annales de Physique (Paris, France), 19(C1):C1–63, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941920/
19941920.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941920.pdf.

Klisnick:1990:MRX

[KSG^{+}90] A. Klisnick, A. Sureau, H. Guennou, C. Möller, and J. Virmont. Modelling of recombination X-ray lasers. Annales de Physique (Paris, France), 15(6):529–??, ???? 1990. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1990/06/anphys_1990__15_6_529_0/anphys_1990__15_6_529_
0.html;   http://www.annphys.org/articles/anphys/pdf/1990/
06/anphys_1990__15_6_529_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1990/06/anphys_1990__15_6_529_0/anphys_
1990__15_6_529_0.html.

Kunz:1990:MTS

[KTB90] Werner Kunz, Pierre Turq, and Josef Barthel. Modern techniques for the study of electrolyte solutions. Annales de Physique (Paris, France), 15(5):447–??, ???? 1990. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1990/05/anphys_1990__15_5_447_0/anphys_1990__15_5_447_
0.html;   http://www.annphys.org/articles/anphys/pdf/1990/
05/anphys_1990__15_5_447_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1990/05/anphys_1990__15_5_447_0/anphys_
1990__15_5_447_0.html.

Khacef:1997:EPS

[KVC^{+}97] A. Khacef, R. Viladrosa, C. Cachoncinlle, E. Robert, and J. M. Pouvesle. Étude paramétrique d’une source X impulsionnelle créée par décharge rapide THT dans le vide ou dans les gaz basse pression. (French) []. Annales de Physique (Paris, France), 22(C1):C1–71, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972211/
19972211.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972211.pdf.

Kanaev:1992:EEP

[KZAK^{+}92] A. Kanaev, V. Zafiropulos, M. Aït-Kaci, L. Museur, H. Nkwawo, and M. C. Castex. Étude des effets de pression sur la dynamique de relaxation du krypton après excitation sélective par un laser ultraviolet à 125 nm. (French) []. Annales de Physique (Paris, France), ??(??):61–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921712/
19921712.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921712.pdf.

Labilloy:2000:EEC

[Lab00] D. Labilloy. Étude expérimentale de cristaux photoniques bidimensionnels. (French) [Experimental study of two-dimensional photonic crystals]. Annales de Physique (Paris, France), 25(3):1– ??, ???? 2000. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/2000/03/labilloy/labilloy.html;   http://www.
annphys.org/articles/anphys/pdf/2000/03/labilloy.pdf.

Lafon:1991:ETE

[Laf91a] Jean-Pierre J. Lafon. Étapes tardives de l’évolution stellaire — préambule. (French) []. Annales de Physique (Paris, France), 16(4):R2, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1991/04/anphys_1991__16_4_360_
0/anphys_1991__16_4_360_0.html;   http://www.annphys.org/
articles/anphys/pdf/1991/04/anphys_1991__16_4_360_0.pdf.

Lafon:1991:LSS

[Laf91b] Jean-Pierre J. Lafon. Late stages of star evolution — foreword. Annales de Physique (Paris, France), 16(4):R1, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1991/04/anphys_1991__16_4_359_0/anphys_1991__16_4_359_
0.html;   http://www.annphys.org/articles/anphys/pdf/1991/
04/anphys_1991__16_4_359_0.pdf.

Lagniel:2000:AHI

[Lag00] J.-M. Lagniel. Accélérateurs de haute intensité. (French) [High intensity accelerators]. Annales de Physique (Paris, France), 25(2):269–??, ???? 2000. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.

annphys.org/articles/anphys/abs/2000/02/chap10/chap10.
html;   http://www.annphys.org/articles/anphys/pdf/2000/
02/chap10.pdf.

Lahaye:2006:RPE

[Lah06] T. Lahaye. Refroidissement par évaporation d’un jet atomique guidé magnétiquement. (French) [Evaporative cooling of a magnetically-guided atomic jet]. Annales de Physique (Paris, France), 31(1):1–??, ???? 2006. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://

www.annphys.org/articles/anphys/abs/2006/01/ann012006/
ann012006.html;   http://www.annphys.org/articles/anphys/
pdf/2006/01/ann012006.pdf;   http://www.annphys.org/articles/
anphys/ref/2006/01/ann012006/ann012006.html.

Laloe:1985:I

[Lal85] Franck Laloë. Introduction. Annales de Physique (Paris, France), 10(6):R1, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1985/06/anphys_1985__10_6_R7_0/anphys_
1985__10_6_R7_0.html;   http://www.annphys.org/articles/
anphys/pdf/1985/06/anphys_1985__10_6_R7_0.pdf.

Larose:2006:MUS

[Lar06] É. Larose. Mesoscopics of ultrasound and seismic waves: application to passive imaging. Annales de Physique (Paris, France), 31(3):1–??, ???? 2006. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://

www.annphys.org/articles/anphys/abs/2006/03/ann032006/
ann032006.html;   http://www.annphys.org/articles/anphys/
pdf/2006/03/ann032006.pdf;   http://www.annphys.org/articles/
anphys/ref/2006/03/ann032006/ann032006.html.

Lasota:1990:XRB

[Las90] J.-P. Lasota. X-ray binaries. Annales de Physique (Paris, France), 15(1):29–??, ???? 1990. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1990/01/anphys_1990__15_1_29_0/anphys_
1990__15_1_29_0.html;   http://www.annphys.org/articles/
anphys/pdf/1990/01/anphys_1990__15_1_29_0.pdf;   http:
//www.annphys.org/articles/anphys/ref/1990/01/anphys_1990_
_15_1_29_0/anphys_1990__15_1_29_0.html.

Lavagna:1988:SML

[Lav88] M. Lavagna. Superconductivity and Mott localization. Annales de Physique (Paris, France), 13(5):441–??, ???? 1988. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1988/05/anphys_1988__13_5_441_0/anphys_1988__13_5_441_
0.html;   http://www.annphys.org/articles/anphys/pdf/1988/
05/anphys_1988__13_5_441_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1988/05/anphys_1988__13_5_441_0/anphys_
1988__13_5_441_0.html.

Lavielle:1989:MF

[Lav89] Lisette Lavielle. La mouillabilité. (French) []. Annales de Physique (Paris, France), 14(1):1–??, ???? 1989. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1989/
01/anphys_1989__14_1_1_0/anphys_1989__14_1_1_0.html;
http://www.annphys.org/articles/anphys/pdf/1989/01/anphys_
1989__14_1_1_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1989/01/anphys_1989__14_1_1_0/anphys_1989__
14_1_1_0.html.

LHuillier:1992:NSC

[LB92] A. L’Huillier and P. Balcou. Une nouvelle source cohérente UVX : la génération d’harmoniques d’ordre élevé d’un rayonnement laser intense. (French) []. Annales de Physique (Paris, France), ??(??):181–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921735/
19921735.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921735.pdf.

Liger-Belair:2002:PBC

[LB02] G. Liger-Belair. La physique des bulles de champagne. — Une première approche des processus physico-chimiques liés à l’effervescence des vins de Champagne. (French) [The physics of champagne bubbles. — A first approach of physicochemical processes related to the effervescence of champagne]. Annales de Physique (Paris, France), 27(4):1–??, ???? 2002. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2002/
04/ann042002/ann042002.html;   http://www.annphys.org/
articles/anphys/pdf/2002/04/ann042002.pdf.

Liger-Belair:2006:NAE

[LB06] G. Liger-Belair. Nucléation, ascension et éclatement d’une bulle de champagne. (French) [Nucleation, rise and bursting of a champagne bubble]. Annales de Physique (Paris, France), 31(2):1–??, ???? 2006. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/2006/02/ann022006/ann022006.html;   http://
www.annphys.org/articles/anphys/pdf/2006/02/ann022006.
pdf;   http://www.annphys.org/articles/anphys/ref/2006/
02/ann022006/ann022006.html.

Lacour:1994:LEF

[LBB^{+}94] B. Lacour, H. Besaucele, H. Brunet, C. Gagnol, and B. Vincent. Lasers à excimères de forte puissance moyenne à grande cadence. (French) []. Annales de Physique (Paris, France), 19(C1):C1–205, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941951/
19941951.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941951.pdf.

Lazare:1994:TPP

[LBDM94] S. Lazare, M. Bolle, D. Drilhole, and S. Mihailov. Traitement des polymères par laser à excimères. (French) []. Annales de Physique (Paris, France), 19(C1):C1–269, October 1994. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1994/
07/19941959/19941959.html;   http://www.annphys.org/articles/
anphys/pdf/1994/07/19941959.pdf.

Lodewyck:2007:QKD

[LBF^{+}07] J. Lodewyck, M. Bloch, S. Fossier, E. Diamanti, T. Debuisschert, R. Tualle-Brouri, and P. Grangier. Quantum key distribution over 25 km using a fibre set-up based on continuous variables. Annales de Physique (Paris, France), 32(2–3):163–??, ???? 2007. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2007/
02/ann02200733/ann02200733.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200733.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200733/
ann02200733.html.

Legrand:1994:ECO

[LBFdR94] F. Legrand, G. Bridenne, G. Faussurier, and M. Sanchez del Rio. Étalonnage de cristaux organiques : simulation par SHADOW. (French) []. Annales de Physique (Paris, France), 19(C1):C1–155, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941932/
19941932.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941932.pdf.

Laporte:1992:SFG

[LBH92] P. Laporte, M. Bon, and M. Hemici. Source flash dans un gaz rare sous pression amorcée par plasma laser. (French) []. Annales de Physique (Paris, France), ??(??):67–??, ???? 1992. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1992/
07/19921714/19921714.html;   http://www.annphys.org/articles/
anphys/pdf/1992/07/19921714.pdf.

Larour:1992:FMQ

[LBS^{+}92] J. Larour, C. Barre, M. Skowronek, P. Romeas, and G. Vu-Tien. Facteurs de mérite de quelques sources X puisées miniatures à cathode creuse. (French) []. Annales de Physique (Paris, France), ??(??):183–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921736/
19921736.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921736.pdf.

Li:1988:SPB

[LCSY88] J. Q. Li, G. H. Chen, D. H. Shen, and Q. S. Yang. Superconducting phases in Bi-Sr-Ca-Cu-O system. Annales de Physique (Paris, France), 13(5):339–??, ???? 1988. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1988/05/anphys_1988__13_5_339_0/anphys_1988__13_5_339_
0.html;   http://www.annphys.org/articles/anphys/pdf/1988/
05/anphys_1988__13_5_339_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1988/05/anphys_1988__13_5_339_0/anphys_
1988__13_5_339_0.html.

Leger:1989:LAM

[Ld89] A. Léger and L. d’Hendecourt. Large aromatic molecules in the interstellar medium. Annales de Physique (Paris, France), 14(2):181–??, ???? 1989. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1989/02/anphys_1989__14_2_181_
0/anphys_1989__14_2_181_0.html;   http://www.annphys.org/
articles/anphys/pdf/1989/02/anphys_1989__14_2_181_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1989/02/anphys_
1989__14_2_181_0/anphys_1989__14_2_181_0.html.

Lehoucq:1991:NRL

[LD91] R. Lehoucq and Ph. Durouchouxa. Nuclear γ ray lines in the supernova SN 1987a. Annales de Physique (Paris, France), 16(4):443–??, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1991/04/anphys_1991__16_4_443_
0/anphys_1991__16_4_443_0.html;   http://www.annphys.org/
articles/anphys/pdf/1991/04/anphys_1991__16_4_443_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1991/04/anphys_
1991__16_4_443_0/anphys_1991__16_4_443_0.html.

Lopez:2007:FDK

[LDC^{+}07] O. Lopez, C. Daussy, C. Chardonnet, G. Santarelli, and A. Amy-Klein. Frequency dissemination with a 86-km optical fibre for fundamental tests of physics. Annales de Physique (Paris, France), 32(2–3):187–??, ???? 2007. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2007/
02/ann02200739/ann02200739.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200739.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200739/
ann02200739.html.

Lalo:1992:DNP

[LDCBA92] C. Lalo, J. Deson, C. Cerveau, and R. I. Ben-Aïm. Détection de NO dans la post-décharge d’un plasma de N_{2} /O_{2} par fluorescence induite par laser ArF. (French) []. Annales de Physique (Paris, France), ??(??):63–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921713/
19921713.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921713.pdf.

Lebrun:2007:SRS

[LDR07] S. Lebrun, P. Delaye, and G. Roosen. Stimulated Raman scattering in hollow core photonic crystal fibres. Annales de Physique (Paris, France), 32(2–3):45–??, ???? 2007. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2007/
02/ann02200705/ann02200705.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200705.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200705/
ann02200705.html.

Lefebvre:1990:DMO

[Lef90] R. Lefebvre. Dynamique moléculaire, oscillateurs couplés, résonances et réactivité. (French) []. Annales de Physique (Paris, France), 15(1):1–??, ???? 1990. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1990/01/anphys_1990__15_1_1_0/
anphys_1990__15_1_1_0.html;   http://www.annphys.org/articles/
anphys/pdf/1990/01/anphys_1990__15_1_1_0.pdf;   http:/
/www.annphys.org/articles/anphys/ref/1990/01/anphys_1990_
_15_1_1_0/anphys_1990__15_1_1_0.html.

Lehmann:1985:AOF

[Leh85] J. C. Lehmann. Allocution d’ouverture. (French) []. Annales de Physique (Paris, France), 10(6):525–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/06/anphys_1985__10_6_525_0/anphys_1985__10_6_525_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
06/anphys_1985__10_6_525_0.pdf.

Lemaire:1992:GRV

[Lem92] J. L. Lemaire. Génération de rayonnement VUV et XUV cohérent par conversion de fréquence dans les gaz et les vapeurs métalliques. Utilisation en physique atomique et moléculaire. (French) []. Annales de Physique (Paris, France), ??(??):159–??, ???? 1992. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1992/
07/19921729/19921729.html;   http://www.annphys.org/articles/
anphys/pdf/1992/07/19921729.pdf.

Lutke-Entrup:2000:AVS

[LEPMS00] N. Lütke-Entrup, B. Plaçais, P. Mathieu, and Y. Simon. Ancrage des vortex dans les supraconducteurs. — Description phénoménologique de la réponse linéaire d’un de vortex ancré. (French) [Anchor vortices in superconductors — phenomenological description of the linear response of an anchored vortex]. Annales de Physique (Paris, France), 25(4):1–??, ???? 2000. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/2000/04/placais/placais.html;   http://www.
annphys.org/articles/anphys/pdf/2000/04/placais.pdf.

Leray:2000:TDP

[Ler00] S. Leray. Transmutation des déchets : les perspectives. (French) [Waste transmutation: Prospects]. Annales de Physique (Paris, France), 25(2):293–??, ???? 2000. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.

annphys.org/articles/anphys/abs/2000/02/chap11/chap11.
html;   http://www.annphys.org/articles/anphys/pdf/2000/
02/chap11.pdf.

Leturcq:1995:EET

[Let95] R. Leturcq. Étude expérimentale de la transition métalisolant en dimension deux. (French) [experimental study of the metal–insulator transition in two dimensions]. Annales de Physique (Paris, France), 29(2–3):1–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2004/
02/ann022004/ann022004.html;   http://www.annphys.org/
articles/anphys/pdf/2004/02/ann022004.pdf;   http://www.
annphys.org/articles/anphys/ref/2004/02/ann022004/ann022004.
html.

Langlot:1995:RUT

[LFTV95] P. Langlot, N. Del Fatti, R. Tommasi, and F. Vallée. Relaxation ultrarapide des trous dans les semi-conducteurs. (French) []. Annales de Physique (Paris, France), 20(5–6):549–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05101995/ann05101995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05101995.pdf.

Lhuillier:1985:THP

[LL85] C. Lhuillier and M. Leduc. Transport dans l’hélium 3 polarisé par pompage optique. (French) []. Annales de Physique (Paris, France), 10(6):859–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1985/06/anphys_1985__10_6_859_
0/anphys_1985__10_6_859_0.html;   http://www.annphys.org/
articles/anphys/pdf/1985/06/anphys_1985__10_6_859_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1985/06/anphys_
1985__10_6_859_0/anphys_1985__10_6_859_0.html.

Landragin:1995:OFV

[LLC^{+}95] A. Landragin, G. Labeyrie, J. Y. Courtois, R. Kaiser, and N. Vansteenkiste. Observation de la force de van der Waals dans le miroir à atomes. (French) []. Annales de Physique (Paris, France), 20(5–6):641–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1995/05/ann05511995/ann05511995.html;
 http://www.annphys.org/articles/anphys/pdf/1995/05/ann05511995.
pdf.

Loiseaux:2000:PBC

[LM00]  J. M. Loiseaux and O. Meplan.  Principes de base et

caractéristiques particulières des systèmes sous-critiques assistés par accélérateur. (French) [Basic principles and specific characteristics of sub-critical systems assisted by accelerator]. Annales de Physique (Paris, France), 25(2):241– ??, ???? 2000. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/2000/02/chap9/chap9.html;   http:/
/www.annphys.org/articles/anphys/pdf/2000/02/chap9.pdf.

Lipavsky:2001:KES

[LMpk01] P. Lipavský, K. Morawetz, and V. pi ka. Kinetic equation for strongly interacting dense Fermi systems. Annales de Physique (Paris, France), 26(1):1–??, ???? 2001. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.

annphys.org/articles/anphys/abs/2001/01/annales012001/
annales012001.html;   http://www.annphys.org/articles/
anphys/pdf/2001/01/annales012001.pdf.

Loriette:1998:IMO

[Lor98] V. Loriette. Instruments de métrologie optique pour le projet Virgo. (French) [Optical metrology instruments for Project Virgo]. Annales de Physique (Paris, France), 23(2):1–??, ???? 1998. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/1998/02/ann021998/ann021998.html;   http://
www.annphys.org/articles/anphys/pdf/1998/02/ann021998.
pdf.

Lalanne:1999:OIS

[LPPC99] Ph. Lalanne, G. Prémont, D. Prévost, and P. Chavel. Optoelectronic implementation of stochastic artificial retinas. Annales de Physique (Paris, France), 24(3):125–??, ???? 1999. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1999/
03/abschap8/abschap8.html;   http://www.annphys.org/articles/
anphys/pdf/1999/03/abschap8.pdf.

Lefkir:1995:SEP

[LPvWR95] M. Lefkir, X. Nguyen Phu, A. J. van Wonderen, and G. Rivoire. Stabilité de l’état de polarisation elliptique en régime picoseconde. (French) []. Annales de Physique (Paris, France), 20(5–6):603–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1995/05/ann05321995/ann05321995.html;   http://www.
annphys.org/articles/anphys/pdf/1995/05/ann05321995.pdf.

Lopez-Rios:1987:SER

[LR87] T. Lopez-Rios. Surface-enhanced Raman scattering: some unexplained effects. Annales de Physique (Paris, France), 12(6):595–??, ???? 1987. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1987/06/anphys_1987__12_6_595_0/anphys_1987__12_6_
595_0.html;   http://www.annphys.org/articles/anphys/pdf/
1987/06/anphys_1987__12_6_595_0.pdf;   http://www.annphys.
org/articles/anphys/ref/1987/06/anphys_1987__12_6_595_
0/anphys_1987__12_6_595_0.html.

Larour:1992:MMG

[LRRVT92] J. Larour, P. Romeas, J. Rous, and G. Vu-Tien. MALIX : une machine 10 GW pour l’émission X intense entre 8 et 10 keV. (French)

[]. Annales de Physique (Paris, France), ??(??):187–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1992/07/19921737/19921737.html;   http://www.annphys.org/
articles/anphys/pdf/1992/07/19921737.pdf.

Lamy:1994:LXI

[LS94] T. Lamy and A. Simionovici. L’émission X des ions multichargés N^{7+}, O^{8+}, en interaction avec une surface métallique. (French) []. Annales de Physique (Paris, France), 19(C1):C1–177, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1994/07/19941943/19941943.html;   http://www.annphys.org/
articles/anphys/pdf/1994/07/19941943.pdf.

Ladieu:1996:PLL

[LS96] François Ladieu and Marc Sanquer. Pertinence et limitations de la loi de Mott dans les isolants désordonnés. (French) [Relevance and limits of Mott’s law in disordered insulators]. Annales de Physique (Paris, France), 21(3):267–??, ???? 1996. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1996/
03/ann031996/ann031996.html;   http://www.annphys.org/
articles/anphys/pdf/1996/03/ann031996.pdf.

leSueur:2008:CAS

[lS08] H. le Sueur. Cryogenic AFM–STM for mesoscopic physics. Annales de Physique (Paris, France), 33(6):1–??, ???? 2008. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2008/
04/ann62008/ann62008.html;   http://www.annphys.org/articles/
anphys/pdf/2008/04/ann62008.pdf;   http://www.annphys.
org/articles/anphys/ref/2008/04/ann62008/ann62008.html.

LHuillier:1994:RUC

[LSB94] A. L’Huillier, P. Salières, and Ph. Balcou. Rayonnement UVX créé par génération d’harmoniques d’un champ laser intense. (French) []. Annales de Physique (Paris, France), 19(C1):C1–111, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1994/07/19941926/19941926.html;   http://www.annphys.org/
articles/anphys/pdf/1994/07/19941926.pdf.

Laurens:1997:ULE

[LSD^{+}97] P. Laurens, B. Sadras, F. Decobert, F. Aréfï, and J. Amouroux. Utilisation des lasers excimères dans le domaine du collage des matériaux. (French) []. Annales de Physique (Paris, France), 22(C1):C1–197, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972240/
19972240.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972240.pdf.

Luminet:1985:EMR

[Lum85] J.-P. Luminet. Effets de marée : rupture explosive d’étoiles par un trou noir géant. (French) []. Annales de Physique (Paris, France), 10(2):101–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1985/02/anphys_1985__10_2_101_
0/anphys_1985__10_2_101_0.html;   http://www.annphys.org/
articles/anphys/pdf/1985/02/anphys_1985__10_2_101_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1985/02/anphys_
1985__10_2_101_0/anphys_1985__10_2_101_0.html.

Leroy-Willig:1991:DMR

[LWMG91] Anne Leroy-Willig, Daniel Morvan, and Geneviève Guillot. Does Magnetic Resonance Imaging belong to physicians or physicists? Annales de Physique (Paris, France), 16(2):133–??, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1991/02/anphys_1991__16_2_133_0/anphys_1991__16_2_133_
0.html;   http://www.annphys.org/articles/anphys/pdf/1991/
02/anphys_1991__16_2_133_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1991/02/anphys_1991__16_2_133_0/anphys_
1991__16_2_133_0.html.

Maal:1997:ORQ

[Maa97] Abdelhamid Maal. Oscillations de Rabi quantiques : test direct de la quantification du champ dans une cavité. (French) [Rabi nutations: immediate test of the field quantization in a cavity]. Annales de Physique (Paris, France), 22(4):405–??, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/1997/04/ann0401/ann0401.html;   http://www.
annphys.org/articles/anphys/pdf/1997/04/ann0401.pdf.

Mathiot:1991:RMD

[Mat91] J.-F. Mathiot. The relevance of mesonic degrees of freedom in nuclei. Annales de Physique (Paris, France), 16(5):527–??, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1991/05/anphys_1991__16_5_527_0/anphys_1991__16_5_527_
0.html;   http://www.annphys.org/articles/anphys/pdf/1991/
05/anphys_1991__16_5_527_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1991/05/anphys_1991__16_5_527_0/anphys_
1991__16_5_527_0.html.

Mathiot:1993:NLU

[Mat93] Jean-François Mathiot. Le noyau, laboratoire unique de la dynamique des quarks ! Défis théoriques. (French) []. Annales de Physique (Paris, France), 18(4):299–??, ???? 1993. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1993/04/anphys_1993__18_4_299_0/anphys_1993__18_4_299_
0.html;   http://www.annphys.org/articles/anphys/pdf/1993/
04/anphys_1993__18_4_299_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1993/04/anphys_1993__18_4_299_0/anphys_
1993__18_4_299_0.html.

Mazure:1985:LGE

[Maz85] A. Mazure. L’Univers à grande échelle les grandes structures et leur formation. (French) []. Annales de Physique (Paris, France), 10(5):415–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1985/05/anphys_1985__10_5_415_
0/anphys_1985__10_5_415_0.html;   http://www.annphys.org/
articles/anphys/pdf/1985/05/anphys_1985__10_5_415_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1985/05/anphys_
1985__10_5_415_0/anphys_1985__10_5_415_0.html.

Mazure:1989:MMU

[Maz89] A. Mazure. La masse manquante dans l’Univers. (French) []. Annales de Physique (Paris, France), 14(3):347–??, ???? 1989. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1989/03/anphys_1989__14_3_347_0/anphys_1989__14_3_347_
0.html;   http://www.annphys.org/articles/anphys/pdf/1989/
03/anphys_1989__14_3_347_0.pdf;   http://www.annphys.org/

articles/anphys/ref/1989/03/anphys_1989__14_3_347_0/anphys_
1989__14_3_347_0.html.

Mialocq:1997:SPA

[MB97] J. C. Mialocq and R. V. Bensasson. Spectroscopie et photochimie des acides nucleiques, des protéines et de leurs composants par éclair laser. (French) []. Annales de Physique (Paris, France), 22(C1):C1–169, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972236/
19972236.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972236.pdf.

Montagne:1992:LEK

[MBIA92] J. E. Montagne, J. Bouesc, G. Inglesakis, and M. Autric. Laser à exciplexe KrF impulsionnel de grande énergie. (French) []. Annales de Physique (Paris, France), ??(??):71–??, ???? 1992. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1992/
07/19921715/19921715.html;   http://www.annphys.org/articles/
anphys/pdf/1992/07/19921715.pdf.

Moliton:1988:ATE

[MDF88] A. Moliton, J. L. Duroux, and G. Froyer. Aspects théoriques et expérimentaux du dopage physique (implantation) des polymères électroactifs. (French) []. Annales de Physique (Paris, France), 13(4):261–??, ???? 1988. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1988/04/anphys_1988__13_4_261_
0/anphys_1988__13_4_261_0.html;   http://www.annphys.org/
articles/anphys/pdf/1988/04/anphys_1988__13_4_261_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1988/04/anphys_
1988__13_4_261_0/anphys_1988__13_4_261_0.html.

Mejean:2005:PIF

[Méj05] G. Méjean. Propagation d’impulsions femtosecondes térawatts dans l’atmosphère et applications. (French) [Propagation of terawatt femtosecond pulses in the atmosphere and applications]. Annales de Physique (Paris, France), 30(6):1–??, ???? 2005. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/2005/06/ann062005/ann062005.html;   http://
www.annphys.org/articles/anphys/pdf/2005/06/ann062005.

pdf;   http://www.annphys.org/articles/anphys/ref/2005/
06/ann062005/ann062005.html.

Melin:1997:SFV

[Mél97] R. Mélin. Systèmes ferromagnétiques vitreux. (French) [Glassy ferromagnetic systems]. Annales de Physique (Paris, France), 22(3):233–??, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1997/03/ann0301/ann0301.html;   http:
//www.annphys.org/articles/anphys/pdf/1997/03/ann0301.
pdf.

Mens:1994:UCC

[Men94] A. Mens. Utilisation de capteurs CCD pour l’imagerie et la spectroscopic XUV. (French) []. Annales de Physique (Paris, France), 19(C1):C1–139, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941930/
19941930.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941930.pdf.

Metcalf:1985:MTN

[Met85] H. Metcalf. Magnetic traps for neutral atoms. Annales de Physique (Paris, France), 10(6):733–??, ???? 1985. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/06/anphys_1985__10_6_733_0/anphys_1985__10_6_733_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
06/anphys_1985__10_6_733_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1985/06/anphys_1985__10_6_733_0/anphys_
1985__10_6_733_0.html.

Metivier:2000:PNP

[Mét00] H. Métivier. Principes et normes de la protection radiologique. (French) [Principles and norms of radiological protection]. Annales de Physique (Paris, France), 25(2):153– ??, ???? 2000. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/2000/02/chap5/chap5.html;   http:/
/www.annphys.org/articles/anphys/pdf/2000/02/chap5.pdf.

Meyer:1984:NLV

[Mey84] M. Meyer. Noyaux loin de la vallée de stabilité. excitations individuelles et collectives à basse énergie. (French) []. Annales de Physique (Paris, France), 9(5):791–??, ???? 1984. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1984/
05/anphys_1984__9_5_791_0/anphys_1984__9_5_791_0.html;
 http://www.annphys.org/articles/anphys/pdf/1984/05/anphys_
1984__9_5_791_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1984/05/anphys_1984__9_5_791_0/anphys_1984_
_9_5_791_0.html.

Meynet:1991:ISM

[Mey91] Georges Meynet. The initial stellar masses for the formation of white dwarfs, neutron stars and black holes. Annales de Physique (Paris, France), 16(4):459–??, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1991/04/anphys_1991__16_4_459_0/anphys_1991__16_4_459_
0.html;   http://www.annphys.org/articles/anphys/pdf/1991/
04/anphys_1991__16_4_459_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1991/04/anphys_1991__16_4_459_0/anphys_
1991__16_4_459_0.html.

Meyer:2003:IET

[Mey03] J. Meyer. Interactions effectives, théories de champ moyen, masses et rayons nucléaires. (French) [Effective interactions, mean-field theories, nuclear masses and radii]. Annales de Physique (Paris, France), 28(3):1–??, ???? 2003. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/2003/03/ann0032003/ann0032003.
html;   http://www.annphys.org/articles/anphys/pdf/2003/
03/ann0032003.pdf.

Morillon:1997:MUP

[MFG^{+}97] P. Morillon, J.-Cl. Farcy, F. Gilbert, G. Hélias, and G. Legeay. Micro-usinage par laser pour interconnexions micro-électroniques. (French) []. Annales de Physique (Paris, France), 22(C1):C1–237, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1997/07/19972246/19972246.html;   http://www.annphys.
org/articles/anphys/pdf/1997/07/19972246.pdf.

Michaud:2007:SLI

[MFM^{+}07a] J. Michaud, G. Fanjoux, H. Maillotte, L. Furfaro, and T. Sylvestre. Slow light induced by stimulated Raman scattering on spatial Kerr soliton. Annales de Physique (Paris, France), 32(2–3):103–??, ???? 2007. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
2007/02/ann02200718/ann02200718.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200718.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200718/
ann02200718.html.

Moldovan:2007:SUM

[MFM^{+}07b] I. C. Moldovan, V. Fesquet, F. Marc, H. Guillet de Chatellus, and J.-P. Pique. Sodium UV modeless laser excitation for PLGS. Annales de Physique (Paris, France), 32(2–3):91–??, ???? 2007. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2007/
02/ann02200715/ann02200715.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200715.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200715/
ann02200715.html.

Motret:1997:MTR

[MHN^{+}97] O. Motret, C. Hibert, M. Nikravech, S. Pellerin, R. Viladrosa, and J. M. Pouvesle. Mesure de la température rotationnelle de OH par spectroscopie UV dans une D.B.D. T.H.T. impulsionnelle. (French) []. Annales de Physique (Paris, France), 22(C1):C1–129, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972226/
19972226.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972226.pdf.

Michaud:1991:DLS

[Mic91] Georges Michaud. Diffusion in late stages of stellar evolution. Annales de Physique (Paris, France), 16(4):481–??, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1991/04/anphys_1991__16_4_481_0/anphys_1991__16_4_481_
0.html;   http://www.annphys.org/articles/anphys/pdf/1991/
04/anphys_1991__16_4_481_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1991/04/anphys_1991__16_4_481_0/anphys_
1991__16_4_481_0.html.

Mineev:1994:SU

[Min94] V. P. Mineev. Superconductivity in UPt_{3}. Annales de Physique (Paris, France), 19(4):367–??, ???? 1994. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1994/04/anphys_1994__19_4_367_0/anphys_1994__19_4_367_
0.html;   http://www.annphys.org/articles/anphys/pdf/1994/
04/anphys_1994__19_4_367_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1994/04/anphys_1994__19_4_367_0/anphys_
1994__19_4_367_0.html.

Miniatura:1995:IAB

[Min95] Ch. Miniatura. Interféromètrie atomique : un bref survol. (French) []. Annales de Physique (Paris, France), 20(5–6):487–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1995/05/ann05021995/ann05021995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05021995.pdf.

Miquet:1997:DRD

[Miq97] E. Miquet. Développements récents dans le domaine du microusinage par lasers excimères. (French) []. Annales de Physique (Paris, France), 22(C1):C1–245, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972247/
19972247.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972247.pdf.

Murray:2007:IPU

[MJM^{+}07] J. D. Murray, V. Jacques, F. Marquier, D. Chauvat, F. Grosshans, F. Treussart, and J.-F. Roch. Improvement of photostability using feedback from single-molecule fluorescence time trace. Annales de Physique (Paris, France), 32(2–3):135–??, ???? 2007. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2007/
02/ann02200726/ann02200726.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200726.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200726/
ann02200726.html.

Motret:1992:CDR

[MKG^{+}92] O. Motret, A. Khacef, M. Ganciu, C. Cachoncinlle, and M. Pouvesle. Caractérisation d’une décharge rapide dans Hg pur. (French) []. Annales de Physique (Paris, France), ??(??):73–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/

1992/07/19921716/19921716.html;   http://www.annphys.org/
articles/anphys/pdf/1992/07/19921716.pdf.

Millon:1997:LIL

[MKM97] E. Millón, B. Kubler, and J.-F. Mullert. L’ablation/ionisation laser appliquée à la préparation et à l’étude des fullerènes. (French) []. Annales de Physique (Paris, France), 22(C1):C1–261, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1997/07/19972250/19972250.html;   http://www.annphys.org/
articles/anphys/pdf/1997/07/19972250.pdf.

Moliton:1994:PES

[ML94] André Moliton and Bruno Lucas. Propriétés électroniques dans les solides désordonnés. III. Les mécanismes de transport par saut faisant appel aux modèles des paires et à la conduction par polarons. (French) []. Annales de Physique (Paris, France), 19(3):299–??, ???? 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1994/03/anphys_1994__19_3_299_
0/anphys_1994__19_3_299_0.html;   http://www.annphys.org/
articles/anphys/pdf/1994/03/anphys_1994__19_3_299_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1994/03/anphys_
1994__19_3_299_0/anphys_1994__19_3_299_0.html.

Mottin:1994:FIP

[MLCJ94] SJ. Mottin, P. Laporte, R. Cespuglio, and M. Jouvet. Fluorescence induite par laser sur les tissus profonds du cerveau du rat : application à l’étude du cycle veille-sommeil. (French) []. Annales de Physique (Paris, France), 19(C1):C1–187, October 1994. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1994/
07/19941948/19941948.html;   http://www.annphys.org/articles/
anphys/pdf/1994/07/19941948.pdf.

Mottin:1997:SED

[MLJC97] S. Mottin, P. Laporte, M. Jouvet, and R. Cespuglio. Spectroscopie d’émission du noyau Raphe Dorsalis chez le rat non anesthésié sous excitation laser UV 300–355 nm. (French) []. Annales de Physique (Paris, France), 22(C1):C1–179, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972237/

19972237.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972237.pdf.

Mercier:1999:LMP

[MLP99] R. Mercier, M. Lamare, and P. Picart. In-line measurement of phase-shift non linearity. Annales de Physique (Paris, France), 24(3):43–??, ???? 1999. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1999/03/abschap4/
abschap4.html;   http://www.annphys.org/articles/anphys/
pdf/1999/03/abschap4.pdf.

Minguzzi:1985:NCM

[MLT85] P. Minguzzi, A. Di Lieto, and M. Tonelli. Non-conventional methods in radiofrequency spectroscopy. Annales de Physique (Paris, France), 10(6):1065–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1985/06/anphys_1985__10_6_1065_
0/anphys_1985__10_6_1065_0.html;   http://www.annphys.
org/articles/anphys/pdf/1985/06/anphys_1985__10_6_1065_
0.pdf;   http://www.annphys.org/articles/anphys/ref/1985/
06/anphys_1985__10_6_1065_0/anphys_1985__10_6_1065_0.html.

Marc:2007:HPM

[MMdC^{+}07] F. Marc, I. C. Moldovan, H. Guillet de Chatellus, V. Fesquet, and J.-P. Pique. High power modeless lasers for sodium laser guide stars. Annales de Physique (Paris, France), 32(2–3):83–??, ???? 2007. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
2007/02/ann02200713/ann02200713.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200713.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200713/
ann02200713.html.

Misbah:1994:DFL

[MMK94] Chaouqi Misbah and Heiner Müller-Kurmbhaar. Dynamique d’une frontière libre : un siège fascinant de morphogénèse. (French) []. Annales de Physique (Paris, France), 19(6):601–??, ???? 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1994/06/anphys_1994__19_6_601_0/anphys_1994__19_6_601_
0.html;   http://www.annphys.org/articles/anphys/pdf/1994/
06/anphys_1994__19_6_601_0.pdf;   http://www.annphys.org/

articles/anphys/ref/1994/06/anphys_1994__19_6_601_0/anphys_
1994__19_6_601_0.html.

Motret:1994:CSP

[MMN^{+}94] O. Motret, V. Massereau, M. Nikravech, C. Hibert, J. M. Pouvesle, and J. Chapelle. Caractérisation spectroscopique d’un plasma de CH_{4} + CO_{2} obtenu par décharge à barrière diélectrique. (French) []. Annales de Physique (Paris, France), 19(C1):C1–157, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1994/07/19941933/19941933.html;   http://www.annphys.org/
articles/anphys/pdf/1994/07/19941933.pdf.

Mevel:1997:CDM

[MMZ^{+}97] E. Mével, I. Mercer, R. Zerne, A. L’Huillier, Ph. Antoine, and C.-G. Wahlström. Contrôle du mode spatial des harmoniques d’ordres élévés. (French) []. Annales de Physique (Paris, France), 22(C1):C1–73, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972212/
19972212.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972212.pdf.

Mazure:1984:EIM

[MN84] A. Mazure and G. Nollez. Effects of ion motion on Stark profiles of overlapping neutral helium lines. Annales de Physique (Paris, France), 9(4):675–??, ???? 1984. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1984/
04/anphys_1984__9_4_675_0/anphys_1984__9_4_675_0.html;
 http://www.annphys.org/articles/anphys/pdf/1984/04/anphys_
1984__9_4_675_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1984/04/anphys_1984__9_4_675_0/anphys_1984_
_9_4_675_0.html.

Museur:1994:OPP

[MNO^{+}94] L. Museur, A. Ndiaye, N. Ould, W. Q. Zheng, A. Kanaev, and M. C. Castex. Un outil performant pour la spectroscopie VUV (et les agrégats de gaz rare) : un laser à 125 nm de ≈ 10 J. (French) []. Annales de Physique (Paris, France), 19(C1):C1–159, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941934/

19941934.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941934.pdf.

Monthus:1995:EQF

[Mon95] Cécile Monthus. Étude de quelques fonctionnelles du mouvement brownien et de certaines propriétés de la diffusion unidimensionnelle en milieu aléatoire. (French) [On some functionals of Brownian motion with applications to winding properties of planar Brownian curves and to transport properties of one-dimensional diffusion in random media]. Annales de Physique (Paris, France), 20(4):341–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
04/ann04011995/ann04011995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/04/ann04011995.pdf.

Mott:1985:MIV

[Mot85] N. F. Mott. Materials with intermediate valence; a comparison with transition metals. Annales de Physique (Paris, France), 10(1):55–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1985/01/anphys_1985__10_1_55_0/anphys_
1985__10_1_55_0.html;   http://www.annphys.org/articles/
anphys/pdf/1985/01/anphys_1985__10_1_55_0.pdf;   http:
//www.annphys.org/articles/anphys/ref/1985/01/anphys_1985_
_10_1_55_0/anphys_1985__10_1_55_0.html.

Mehnaoui:2007:KPD

[MPT^{+}07] M. Mehnaoui, G. Panczer, R. Ternane, M. Trabelsi-Ayedi, and G. Boulon. Kinetics of Pb^{2+}-doped calcium hydroxyapatites fluorescence. Annales de Physique (Paris, France), 32(2–3):87–??, ???? 2007. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/2007/02/ann02200714/ann02200714.html;   http://www.
annphys.org/articles/anphys/pdf/2007/02/ann02200714.pdf;
 http://www.annphys.org/articles/anphys/ref/2007/02/ann02200714/
ann02200714.html.

Monchicourt:1984:CIU

[MPW84] P. Monchicourt, P. Pradel, and G. Watel. Collisional ionization under strong laser irradiation. Annales de Physique (Paris, France), 9(4):647–??, ???? 1984. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1984/04/anphys_1984__9_4_647_0/anphys_
1984__9_4_647_0.html;   http://www.annphys.org/articles/
anphys/pdf/1984/04/anphys_1984__9_4_647_0.pdf;   http:
//www.annphys.org/articles/anphys/ref/1984/04/anphys_1984_
_9_4_647_0/anphys_1984__9_4_647_0.html.

Moliton:1991:PESa

[MR91a] André Moliton and Bernard Ratier. Propriétés électroniques et schémas de bandes dans les semi-conducteurs amorphes. I. Concepts fondamentaux. (French) []. Annales de Physique (Paris, France), 16(3):261–??, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1991/03/anphys_1991__16_3_261_
0/anphys_1991__16_3_261_0.html;   http://www.annphys.org/
articles/anphys/pdf/1991/03/anphys_1991__16_3_261_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1991/03/anphys_
1991__16_3_261_0/anphys_1991__16_3_261_0.html.

Moliton:1991:PESb

[MR91b] André Moliton and Bernard Ratier. Propriétés électroniques et schémas de bandes dans les semi-conducteurs amorphes. II. Étude des phénomènes de transport. (French) []. Annales de Physique (Paris, France), 16(3):305–??, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1991/03/anphys_1991__16_3_305_0/anphys_1991__16_3_305_
0.html;   http://www.annphys.org/articles/anphys/pdf/1991/
03/anphys_1991__16_3_305_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1991/03/anphys_1991__16_3_305_0/anphys_
1991__16_3_305_0.html.

Misguich:2003:NIT

[MRC^{+}03] J. H. Misguich, J.-D. Reuss, D. Constantinescu, G. Steinbrecher, M. Vlad, F. Spineanu, B. Weyssow, and R. Balescu. Noble internal transport barriers and radial subdiffusion of toroidal magnetic lines. Annales de Physique (Paris, France), 28(6):1–??, ???? 2003. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/2003/06/ann062003/ann062003.html;   http://
www.annphys.org/articles/anphys/pdf/2003/06/ann062003.
pdf;   http://www.annphys.org/articles/anphys/ref/2003/
06/ann062003/ann062003.html.

Maurer:1988:ESY

[MRG^{+}88] M. Maurer, M. F. Ravet, T. Gourieux, G. Krill, A. Fontaine, H. Tolentino, E. Dartyge, and G. Tourillon. Electronic structure of YBa_{2}Cu_{3}O_{7−δ} as a function of δ: determination of the valency of copper atoms by X-ray absorption and photoemission experiments. Annales de Physique (Paris, France), 13(5):381–??, ???? 1988. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1988/05/anphys_1988__13_5_381_0/anphys_1988__13_5_381_
0.html;   http://www.annphys.org/articles/anphys/pdf/1988/
05/anphys_1988__13_5_381_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1988/05/anphys_1988__13_5_381_0/anphys_
1988__13_5_381_0.html.

Maleck-Rassoul:1995:EDD

[MRIC^{+}95] R. Maleck-Rassoul, A. Ivanov, F. Capelle, A. Le Calvez, E. Freysz, and A. Ducasse. Étude du doublage en fréquence d’impulsions femtosecondes issues d’un amplificateur régénératif Ti:saphir. (French) []. Annales de Physique (Paris, France), 20(5–6):651–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1995/05/ann05561995/ann05561995.html;   http://www.
annphys.org/articles/anphys/pdf/1995/05/ann05561995.pdf.

Mandon:2007:IFC

[MSS^{+}07] J. Mandon, E. Sorokin, I. T. Sorokina, G. Guelachvili, and N. Picqué. Infrared frequency combs and supercontinua for multiplex high sensitivity spectroscopy. Annales de Physique (Paris, France), 32(2–3):199–??, ???? 2007. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.

annphys.org/articles/anphys/abs/2007/02/ann02200742/ann02200742.
html;   http://www.annphys.org/articles/anphys/pdf/2007/
02/ann02200742.pdf;   http://www.annphys.org/articles/
anphys/ref/2007/02/ann02200742/ann02200742.html.

Muzy:1995:ADF

[Muz95] Jean-François Muzy. Analyse de distributions fractales à partir de leur transformée en ondelettes. (French) []. Annales de Physique (Paris, France), 20(2):63–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
02/ann02011995/ann02011995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/02/ann02011995.pdf.

Meyer:1992:SPA

[MV92] M. Meyer and J. P. Vivien. The superdeformation phenomenon in atomic nuclei. Annales de Physique (Paris, France), 17(1):11–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/1992/01/anphys_1992__17_1_11_0/anphys_1992_
_17_1_11_0.html;   http://www.annphys.org/articles/anphys/
pdf/1992/01/anphys_1992__17_1_11_0.pdf;   http://www.annphys.
org/articles/anphys/ref/1992/01/anphys_1992__17_1_11_0/
anphys_1992__17_1_11_0.html.

Menard:1995:REC

[MVCF95] S. Menard, M. Vampouille, B. Colombeau, and C. Froehly. Recombinaison extra-cavité de faisceaux lasers synchronisés yag/nd, pompés par diodes. (French) []. Annales de Physique (Paris, France), 20(5–6):571–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.

annphys.org/articles/anphys/abs/1995/05/ann05161995/ann05161995.
html;   http://www.annphys.org/articles/anphys/pdf/1995/
05/ann05161995.pdf.

Nicolas:1997:EPH

[NA97] G. Nicolas and M. Autric. Étude des phénomènes hydrodynamiques induits à la surface de matériaux solides par un rayonnement laser uv. (French) []. Annales de Physique (Paris, France), 22(C1):C1–131, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972227/
19972227.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972227.pdf.

Naud:2002:TQN

[Nau02] C. Naud. Transport quantique dans des nanostructures. (French) [Quantum transport in nanostructures]. Annales de Physique (Paris, France), 27(5):1–??, ???? 2002. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2002/
05/ann052002/ann052002.html;   http://www.annphys.org/
articles/anphys/pdf/2002/05/ann052002.pdf.

Nowicki-Bringuier:2007:HWG

[NBCB^{+}07] Y.-R. Nowicki-Bringuier, J. Claudon, C. Böckler, S. Reitzenstein, M. Kamp, A. Forchel, and J. M. Gérard. High-Q whispering gallery modes in pillar microcavities. Annales de Physique (Paris, France), 32(2–3):123–??, ???? 2007. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2007/
02/ann02200723/ann02200723.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200723.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200723/
ann02200723.html.

Nowicki-Bringuier:2007:NHE

[NBHC^{+}07] Y.-R. Nowicki-Bringuier, R. Hahner, J. Claudon, G. Lecamp, P. Lalanne, and J. M. Gérard. A novel high-efficiency singlemode single photon source. Annales de Physique (Paris, France), 32(2–3):151–??, ???? 2007. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/2007/02/ann02200730/ann02200730.html;
 http://www.annphys.org/articles/anphys/pdf/2007/02/ann02200730.
pdf;   http://www.annphys.org/articles/anphys/ref/2007/
02/ann02200730/ann02200730.html.

Nenner:1992:RSD

[Nen92] I. Nenner. Le rayonnement synchrotron dans le domaine XUV présent et futur. (French) []. Annales de Physique (Paris, France), ??(??):123–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921725/
19921725.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921725.pdf.

Nenner:1997:PSF

[Nen97] I. Nenner. Le projet SOLEIL. (French) [The SOLEIL [sun] project]. Annales de Physique (Paris, France), 22(C1):C1–11, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1997/07/19972202/19972202.html;   http://www.annphys.org/
articles/anphys/pdf/1997/07/19972202.pdf.

Neufeld:1994:MEQ

[NF94] M. Neufeld and R. Friedrich. Model equations for quasi-two-dimensional pattern forming systems. Annales de Physique (Paris, France), 19(6):721–??, ???? 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1994/06/anphys_1994__19_6_721_
0/anphys_1994__19_6_721_0.html;   http://www.annphys.org/
articles/anphys/pdf/1994/06/anphys_1994__19_6_721_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1994/06/anphys_
1994__19_6_721_0/anphys_1994__19_6_721_0.html.

Nissim:1994:SVP

[NFV^{+}94] Y. I. Nissim, J. Flicstein, Y. Vitel, O. Dulac, C. Debauche, and C. Licoppe. Source VUV pour les dépots photochimiques de couches minces diélectriques. (French) []. Annales de Physique (Paris, France), 19(C1):C1–229, October 1994. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1994/
07/19941954/19941954.html;   http://www.annphys.org/articles/
anphys/pdf/1994/07/19941954.pdf.

Ney:1994:ASX

[NG94] P. Ney and M. Rabec Le Gloahec. Applications des sources X et XUV à l’étude des plasmas créés par laser. (French) []. Annales de Physique (Paris, France), 19(C1):C1–87, October 1994. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1994/
07/19941923/19941923.html;   http://www.annphys.org/articles/
anphys/pdf/1994/07/19941923.pdf.

Nahon:1994:SML

[NLM^{+}94] L. Nahon, J. Lacoursière, M. Meyer, M. Larzillière, and P. Morin. Synchronization of a mode-locked Ar^{+} laser with the synchrotron radiation: application to the two-photon ionization (VUV + visible) of helium. Annales de Physique (Paris, France), 19(C1):C1–161, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941935/
19941935.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941935.pdf.

Nessib:1994:MES

[NLNA94] N. Ben Nessib, Z. Ben Lakhdar, H. Nguyen, and J. P. Arranz. Modélisation de l’élargissement Stark dans les lasers X. (French) []. Annales de Physique (Paris, France), 19(C1):C1–45, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1994/07/19941911/19941911.html;   http://www.annphys.org/
articles/anphys/pdf/1994/07/19941911.pdf.

Naudan:1995:FML

[NNH95] F. Naudan, C. Nicolas, and J.-P. Huignard. Filtrage des modes latéraux d’un laser semi-conducteur à cavité large par rétroinjection conjuguée. (French) []. Annales de Physique (Paris, France), 20(5–6):599–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.

annphys.org/articles/anphys/abs/1995/05/ann05301995/ann05301995.
html;   http://www.annphys.org/articles/anphys/pdf/1995/
05/ann05301995.pdf.

Nottale:1988:GLO

[Not88] L. Nottale. Gravitational lensing in observational cosmology. Annales de Physique (Paris, France), 13(4):223–??, ???? 1988. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1988/04/anphys_1988__13_4_223_0/anphys_1988__13_4_223_
0.html;   http://www.annphys.org/articles/anphys/pdf/1988/
04/anphys_1988__13_4_223_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1988/04/anphys_1988__13_4_223_0/anphys_
1988__13_4_223_0.html.

Nozieres:1985:IME

[Noz85] P. Nozières. Impuretés magnétiques et effet Kondo. (French) []. Annales de Physique (Paris, France), 10(1):19–??, ???? 1985. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1985/
01/anphys_1985__10_1_19_0/anphys_1985__10_1_19_0.html;
 http://www.annphys.org/articles/anphys/pdf/1985/01/anphys_
1985__10_1_19_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1985/01/anphys_1985__10_1_19_0/anphys_1985_
_10_1_19_0.html.

Neve:1995:ESD

[NVJ^{+}95] M. Le Nevé, J. C. Vial, A. Jolly, H. Coïc, P. Féru, P. Maruenda, and J. Vicrey. Étude de la saturation du Cr^{4+}: YAG pompé à 1,06 μ m. (French) []. Annales de Physique (Paris, France), 20 (5–6):623–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1995/05/ann05421995/ann05421995.html;
 http://www.annphys.org/articles/anphys/pdf/1995/05/ann05421995.
pdf.

Ousset:1988:UCF

[OBU^{+}88] J. C. Ousset, J. F. Bobo, J. P. Ulmet, H. Rakoto, and N. Cheggour. Upper critical field measurements in high-T_{c} superconducting oxides. Annales de Physique (Paris, France), 13(5):387–??, ???? 1988. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1988/05/anphys_1988__13_5_387_0/anphys_1988__13_5_387_
0.html;   http://www.annphys.org/articles/anphys/pdf/1988/
05/anphys_1988__13_5_387_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1988/05/anphys_1988__13_5_387_0/anphys_
1988__13_5_387_0.html.

Ohayon:1994:TMM

[OSA94] E. Ohayon, Th. Samet, and M. Autric. Traitement de matériaux métalliques par choc laser en géométrie confinée à 248 nm. (French) []. Annales de Physique (Paris, France), 19(C1):C1–163, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1994/07/19941936/19941936.html;   http://www.annphys.org/
articles/anphys/pdf/1994/07/19941936.pdf.

Otten:1985:OPU

[Ott85] E. W. Otten. Optical pumping and unstable nuclei. Annales de Physique (Paris, France), 10(6):659–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/06/anphys_1985__10_6_659_0/anphys_1985__10_6_659_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
06/anphys_1985__10_6_659_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1985/06/anphys_1985__10_6_659_0/anphys_
1985__10_6_659_0.html.

Ott:1997:MIP

[Ott97] A. Ott. Micromanipulation intracellulaire par pinces optiques. Annales de Physique (Paris, France), 22(C1):C1–181, ???? 1997. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1997/

07/19972238/19972238.html;   http://www.annphys.org/articles/
anphys/pdf/1997/07/19972238.pdf.

Ourjoumtsev:2007:ETE

[Our07] A. Ourjoumtsev. Étude théorique et expérimentale de superpositions quantiques cohérentes et d’états intriqués non-gaussiens de la lumière. (French) [Theoretical and experimental study of coherent quantum superpositions and entangled non-Gaussian states of light]. Annales de Physique (Paris, France), 32(2–3):1–??, ???? 2007. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/2007/03/ann042007/ann042007.html;   http://
www.annphys.org/articles/anphys/pdf/2007/03/ann042007.
pdf;   http://www.annphys.org/articles/anphys/ref/2007/
03/ann042007/ann042007.html.

Ovchinnikov:1994:PCS

[Ovc94] Yu. N. Ovchinnikov. Pinning and creep in superconductors. Annales de Physique (Paris, France), 19(4):385–??, ???? 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1994/04/anphys_1994__19_4_385_0/anphys_1994__19_4_385_
0.html;   http://www.annphys.org/articles/anphys/pdf/1994/
04/anphys_1994__19_4_385_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1994/04/anphys_1994__19_4_385_0/anphys_
1994__19_4_385_0.html.

Owen:1986:LSM

[Owe86] D. A. Owen. Lamb shift in muonium: a review of experiment and theory. Annales de Physique (Paris, France), 11(3):249–??, ???? 1986. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1986/03/anphys_1986__11_3_249_0/anphys_1986__11_3_249_
0.html;   http://www.annphys.org/articles/anphys/pdf/1986/
03/anphys_1986__11_3_249_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1986/03/anphys_1986__11_3_249_0/anphys_
1986__11_3_249_0.html.

Paris:1984:RIM

[Par84] R. B. Paris. Resistive instabilities in MHD. Annales de Physique (Paris, France), 9(2):347–??, ???? 1984. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1984/
02/anphys_1984__9_2_347_0/anphys_1984__9_2_347_0.html;
 http://www.annphys.org/articles/anphys/pdf/1984/02/anphys_
1984__9_2_347_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1984/02/anphys_1984__9_2_347_0/anphys_1984_
_9_2_347_0.html.

Pascoli:1986:AGE

[Pas86] G. Pascoli. Les anisotropies à grande échelle dans les nébuleuses planétaires. (French) []. Annales de Physique (Paris, France), 11(3):297–??, ???? 1986. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1986/03/anphys_1986__11_3_297_
0/anphys_1986__11_3_297_0.html;   http://www.annphys.org/
articles/anphys/pdf/1986/03/anphys_1986__11_3_297_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1986/03/anphys_
1986__11_3_297_0/anphys_1986__11_3_297_0.html.

Paterson:1996:MSH

[Pat96] Agnès Paterson. Mouillage de surfaces hétérogènes: cellule de Hele–Shaw imparfaite. (French) [Wetting on heterogeneous surfaces: imperfect Hele–Shaw cell]. Annales de Physique (Paris, France), 21(4):337–??, ???? 1996. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1996/
04/ann041996/ann041996.html;   http://www.annphys.org/
articles/anphys/pdf/1996/04/ann041996.pdf.

Presles:1989:ERP

[PBD89] H. N. Presles, P. Bauer, and D. Desbordes. Étude de la répartition des perturbations sur un front d’onde de détonation dans un mélange explosif gazeux. (French) []. Annales de Physique (Paris, France), 14(6):623–??, ???? 1989. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1989/06/anphys_1989__14_6_623_0/anphys_1989__14_6_623_
0.html;   http://www.annphys.org/articles/anphys/pdf/1989/
06/anphys_1989__14_6_623_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1989/06/anphys_1989__14_6_623_0/anphys_
1989__14_6_623_0.html.

Papageorgiou:1995:EAB

[PBD^{+}95] N. Papageorgiou, D. Bloch, M. Ducloy, B. Gross, F. Baur, and A. Weis. Étude de l’anisotropie (biréfringinces) induite par les intéractions de surface. (French) []. Annales de Physique (Paris, France), 20(5–6):611–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05361995/ann05361995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05361995.pdf.

Parizet:1984:EPL

[PBG^{+}84] M. J. Parizet, E. Borie, B. Grossetête, D. B. Isabelle, and J. Proriol. Electromagnetic processes at low momentum transfer: a review for users. Annales de Physique (Paris, France), 9(1):103–??, ???? 1984. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1984/01/anphys_1984__9_1_103_0/anphys_1984__9_1_103_0.
html;   http://www.annphys.org/articles/anphys/pdf/1984/
01/anphys_1984__9_1_103_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1984/01/anphys_1984__9_1_103_0/anphys_
1984__9_1_103_0.html.

Pasquiers:1994:EEM

[PBL^{+}94] S. Pasquiers, M.-C. Bordage, M. Legentil, V. Puech, R. Riva, and P. Ségur. Étude expérimentale et modélisation des lasers à excimère XeCl photo-déclenchés. (French) []. Annales de Physique (Paris, France), 19(C1):C1–213, October 1994. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1994/
07/19941952/19941952.html;   http://www.annphys.org/articles/
anphys/pdf/1994/07/19941952.pdf.

Pouvesle:1992:CEU

[PC92] J. M. Pouvesle and C. Cachoncinlle. Continu d’émission UV– VUV du néon dans des plasmas à haute pression. (French) []. Annales de Physique (Paris, France), ??(??):77–??, ???? 1992. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1992/
07/19921717/19921717.html;   http://www.annphys.org/articles/
anphys/pdf/1992/07/19921717.pdf.

Philippet:2007:SLV

[PCFC07] L. Philippet, S. Chénais, S. Forget, and M.-C. Castex. Study of laser vaporization of mercury: application to the energy enhancement of a VUV coherent source. Annales de Physique (Paris, France), 32(2–3):95–??, ???? 2007. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2007/
02/ann02200716/ann02200716.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200716.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200716/
ann02200716.html.

Pankoke:1995:GAP

[PCS^{+}95] R. Pankoke, A. Cassineli, P. Sillard, C. Dorrer, P. Leisching, and R. Frey. Générateurs et amplificateurs paramétriques optiques monomodes transverse. (French) []. Annales de Physique (Paris, France), 20(5–6):593–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05271995/ann05271995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05271995.pdf.

Paillard:1989:ICC

[PDAYS89] C. Paillard, G. Dupré, H. Al Aiteh, and S. Youssefi-Stitou. Influence de la cinétique chimique sur la structure des détonations : cas du bioxyde de chlore. (French) []. Annales de Physique (Paris, France), 14(6):641–??, ???? 1989. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1989/06/anphys_1989__14_6_641_
0/anphys_1989__14_6_641_0.html;   http://www.annphys.org/
articles/anphys/pdf/1989/06/anphys_1989__14_6_641_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1989/06/anphys_
1989__14_6_641_0/anphys_1989__14_6_641_0.html.

Petite:1997:PFA

[PDGM97] G. Petite, Ph. Daguzan, S. Guizard, and Ph. Martin. Processus fondamentaux d’absorption et de relaxation de l’énergie dans les interactions laser-surface. (French) []. Annales de Physique (Paris, France), 22(C1):C1–81, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972215/
19972215.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972215.pdf.

Pecker:1985:PON

[Pec85] J.-C. Pecker. Le pompage optique naturel dans le milieu astrophysique. (French) []. Annales de Physique (Paris, France), 10(6):631–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1985/06/anphys_1985__10_6_631_
0/anphys_1985__10_6_631_0.html;   http://www.annphys.org/
articles/anphys/pdf/1985/06/anphys_1985__10_6_631_0.pdf.

Pelce:2004:CFH

[Pel04] P. Pelcé. Croissance et forme de l’hyphe modelées par une instabilité. (French) [Growth and shape of the peg modeled by an instability]. Annales de Physique (Paris, France), 29(4):1–??, ???? 2004. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/2004/04/ann2004052002/ann2004052002.html;   http:/
/www.annphys.org/articles/anphys/pdf/2004/04/ann2004052002.
pdf.

Perronet:2005:EPD

[Per05] K. Perronet. Étude par détection de photons des processus électroniques au sein d’une jonction tunnel dans un milieu moléculaire. (French) [Study of photon detection by electronic processes within a tunnel junction in a molecular environment]. Annales de Physique (Paris, France), 30(2):1–??, ???? 2005. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/2005/02/ann022005/ann022005.html;   http://
www.annphys.org/articles/anphys/pdf/2005/02/ann022005.
pdf;   http://www.annphys.org/articles/anphys/ref/2005/
02/ann022005/ann022005.html.

Petit:1991:EMS

[Pet91] L. Petit. Écoulement et mise en structure de suspensions macroscopiques. (French) []. Annales de Physique (Paris, France), 16(2):155–??, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1991/02/anphys_1991__16_2_155_
0/anphys_1991__16_2_155_0.html;   http://www.annphys.org/
articles/anphys/pdf/1991/02/anphys_1991__16_2_155_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1991/02/anphys_
1991__16_2_155_0/anphys_1991__16_2_155_0.html.

Petrignani:1994:POA

[Pet94] J. F. Petrignani. Procédés d’oxydation avancés : utilisation de rayonnement UV dans le domaine de l’environnement. (French) []. Annales de Physique (Paris, France), 19(C1):C1–221, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1994/07/19941953/19941953.html;   http://www.annphys.org/
articles/anphys/pdf/1994/07/19941953.pdf.

Petitjean:1999:CBU

[Pet99] P. Petitjean. Le contenu baryonique de l’univers révélé par les raies d’absorption dans le spectre des quasars. (French) [Baryon content of the Universe revealed by the absorption lines in the quasar spectrum]. Annales de Physique (Paris, France), 24(1):1– ??, ???? 1999. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/1999/01/ann0199/ann0199.html;   http://www.
annphys.org/articles/anphys/pdf/1999/01/ann0199.pdf.

Pellat-Finet:1993:ROF

[PF93] Pierre Pellat-Finet. Sur les rotations de 2 et 4 en optique. (French) []. Annales de Physique (Paris, France), 18(3):269–??, ???? 1993. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1993/03/anphys_1993__18_3_269_0/anphys_1993__18_3_269_
0.html;   http://www.annphys.org/articles/anphys/pdf/1993/
03/anphys_1993__18_3_269_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1993/03/anphys_1993__18_3_269_0/anphys_
1993__18_3_269_0.html.

Perrier:1991:VII

[PFC91] Ch. Perrier, R. Foy, and A. Chalabaev. Visible and infrared interferometry before and with the VLT. Annales de Physique (Paris, France), 16(4):497–??, ???? 1991. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1991/04/anphys_1991__16_4_497_0/anphys_1991__16_4_497_
0.html;   http://www.annphys.org/articles/anphys/pdf/1991/
04/anphys_1991__16_4_497_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1991/04/anphys_1991__16_4_497_0/anphys_
1991__16_4_497_0.html.

Phillips:1985:LCT

[Phi85] W. D. Phillips. Laser-cooling and trapping neutral atoms. Annales de Physique (Paris, France), 10(6):717–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/06/anphys_1985__10_6_717_0/anphys_1985__10_6_717_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
06/anphys_1985__10_6_717_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1985/06/anphys_1985__10_6_717_0/anphys_
1985__10_6_717_0.html.

Pierre:2000:PAG

[Pie00] M. Pierre. Le point sur les amas de galaxies. (French) [The point of clusters of galaxies]. Annales de Physique (Paris, France), 25(5):1–??, ???? 2000. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/2000/05/ann05/ann05.html;   http:/
/www.annphys.org/articles/anphys/pdf/2000/05/ann05.pdf.

Pierre:2001:IEE

[Pie01] F. Pierre. Interactions électron-électron dans les fils mésoscopiques. (French) [Electron–electron interactions in mesoscopic wires]. Annales de Physique (Paris, France), 26(4):1–??, ???? 2001. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic).  URL http://www.annphys.org/articles/anphys/
abs/2001/04/annales042001/annales042001.html;   http:/
/www.annphys.org/articles/anphys/pdf/2001/04/annales042001.
pdf.

Pinard:1985:SAI

[PJ85] M. Pinard and L. Julien. La sélection d’atomes immobiles par pompage optique. (French) []. Annales de Physique (Paris, France), 10(6):1007–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1985/06/anphys_1985__10_6_1007_
0/anphys_1985__10_6_1007_0.html;   http://www.annphys.
org/articles/anphys/pdf/1985/06/anphys_1985__10_6_1007_
0.pdf;   http://www.annphys.org/articles/anphys/ref/1985/
06/anphys_1985__10_6_1007_0/anphys_1985__10_6_1007_0.html.

Polack:1997:SPT

[PJ97]  F. Polack and D. Joyeux.  Spectrométrie par TF dans

les VUV. Perspectives offertes par des interféromètres sans lames séparatrices. (French) []. Annales de Physique (Paris, France), 22(C1):C1–151, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972233/
19972233.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972233.pdf.

Pribat:1997:CDS

[PLP^{+}97] D. Pribat, P. Legagneux, F. Plais, F. Petinot, O. Huet, and C. Reita. Cristallisation du silicium amorphe par laser à excimères. (French) []. Annales de Physique (Paris, France), 22(C1):C1–213, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1997/07/19972242/19972242.html;   http://www.annphys.
org/articles/anphys/pdf/1997/07/19972242.pdf.

Prat:1994:CMC

[PNCA94] Ch. Prat, G. Nicolas-Costa, and M. Autric. Caractérisation de matériaux céramiques soumis à rayonnements laser UV : applications au traitement de surface. (French) []. Annales de Physique (Paris, France), 19(C1):C1–165, October 1994. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1994/
07/19941937/19941937.html;   http://www.annphys.org/articles/
anphys/pdf/1994/07/19941937.pdf.

Ponson:2007:CPD

[Pon07] L. Ponson. Crack propagation in disordered materials: how to decipher fracture surfaces. Annales de Physique (Paris, France), 32(1):1–??, ???? 2007. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://

www.annphys.org/articles/anphys/abs/2007/01/ann012007/
ann012007.html;   http://www.annphys.org/articles/anphys/
pdf/2007/01/ann012007.pdf;   http://www.annphys.org/articles/
anphys/ref/2007/01/ann012007/ann012007.html.

Pouey:1992:ILP

[Pou92] M. Pouey. Intensité lumineuse et photométrie. (French) []. Annales de Physique (Paris, France), ??(??):95–??, ???? 1992. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1992/
07/19921722/19921722.html;   http://www.annphys.org/articles/
anphys/pdf/1992/07/19921722.pdf.

Pouvesle:1997:GRX

[Pou97] J. M. Pouvesle. Génération de rayonnement X cohérent et incohérent par décharges et micro-décharges capillaires : progrès récents et perspectives. (French) []. Annales de Physique (Paris, France), 22(C1):C1–53, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972208/
19972208.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972208.pdf.

Pebay-Peyroula:1985:QEP

[PP85] J. C. Pébay-Peyroula. Quelques expériences de physique moléculaire. (French) []. Annales de Physique (Paris, France), 10(6):1087–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1985/06/anphys_1985__10_6_1087_0/anphys_
1985__10_6_1087_0.html;   http://www.annphys.org/articles/
anphys/pdf/1985/06/anphys_1985__10_6_1087_0.pdf;   http:
//www.annphys.org/articles/anphys/ref/1985/06/anphys_1985_
_10_6_1087_0/anphys_1985__10_6_1087_0.html.

Peres:1992:MPP

[PP92a] I. Pérès and L. C. Pitchford. Modélisation des plasmas produits par décharge électrique. (French) []. Annales de Physique (Paris, France), ??(??):21–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921704/
19921704.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921704.pdf.

Pouvesle:1992:GUX

[PP92b] J. M. Pouvesle and D. Pigache. Génération d’UV-X mous par décharge électrique. (French) []. Annales de Physique (Paris, France), ??(??):229–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921743/
19921743.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921743.pdf.

Pousset:1997:EDR

[PPEZ97] T. Pousset, D. Porras, H. Elloumi, and G. Zissis. Étude du rayonnement UV d’une lampe à décharge mercure haute pression en fonction de son mode d’alimentation. (French) []. Annales de Physique (Paris, France), 22(C1):C1–75, ???? 1997. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1997/
07/19972213/19972213.html;   http://www.annphys.org/articles/
anphys/pdf/1997/07/19972213.pdf.

Poizat:1994:CQN

[PRG94] J.-Ph. Poizat, J.-F. Roch, and P. Grangier. Characterization of quantum non-demolition measurements in optics. Annales de Physique (Paris, France), 19(3):265–??, ???? 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1994/03/anphys_1994__19_3_265_0/anphys_1994__19_3_265_
0.html;   http://www.annphys.org/articles/anphys/pdf/1994/
03/anphys_1994__19_3_265_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1994/03/anphys_1994__19_3_265_0/anphys_
1994__19_3_265_0.html.

Peyranere:1988:WRC

[PT88] M. Capdequi Peyranere and M. Talon. Weak radiative corrections at LEP. Annales de Physique (Paris, France), 13(1):37–??, ???? 1988. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1988/01/anphys_1988__13_1_37_0/anphys_1988__13_1_37_0.
html;   http://www.annphys.org/articles/anphys/pdf/1988/
01/anphys_1988__13_1_37_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1988/01/anphys_1988__13_1_37_0/anphys_
1988__13_1_37_0.html.

Paulin:1999:TNL

[PTC99] C. Paulin, J. Taboury, and P. Chavel. Traitement non linéaire d’images stéréo par corrélation de phase locale. (French) [Nonlinear stereo image processing by local phase correlation]. Annales de Physique (Paris, France), 24(3):103–??, ???? 1999. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1999/
03/abschap7/abschap7.html;   http://www.annphys.org/articles/
anphys/pdf/1999/03/abschap7.pdf.

Peremans:1995:SVI

[PTZR95] A. Peremans, A. Tadjeddine, W.-Q. Zheng, and A. Le Rille. Spectroscopie vibrationelle d’interface par génération de fréquencesomme. (French) []. Annales de Physique (Paris, France), 20 (5–6):527–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1995/05/ann05071995/ann05071995.html;
 http://www.annphys.org/articles/anphys/pdf/1995/05/ann05071995.
pdf.

Pouvesle:1992:ECO

[PV92] J. M. Pouvesle and R. Viladrosa. Étude, caractérisation et optimisation de lampes flash microseconde et submicroseconde dans le domaine 200-300 nm. (French) []. Annales de Physique (Paris, France), ??(??):79–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921718/
19921718.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921718.pdf.

Poirson:1995:MFA

[PVJ^{+}95] J. Poirson, M. Vallet, D. Jacobs, F. Bretenaker, and A. Le Floch. Mesures de faibles anisotropies intracavités au moyen de cavités fabry-perot de grande finesse. (French) []. Annales de Physique (Paris, France), 20(5–6):607–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05341995/ann05341995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05341995.pdf.

Quere:1998:MDF

[QdR98] D. Quéré and A. de Ryck. Le mouillage dynamique des fibres. (French) [Dynamic wetting of fibers]. Annales de Physique (Paris, France), 23(1):1–??, ???? 1998. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1998/
01/ann011998/ann011998.html;   http://www.annphys.org/
articles/anphys/pdf/1998/01/ann011998.pdf.

Quentrec:1984:VDD

[Que84] B. Quentrec. Viscoélasticité et diffusion dépolarisée de la lumière dans les liquides. (French) []. Annales de Physique (Paris, France), 9(1):1–??, ???? 1984. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1984/01/anphys_1984__9_1_1_0/anphys_
1984__9_1_1_0.html;   http://www.annphys.org/articles/
anphys/pdf/1984/01/anphys_1984__9_1_1_0.pdf;   http://

www.annphys.org/articles/anphys/ref/1984/01/anphys_1984_
_9_1_1_0/anphys_1984__9_1_1_0.html.

Quebert:1992:FCI

[Qué92] Jean Québert. Les fonctions de corrélation d’intensités en physique nucléaire : un accès possible à la dynamique des collisions. (French) []. Annales de Physique (Paris, France), 17(2):99–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1992/02/anphys_1992__17_2_99_0/anphys_1992__17_2_99_0.
html;   http://www.annphys.org/articles/anphys/pdf/1992/
02/anphys_1992__17_2_99_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1992/02/anphys_1992__17_2_99_0/anphys_
1992__17_2_99_0.html.

Quillec:1995:PQS

[Qui95] M. Quillec. Puits quantiques et superréseaux à base de III–V pour l’optoélectronique: fabrication et usages. (French) []. Annales de Physique (Paris, France), 20(5–6):735–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05681995/ann05681995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05681995.pdf.

Rabaud:1994:DII

[Rab94] Marc Rabaud. Dynamiques interfaciales dans l’instabilité de l’imprimeur. (French) []. Annales de Physique (Paris, France), 19(6):659–??, ???? 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1994/06/anphys_1994__19_6_659_
0/anphys_1994__19_6_659_0.html;   http://www.annphys.org/
articles/anphys/pdf/1994/06/anphys_1994__19_6_659_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1994/06/anphys_
1994__19_6_659_0/anphys_1994__19_6_659_0.html.

Ramsey:1985:NMR

[Ram85] N. F. Ramsey. Neutron magnetic resonance. Annales de Physique (Paris, France), 10(6):945–??, ???? 1985. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/06/anphys_1985__10_6_945_0/anphys_1985__10_6_945_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
06/anphys_1985__10_6_945_0.pdf;   http://www.annphys.org/

articles/anphys/ref/1985/06/anphys_1985__10_6_945_0/anphys_
1985__10_6_945_0.html.

Ranson:1984:AP

[Ran84] P. Ranson. Avant-propos. Annales de Physique (Paris, France), 9(4):R1, ???? 1984. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1984/04/anphys_1984__9_4_R9_0/
anphys_1984__9_4_R9_0.html;   http://www.annphys.org/articles/
anphys/pdf/1984/04/anphys_1984__9_4_R9_0.pdf;   http:/
/www.annphys.org/articles/anphys/ref/1984/04/anphys_1984_
_9_4_R9_0/anphys_1984__9_4_R9_0.html.

Rao:1988:SSS

[Rao88] C. N. R. Rao. Structure, stoichiometry and spectroscopy of oxide superconductors. Annales de Physique (Paris, France), 13(5):327–??, ???? 1988. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1988/05/anphys_1988__13_5_327_
0/anphys_1988__13_5_327_0.html;   http://www.annphys.org/
articles/anphys/pdf/1988/05/anphys_1988__13_5_327_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1988/05/anphys_
1988__13_5_327_0/anphys_1988__13_5_327_0.html.

Rigneault:1995:PRA

[RAR^{+}95] H. Rigneault, C. Amra, S. Robert, S. Monneret, and F. Flory. Propriétés radiatives d’atomes luminescents placés dans une microcavité planaire. (French) []. Annales de Physique (Paris, France), 20(5–6):625–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1995/05/ann05431995/ann05431995.html;
 http://www.annphys.org/articles/anphys/pdf/1995/05/ann05431995.
pdf.

Raulin:1989:COP

[Rau89] François Raulin. Chimie organique prébiotique dans des environnements planétaires. (French) []. Annales de Physique (Paris, France), 14(2):207–??, ???? 1989. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1989/02/anphys_1989__14_2_207_
0/anphys_1989__14_2_207_0.html;   http://www.annphys.org/
articles/anphys/pdf/1989/02/anphys_1989__14_2_207_0.pdf;

 http://www.annphys.org/articles/anphys/ref/1989/02/anphys_
1989__14_2_207_0/anphys_1989__14_2_207_0.html.

Rayez:1991:EDR

[RB91] J. C. Rayez and L. Bonnet. Erratum — dynamique de réactions indirectes du type A + BC → ABC → AB + C et chaos conservatif. (French) []. Annales de Physique (Paris, France), 16(2):259–??, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1991/02/anphys_1991__16_2_259_0/anphys_1991__16_2_
259_0.html;   http://www.annphys.org/articles/anphys/pdf/
1991/02/anphys_1991__16_2_259_0.pdf.

Roisse:1995:ALM

[RB95] E. Roissé and A. Barthélémy. Amplificateur laser multipassage pompé par diode. (French) []. Annales de Physique (Paris, France), 20(5–6):627–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1995/05/ann05441995/ann05441995.html;
 http://www.annphys.org/articles/anphys/pdf/1995/05/ann05441995.
pdf.

Rowe:1997:ERR

[RC97] B. R. Rowe and A. Canosa. Étude de réactions radicalaires à très basses températures. Implications astrochimiques. (French) []. Annales de Physique (Paris, France), 22(C1):C1–89, ???? 1997. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1997/
07/19972216/19972216.html;   http://www.annphys.org/articles/
anphys/pdf/1997/07/19972216.pdf.

Robert:1994:SFX

[RCK^{+}94] E. Robert, C. Cachoncinlle, A. Khacef, R. Viladrosa, J. M. Pouvesle, C. B. Collins, and F. Davanloo. Source flash X compacte impulsionnelle à haut taux de répétition. (French) []. Annales de Physique (Paris, France), 19(C1):C1–167, October 1994. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1994/
07/19941938/19941938.html;   http://www.annphys.org/articles/
anphys/pdf/1994/07/19941938.pdf.

Regueiro:1988:TSH

[REI^{+}88] M. Nunez Regueiro, P. Esquinazi, M. A. Izbizky, C. Duran, D. Castello, and J. Luzuriaga. Tunneling systems in high temperature superconductors. Annales de Physique (Paris, France), 13(5):401–??, ???? 1988. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1988/05/anphys_1988__13_5_401_
0/anphys_1988__13_5_401_0.html;   http://www.annphys.org/
articles/anphys/pdf/1988/05/anphys_1988__13_5_401_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1988/05/anphys_
1988__13_5_401_0/anphys_1988__13_5_401_0.html.

Reinholz:2005:DOP

[Rei05] H. Reinholz. Dielectric and optical properties of dense plasmas. Annales de Physique (Paris, France), 30(4–5):1–??, ???? 2005. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/2005/04/ann042005/ann042005.html;   http://
www.annphys.org/articles/anphys/pdf/2005/04/ann042005.
pdf;   http://www.annphys.org/articles/anphys/ref/2005/
04/ann042005/ann042005.html.

Remaud:1996:PPM

[Rem96] Bernard Remaud. Pseudo-particle models for nuclear reactions with heavy ions. Annales de Physique (Paris, France), 21(5):503– ??, ???? 1996. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/1996/05/ann05031996/ann05031996.html;   http:
//www.annphys.org/articles/anphys/pdf/1996/05/ann05031996.
pdf.

Renard:1992:SLT

[Ren92] F. M. Renard. Standardissimo. Les limitations théoriques du Modèle Standard. Quelles réponses y apporter?. (French) []. Annales de Physique (Paris, France), 17(4):301–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1992/04/anphys_1992__17_4_301_0/anphys_1992__17_4_301_
0.html;   http://www.annphys.org/articles/anphys/pdf/1992/
04/anphys_1992__17_4_301_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1992/04/anphys_1992__17_4_301_0/anphys_
1992__17_4_301_0.html.

Reulet:1995:SMO

[Reu95] Bertrand Reulet. Susceptibilitée magnéetique orbitale et conductance AC d’anneaux méesoscopiques isolées. (French) [Orbital magnetic susceptibility and AC conductance of isolated mesoscopic rings]. Annales de Physique (Paris, France), 20(3):233–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1995/03/ann03011995/ann03011995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/03/ann03011995.pdf.

Reuss:2000:BNC

[Reu00] P. Reuss. Base de neutronique: coefficients de réactivité dans les réacteurs nucléaires. (French) [Basic neutronics: reactivity coefficients in nuclear reactors]. Annales de Physique (Paris, France), 25(2):67–??, ???? 2000. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/2000/02/chap2/chap2.html;   http:/
/www.annphys.org/articles/anphys/pdf/2000/02/chap2.pdf.

Reynaud:1990:IRD

[Rey90] S. Reynaud. Introduction à la réduction du bruit quantique. (French) []. Annales de Physique (Paris, France), 15(2):63–??, ???? 1990. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1990/02/anphys_1990__15_2_63_0/anphys_1990__15_2_63_0.
html;   http://www.annphys.org/articles/anphys/pdf/1990/
02/anphys_1990__15_2_63_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1990/02/anphys_1990__15_2_63_0/anphys_
1990__15_2_63_0.html.

Reynaud:1985:CPN

[RH85] S. Reynaud and A. Heidmann. Can photon noise be reduced? Annales de Physique (Paris, France), 10(3):227–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/03/anphys_1985__10_3_227_0/anphys_1985__10_3_227_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
03/anphys_1985__10_3_227_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1985/03/anphys_1985__10_3_227_0/anphys_
1985__10_3_227_0.html.

Riboulet:1997:LSU

[Rib97] G. Riboulet. Les lasers à solides UV à impulsions courtes. (French) []. Annales de Physique (Paris, France), 22(C1):C1–33, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1997/07/19972205/19972205.html;   http://www.annphys.org/
articles/anphys/pdf/1997/07/19972205.pdf.

Rigaux:1995:SGD

[Rig95] Claudette Rigaux. Spin glass dynamics in semimagnetic semiconductors. Annales de Physique (Paris, France), 20(4):445–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1995/04/ann04021995/ann04021995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/04/ann04021995.pdf.

Robert:1997:DEP

[RKCP97] E. Robert, A. Khacef, C. Cachoncinlle, and J. M. Pouvesle. Dépôt d’énergie par rayonnement X dans les gaz rares à haute pression. (French) []. Annales de Physique (Paris, France), 22(C1):C1–135, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972229/
19972229.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972229.pdf.

Rey:1994:IEX

[RMA94] G. Rey, A. Mysyrowicz, and A. Antonetti. Ionisation et émission X d’un jet puisé d’argon irradié par des impulsions laser femtosecondes. (French) []. Annales de Physique (Paris, France), 19(C1):C1–171, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941940/
19941940.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941940.pdf.

Ranninger:1988:SML

[RMR88] J. Ranninger, R. Micnas, and S. Robaszkiewicz. Superconductivity of a mixture of local pairs and quasi free electrons. Annales de Physique (Paris, France), 13(5):455–??, ???? 1988. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/

1988/05/anphys_1988__13_5_455_0/anphys_1988__13_5_455_
0.html;   http://www.annphys.org/articles/anphys/pdf/1988/
05/anphys_1988__13_5_455_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1988/05/anphys_1988__13_5_455_0/anphys_
1988__13_5_455_0.html.

Romanelli:2007:NAA

[RMT^{+}07] M. Romanelli, I. Maurin, P. Todorov, Chia-Hua Chan, and D. Bloch. A 2D nanosphere array for atomic spectroscopy. Annales de Physique (Paris, France), 32(2–3):127–??, ???? 2007. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2007/
02/ann02200724/ann02200724.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200724.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200724/
ann02200724.html.

Robieux:1985:IAK

[Rob85] J. Robieux. Influence d’Alfred Kastler sur les applications des sciences de l’optique. (French) []. Annales de Physique (Paris, France), 10(6):613–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1985/06/anphys_1985__10_6_613_
0/anphys_1985__10_6_613_0.html;   http://www.annphys.org/
articles/anphys/pdf/1985/06/anphys_1985__10_6_613_0.pdf.

Rocard:1985:LPE

[Roc85] Y. Rocard. Le laboratoire de physique de l’E.N.S. à l’époque de Kastler. (French) []. Annales de Physique (Paris, France), 10(6):541–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1985/06/anphys_1985__10_6_541_
0/anphys_1985__10_6_541_0.html;   http://www.annphys.org/
articles/anphys/pdf/1985/06/anphys_1985__10_6_541_0.pdf.

Romanini:1995:SAQ

[Rom95] D. Romanini. Spectroscopie d’absorption quantitative à très haute sensibilité : ‘ring down’ cavité et absorption intracavité laser. (French) []. Annales de Physique (Paris, France), 20(5–6):655–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1995/05/ann05591995/ann05591995.html;   http://www.
annphys.org/articles/anphys/pdf/1995/05/ann05591995.pdf.

Rousseaux:1992:ISX

[Rou92] F. Rousseaux. Intérêt des sources X pour la production en microélectronique. (French) []. Annales de Physique (Paris, France), ??(??):133–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921726/
19921726.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921726.pdf.

Razet:1995:CCH

[RP95a] A. Razet and S. Picard. Calculs de constantes hyperfines de quelques transitions de la molécule d’iode ^{129}I_{2}. (French) []. Annales de Physique (Paris, France), 20(5–6):633–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05471995/ann05471995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05471995.pdf.

Razet:1995:NFI

[RP95b] A. Razet and S. Picard. Nouvelles formules d’interpolation concernant les constantes hyperfines de l’iode ^{127}I_{2}. (French) []. Annales de Physique (Paris, France), 20(5–6):583–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05221995/ann05221995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05221995.pdf.

Roy-Stephan:1990:DNE

[RS90] Michèle Roy-Stéphan. Le delta dans les noyaux. Expériences. (French) []. Annales de Physique (Paris, France), 15(5):393–??, ???? 1990. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1990/05/anphys_1990__15_5_393_0/anphys_1990__15_5_
393_0.html;   http://www.annphys.org/articles/anphys/pdf/
1990/05/anphys_1990__15_5_393_0.pdf;   http://www.annphys.
org/articles/anphys/ref/1990/05/anphys_1990__15_5_393_
0/anphys_1990__15_5_393_0.html.

Ryter:1991:CCC

[Ryt91] C. Ryter. Carbonaceous compounds in carbon stars and planetary nebulae. Annales de Physique (Paris, France), 16(4):507–??, ????

1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1991/04/anphys_1991__16_4_507_0/anphys_1991__16_4_507_
0.html;   http://www.annphys.org/articles/anphys/pdf/1991/
04/anphys_1991__16_4_507_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1991/04/anphys_1991__16_4_507_0/anphys_
1991__16_4_507_0.html.

Schirmann:1997:PIL

[SA97] D. Schirmann and M. André. Le projet d’installation laser mégajoule (lmj) du CEA-DAM. (French) []. Annales de Physique (Paris, France), 22(C1):C1–3, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972201/
19972201.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972201.pdf.

Szriftgiser:1995:CAF

[SAD^{+}95] P. Szriftgiser, M. Arndt, P. Desbiolles, A. Steane, and J. Dalibard. Cavités atomiques. (French) []. Annales de Physique (Paris, France), 20(5–6):681–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.

annphys.org/articles/anphys/abs/1995/05/ann05611995/ann05611995.
html;   http://www.annphys.org/articles/anphys/pdf/1995/
05/ann05611995.pdf.

Safi:1997:PFQ

[Saf97] Inès Safi. Propriétés d’un fil quantique connecté à des fils de mesure. (French) [Transport through a quantum wire connected to measuring leads]. Annales de Physique (Paris, France), 22(5):463– ??, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/1997/05/ann0501/ann0501.html;   http://www.
annphys.org/articles/anphys/pdf/1997/05/ann0501.pdf.

Salam:1985:PAK

[Sal85] A. Salam. Professor Alfred Kastler and the International Centre for Theoretical Physics. Annales de Physique (Paris, France), 10(6):595–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1985/06/anphys_1985__10_6_595_
0/anphys_1985__10_6_595_0.html;   http://www.annphys.org/
articles/anphys/pdf/1985/06/anphys_1985__10_6_595_0.pdf.

Salin:1992:LFU

[Sal92] F. Salin. Les lasers femtosecondes ultra-intenses. Annales de Physique (Paris, France), ??(??):91–??, ???? 1992. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1992/
07/19921721/19921721.html;   http://www.annphys.org/articles/
anphys/pdf/1992/07/19921721.pdf.

Sauneuf:1992:OXA

[Sau92] R. Sauneuf. Les optiques XUV adaptées à l’étude des plasmas créés par laser. (French) []. Annales de Physique (Paris, France), ??(??):193–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921739/
19921739.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921739.pdf.

Sauty:1994:VJA

[Sau94] C. Sauty. Des vents et des jets astrophysiques. (French) []. Annales de Physique (Paris, France), 19(5):459–??, ???? 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1994/05/anphys_1994__19_5_459_0/anphys_1994__19_5_459_
0.html;   http://www.annphys.org/articles/anphys/pdf/1994/
05/anphys_1994__19_5_459_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1994/05/anphys_1994__19_5_459_0/anphys_
1994__19_5_459_0.html.

Sauteret:1995:GAA

[Sau95] C. Sauteret. Génération et amplification au niveau de 100 TW d’impulsions femtosecondes. (French) []. Annales de Physique (Paris, France), 20(5–6):687–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05621995/ann05621995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05621995.pdf.

Sahal-Brechot:1984:APD

[SB84] S. Sahal-Bréchot. Atomic physics developments for interpretation of spectral line polarization measurements in astrophysics. Annales de Physique (Paris, France), 9(4):705–??, ???? 1984. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1984/
04/anphys_1984__9_4_705_0/anphys_1984__9_4_705_0.html;
 http://www.annphys.org/articles/anphys/pdf/1984/04/anphys_
1984__9_4_705_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1984/04/anphys_1984__9_4_705_0/anphys_1984_
_9_4_705_0.html.

Silvestre-Brac:1988:CCP

[SB88] B. Silvestre-Brac. Corrélations à courte portée : systèmes à peu de particules. (French) []. Annales de Physique (Paris, France), 13(2):69–??, ???? 1988. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1988/02/anphys_1988__13_2_69_0/anphys_
1988__13_2_69_0.html;   http://www.annphys.org/articles/
anphys/pdf/1988/02/anphys_1988__13_2_69_0.pdf;   http:
//www.annphys.org/articles/anphys/ref/1988/02/anphys_1988_
_13_2_69_0/anphys_1988__13_2_69_0.html.

Salin:1992:SLP

[SBE92] F. Salin, G. Brassart, and F. Estable. Source laser pour la production de rayons X de forte puissance moyenne par interaction laser-matière. (French) []. Annales de Physique (Paris, France), ??(??):191–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921738/
19921738.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921738.pdf.

Schuller:1984:OCR

[Sch84] F. Schuller. Optical collisions and redistribution in the presence of an intense driving field. Annales de Physique (Paris, France), 9(4):609–??, ???? 1984. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1984/04/anphys_1984__9_4_609_0/anphys_
1984__9_4_609_0.html;   http://www.annphys.org/articles/
anphys/pdf/1984/04/anphys_1984__9_4_609_0.pdf;   http:
//www.annphys.org/articles/anphys/ref/1984/04/anphys_1984_
_9_4_609_0/anphys_1984__9_4_609_0.html.

Schearer:1985:OPH

[Sch85] L. D. Schearer. Optically pumped helium — a spin dependent probe. Annales de Physique (Paris, France), 10(6):845–??, ????

1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/06/anphys_1985__10_6_845_0/anphys_1985__10_6_845_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
06/anphys_1985__10_6_845_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1985/06/anphys_1985__10_6_845_0/anphys_
1985__10_6_845_0.html.

Schirmann:1992:LDX

[Sch92] D. Schirmann. L’instrumentation dans le domaine X-UV. (French) [instrumentation in the X-UV domain]. Annales de Physique (Paris, France), ??(??):109–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921723/
19921723.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921723.pdf.

Schapira:2000:DAD

[Sch00] J. P. Schapira. Les differents aspects du cycle actuel, de l’extraction de l’uranium au stockage des déchets. (French) [different aspects of the current cycle of the extraction of uranium for waste storage]. Annales de Physique (Paris, France), 25 (2):175–??, ???? 2000. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/2000/02/chap6/chap6.html;   http:/
/www.annphys.org/articles/anphys/pdf/2000/02/chap6.pdf.

Schreck:2003:MUG

[Sch03] F. Schreck. Mixtures of ultracold gases: Fermi sea and Bose–Einstein condensate of lithium isotopes. Annales de Physique (Paris, France), 28(2):1–??, ???? 2003. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2003/
02/ann022003/ann022003.html;   http://www.annphys.org/
articles/anphys/pdf/2003/02/ann022003.pdf.

Schwartz:2008:SSR

[Sch08] S. Schwartz. Solid-state ring laser gyroscope. Annales de Physique (Paris, France), 33(1–2):1–??, ???? 2008. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2008/
01/ann012008/ann012008.html;   http://www.annphys.org/
articles/anphys/pdf/2008/01/ann012008.pdf;   http://www.

annphys.org/articles/anphys/ref/2008/01/ann012008/ann012008.
html.

Sentis:1994:GRU

[SCR^{+}94] M. L. Sentis, F. Chazaud, H. Rigneault, W. I. Marine, and A. Paramonov. Génération de rayonnements UV et VUV par décharge de surface. (French) []. Annales de Physique (Paris, France), 19(C1):C1–1, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941901/
19941901.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941901.pdf.

Serot:1993:TPP

[SCSN93] Olivier Serot, Nicolae Carjan, Dan Strottman, and Michael Martin Nieto. Tunneling dans les potentiels à plusieurs minima. (French) []. Annales de Physique (Paris, France), 18(3):275–??, ???? 1993. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1993/03/anphys_1993__18_3_275_0/anphys_1993__18_3_275_
0.html;   http://www.annphys.org/articles/anphys/pdf/1993/
03/anphys_1993__18_3_275_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1993/03/anphys_1993__18_3_275_0/anphys_
1993__18_3_275_0.html.

Senellart:2003:PMR

[Sen03] P. Senellart. Polaritons de microcavité: relaxation spontanée et stimulée. (French) [Microcavity polaritons: spontaneous and stimulated relaxation]. Annales de Physique (Paris, France), 28(4):1– ??, ???? 2003. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/2003/04/ann042003/ann042003.html;   http://
www.annphys.org/articles/anphys/pdf/2003/04/ann042003.
pdf.

Series:1985:SDR

[Ser85] G. W. Series. Studies in double resonance and optical pumping. Annales de Physique (Paris, France), 10(6):553–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/06/anphys_1985__10_6_553_0/anphys_1985__10_6_553_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
06/anphys_1985__10_6_553_0.pdf;   http://www.annphys.org/

articles/anphys/ref/1985/06/anphys_1985__10_6_553_0/anphys_
1985__10_6_553_0.html.

Stehle:1984:HLP

[SF84] C. Stehlé and N. Feautrier. Hydrogen line profiles at low densities. Annales de Physique (Paris, France), 9(4):697–??, ???? 1984. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1984/
04/anphys_1984__9_4_697_0/anphys_1984__9_4_697_0.html;
 http://www.annphys.org/articles/anphys/pdf/1984/04/anphys_
1984__9_4_697_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1984/04/anphys_1984__9_4_697_0/anphys_1984_
_9_4_697_0.html.

Stauter:1994:CPD

[SFB94] C. Stauter, J. Fontaine, and A. Biernaux. Contrôle de procédés durant le micro-usinage de céramique par laser à excimère. (French) []. Annales de Physique (Paris, France), 19(C1):C1–253, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1994/07/19941957/19941957.html;   http://www.annphys.org/
articles/anphys/pdf/1994/07/19941957.pdf.

Stauter:1995:MUPa

[SFE95] C. Stauter, J. Fontaine, and Th. Engel. Micro-usinage par laser à impulsions courtes : contrôle de procédé et applications. (French) []. Annales de Physique (Paris, France), 20(5–6):567–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1995/05/ann05141995/ann05141995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05141995.pdf.

Stauter:1997:MUPb

[SGFE97] C. Stauter, P. Gérard, J. Fontaine, and T. Engel. Micro-usinage par impulsions laser UV : diagnostic d’interaction. (French) []. Annales de Physique (Paris, France), 22(C1):C1–249, ???? 1997. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1997/
07/19972248/19972248.html;   http://www.annphys.org/articles/
anphys/pdf/1997/07/19972248.pdf.

Stehle:1997:LEF

[SGZ97] M. Stehle, B. Godard, and D. Zahorski. Laser à excimère de forte puissance pour le recuit du silicium : perspectives d’un laser de 50J. (French) []. Annales de Physique (Paris, France), 22(C1):C1–31, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972204/
19972204.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972204.pdf.

Salvan:1988:STM

[SHDT88] F. Salvan, A. Humbert, P. Dumas, and F. Thibaudau. Scanning tunneling microscopy (S.T.M.) of semiconductor surfaces and metal-semiconductor interfaces. Annales de Physique (Paris, France), 13(3):133–??, ???? 1988. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1988/03/anphys_1988__13_3_133_
0/anphys_1988__13_3_133_0.html;   http://www.annphys.org/
articles/anphys/pdf/1988/03/anphys_1988__13_3_133_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1988/03/anphys_
1988__13_3_133_0/anphys_1988__13_3_133_0.html.

Sommers:1985:AP

[SK85] C. Sommers and M. Kléman. Avant-propos. Annales de Physique (Paris, France), 10(1):6–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1985/
01/anphys_1985__10_1_R6_0/anphys_1985__10_1_R6_0.html;
 http://www.annphys.org/articles/anphys/pdf/1985/01/anphys_
1985__10_1_R6_0.pdf.

Schanne-Klein:1995:MDT

[SKGRF95] M. C. Schanne-Klein, M. Ghanassi, D. Ricard, and C. Flyzanis. Mécanismes et dépendance en taille des non-linéarités optiques résonances des verres dopés par nanocristaux de Cd(S,Se). (French) []. Annales de Physique (Paris, France), 20(5–6):591–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1995/05/ann05261995/ann05261995.html;   http://www.
annphys.org/articles/anphys/pdf/1995/05/ann05261995.pdf.

ski:1985:AKM

[ski85] T. Skali ski. Alfred Kastler in memoriam. Annales de Physique (Paris, France), 10(6):601–??, ???? 1985. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/06/anphys_1985__10_6_601_0/anphys_1985__10_6_601_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
06/anphys_1985__10_6_601_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1985/06/anphys_1985__10_6_601_0/anphys_
1985__10_6_601_0.html.

Scherrer:1995:MIT

[SL95] B. Scherrer and M. Lefebvre. Mesures instantanées de températures et de vitesse par D.R.A.S.C. (French) []. Annales de Physique (Paris, France), 20(5–6):605–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.

annphys.org/articles/anphys/abs/1995/05/ann05331995/ann05331995.
html;   http://www.annphys.org/articles/anphys/pdf/1995/
05/ann05331995.pdf.

Soullie:1994:OBF

[SLI^{+}94] G. Soullié, F. Legrand, M. Idir, A. Mirone, P. Dhez, and P. Chevallier. Optiques de Bragg–Fresnel pour la microfocalisation de rayons X. (French) []. Annales de Physique (Paris, France), 19(C1):C1–169, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941939/
19941939.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941939.pdf.

Spach:1989:MAB

[SM89] Gérard Spach and Liliane Merle. La membrane ancestrale : barrière asymétrique à l’origine de l’activité optique dans les milieux vivants ?. (French) []. Annales de Physique (Paris, France), 14(2):235–??, ???? 1989. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1989/02/anphys_1989__14_2_235_
0/anphys_1989__14_2_235_0.html;   http://www.annphys.org/
articles/anphys/pdf/1989/02/anphys_1989__14_2_235_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1989/02/anphys_
1989__14_2_235_0/anphys_1989__14_2_235_0.html.

Salamero:1997:SCG

[SM97] Y. Salamero and P. Millet. Spectroscopie et cinétique des gaz rares et leurs mélanges sous excitation multiphotonique. (French) []. Annales de Physique (Paris, France), 22(C1):C1–97, ???? 1997. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1997/
07/19972217/19972217.html;   http://www.annphys.org/articles/
anphys/pdf/1997/07/19972217.pdf.

Souletie:1985:SGT

[Sou85] J. Souletie. Spin glasses — a transition in plain ordinary space. Annales de Physique (Paris, France), 10(1):69–??, ???? 1985. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1985/
01/anphys_1985__10_1_69_0/anphys_1985__10_1_69_0.html;
 http://www.annphys.org/articles/anphys/pdf/1985/01/anphys_
1985__10_1_69_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1985/01/anphys_1985__10_1_69_0/anphys_1985_
_10_1_69_0.html.

Souletie:1994:RHE

[Sou94] J. Souletie. Renormalisation hiérarchique et exponentielles étirées. (French) []. Annales de Physique (Paris, France), 19(6):765–??, ???? 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1994/06/anphys_1994__19_6_765_0/anphys_1994__19_6_
765_0.html;   http://www.annphys.org/articles/anphys/pdf/
1994/06/anphys_1994__19_6_765_0.pdf;   http://www.annphys.
org/articles/anphys/ref/1994/06/anphys_1994__19_6_765_
0/anphys_1994__19_6_765_0.html.

Svatos:1992:MIR

[SPJP92] J. Svatos, F. Polack, D. Joyeux, and D. Phalippou. Mesure de l’indice de réfraction par interférométrie en X mous. (French) []. Annales de Physique (Paris, France), ??(??):195–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1992/07/19921740/19921740.html;   http://www.annphys.org/
articles/anphys/pdf/1992/07/19921740.pdf.

Schwob:1995:SLC

[SRF95] C. Schwob, C. Richy, and C. Fabre. Spectroscopie en lumière comprimée. (French) []. Annales de Physique (Paris, France), 20 (5–6):581–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1995/05/ann05211995/ann05211995.html;
 http://www.annphys.org/articles/anphys/pdf/1995/05/ann05211995.
pdf.

Stamm:1984:MSA

[ST84] R. Stamm and B. Talin. Microfield simulation applied to the calculation of Stark profiles. Annales de Physique (Paris, France), 9(4):687–??, ???? 1984. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/1984/04/anphys_1984__9_4_687_0/anphys_
1984__9_4_687_0.html;   http://www.annphys.org/articles/
anphys/pdf/1984/04/anphys_1984__9_4_687_0.pdf;   http:
//www.annphys.org/articles/anphys/ref/1984/04/anphys_1984_
_9_4_687_0/anphys_1984__9_4_687_0.html.

Sungur:2007:MMP

[STC^{+}07] E. Sungur, G. Taupier, O. Crégut, D. Gindre, L. Pieuchot, J.-L. Evrard, A.-C. Schmit, L. Mager, and K. D. Dorkenoo. Multiphotonic microscopy as a processing tool for optical data storage and imaging in biophysics. Annales de Physique (Paris, France), 32(2–3):147–??, ???? 2007. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2007/
02/ann02200729/ann02200729.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200729.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200729/
ann02200729.html.

Stenholm:1985:CLQ

[Ste85] S. Stenholm. On the classical limit of quantum electrodynamics. Annales de Physique (Paris, France), 10(6):817–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/06/anphys_1985__10_6_817_0/anphys_1985__10_6_817_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
06/anphys_1985__10_6_817_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1985/06/anphys_1985__10_6_817_0/anphys_
1985__10_6_817_0.html.

Stehle:1992:EAD

[Ste92] M. Stehle. État d’avancement du programme Eurêka EU205 ‘laser à excimères de fortes puissances moyennes pour usage industriel’. (French) []. Annales de Physique (Paris, France), ??(??):9–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1992/07/19921702/19921702.html;   http://www.annphys.org/
articles/anphys/pdf/1992/07/19921702.pdf.

Stroke:1985:BBK

[Str85] H. H. Stroke. Bitter, Brossel, Kastler and beyond: thirty years of U.S.–French collaboration in optical resonance. Annales de Physique (Paris, France), 10(6):589–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/06/anphys_1985__10_6_589_0/anphys_1985__10_6_589_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
06/anphys_1985__10_6_589_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1985/06/anphys_1985__10_6_589_0/anphys_
1985__10_6_589_0.html.

Suraud:1991:LEN

[Sur91] E. Suraud. L’équation d’état nucléaire. (French) []. Annales de Physique (Paris, France), 16(2):193–??, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1991/02/anphys_1991__16_2_193_0/anphys_1991__16_2_193_
0.html;   http://www.annphys.org/articles/anphys/pdf/1991/
02/anphys_1991__16_2_193_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1991/02/anphys_1991__16_2_193_0/anphys_
1991__16_2_193_0.html.

Suraud:1996:KEH

[Sur96] E. Suraud. On kinetic equations in heavy-ion physics. Annales de Physique (Paris, France), 21(5):461–??, ???? 1996. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1996/
05/ann05021996/ann05021996.html;   http://www.annphys.
org/articles/anphys/pdf/1996/05/ann05021996.pdf.

Suzuki:1984:SRT

[Suz84] T. Suzuki. Sum rules and their applications in nuclear physics. Annales de Physique (Paris, France), 9(3):535–??, ???? 1984. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1984/
03/anphys_1984__9_3_535_0/anphys_1984__9_3_535_0.html;

 http://www.annphys.org/articles/anphys/pdf/1984/03/anphys_
1984__9_3_535_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1984/03/anphys_1984__9_3_535_0/anphys_1984_
_9_3_535_0.html.

Sipp:1987:EDTa

[SV87a] B. Sipp and R. Voltz. Étude du transfert d’excitation dans les milieux moléculaires fluides. I. Description théorique. (French) []. Annales de Physique (Paris, France), 12(1):19–??, ???? 1987. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1987/
01/anphys_1987__12_1_19_0/anphys_1987__12_1_19_0.html;
 http://www.annphys.org/articles/anphys/pdf/1987/01/anphys_
1987__12_1_19_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1987/01/anphys_1987__12_1_19_0/anphys_1987_
_12_1_19_0.html.

Sipp:1987:EDTb

[SV87b] B. Sipp and R. Voltz. Étude du transfert d’excitation dans les milieux moléculaires fluides. II. Aspects expérimentaux. (French) []. Annales de Physique (Paris, France), 12(1):39–??, ???? 1987. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1987/
01/anphys_1987__12_1_39_0/anphys_1987__12_1_39_0.html;
 http://www.annphys.org/articles/anphys/pdf/1987/01/anphys_
1987__12_1_39_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1987/01/anphys_1987__12_1_39_0/anphys_1987_
_12_1_39_0.html.

Subtil:1994:RTP

[SVDL^{+}94] J.-L. Subtil, C. Vincent-Donnet, P. Laporte, R. Reininger, and P. Gürtler. Relation température-profil d’émission dans la luminescence de l’exciplexe Kr^{∗} Ar. (French) []. Annales de Physique (Paris, France), 19(C1):C1–173, October 1994. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1994/
07/19941941/19941941.html;   http://www.annphys.org/articles/
anphys/pdf/1994/07/19941941.pdf.

Schaaf:1998:IDA

[SVS98] P. Schaaf, J.-C. Voegel, and B. Senger. Irreversible deposition/adsorption processes on solid surfaces. Annales de Physique (Paris, France), 23(6):1–??, ???? 1998. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1998/
06/ann061998/ann061998.html;   http://www.annphys.org/
articles/anphys/pdf/1998/06/ann061998.pdf.

Thomann:1994:ESF

[TBLH^{+}94] A. L. Thomann, C. Boulmer-Leborgne, J. Hermann, C. Andreazza-Vignolle, P. Andreazza, G. Blondiaux, and F. Degas. Étude de la synthèse de films produits par irradiation laser-UV directe. (French) []. Annales de Physique (Paris, France), 19(C1):C1–175, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/1994/07/19941942/19941942.html;   http://www.
annphys.org/articles/anphys/pdf/1994/07/19941942.pdf.

Tourrenc:1985:ETD

[TD85] Ph. Tourrenc and N. Deruelle. Effects of the time delays in a non linear pendular Fabry–Perot. Annales de Physique (Paris, France), 10(3):241–??, ???? 1985. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/03/anphys_1985__10_3_241_0/anphys_1985__10_3_241_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
03/anphys_1985__10_3_241_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1985/03/anphys_1985__10_3_241_0/anphys_
1985__10_3_241_0.html.

Torchet:1989:SSC

[TFdFR89] G. Torchet, J. Farges, M. F. de Feraudy, and B. Raoult. Structural study of CH_{4}, CO_{2} and H_{2}O clusters containing from several tens to several thousands of molecules. Annales de Physique (Paris, France), 14(2):245–??, ???? 1989. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1989/02/anphys_1989__14_2_245_0/anphys_1989__14_2_245_
0.html;   http://www.annphys.org/articles/anphys/pdf/1989/
02/anphys_1989__14_2_245_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1989/02/anphys_1989__14_2_245_0/anphys_
1989__14_2_245_0.html.

Tchelnokov:1995:DSS

[TLG95] A. V. Tchelnokov, J.-M. Lourtioz, and P. Gavrilovic. Dynamique saptiale et spectrale de l’émission optique de lasers/amplificateurs à semiconducteurs à large ouverture. (French) []. Annales de Physique (Paris, France), 20(5–6):637–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05491995/ann05491995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05491995.pdf.

Torres:2002:TBQ

[Tor02] J. Torrès. Transport et bruit quantique dans les fils mésoscopiques. (French) [Transport and quantum noise in mesoscopic wires]. Annales de Physique (Paris, France), 27(1):1–??, ???? 2002. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2002/
01/ann012002/ann012002.html;   http://www.annphys.org/
articles/anphys/pdf/2002/01/ann012002.pdf.

Torres:2004:EGS

[Tor04] J. Torres. Exaltation de la génération de second harmonique dans les cristaux photoniques à base de nitrures. (French) [Excitation of the generation of second harmonics in nitride-based photonic crystals]. Annales de Physique (Paris, France), 29(5):1–??, ???? 2004. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/2004/05/ann052004/ann052004.html;   http://
www.annphys.org/articles/anphys/pdf/2004/05/ann052004.
pdf.

Toschek:1985:LRF

[Tos85] P. E. Toschek. Lumino-réfrigération. (French) []. Annales de Physique (Paris, France), 10(6):761–??, ???? 1985. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/06/anphys_1985__10_6_761_0/anphys_1985__10_6_761_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/
06/anphys_1985__10_6_761_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1985/06/anphys_1985__10_6_761_0/anphys_
1985__10_6_761_0.html.

Toulouse:1985:ABP

[Tou85] G. Toulouse. André Blandin et la physique des verres de spin. Trois étés alpins : 1958, 1968, 1978. (French) []. Annales de Physique (Paris, France), 10(1):85–??, ???? 1985. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1985/
01/anphys_1985__10_1_85_0/anphys_1985__10_1_85_0.html;
 http://www.annphys.org/articles/anphys/pdf/1985/01/anphys_
1985__10_1_85_0.pdf;   http://www.annphys.org/articles/
anphys/ref/1985/01/anphys_1985__10_1_85_0/anphys_1985_
_10_1_85_0.html.

Troussel:1997:IMA

[TPD^{+}97] P. Troussel, M. Pichet, J. M. Dalmasso, M. Idir, T. Jalinaud, M. Ollivier, J. P. Le Breton, and R. Sauneuf. Imagerie micronique à l’aide de lentilles à zones de Fresnel. (French) []. Annales de Physique (Paris, France), 22(C1):C1–153, ???? 1997. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1997/
07/19972234/19972234.html;   http://www.annphys.org/articles/
anphys/pdf/1997/07/19972234.pdf.

Troussel:1994:EAD

[Tro94] P. Troussel. Étalonnages absolus dans le domaine X-UV. (French) []. Annales de Physique (Paris, France), 19(C1):C1–147, October 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1994/07/19941931/19941931.html;   http://www.annphys.org/
articles/anphys/pdf/1994/07/19941931.pdf.

Troussel:2002:QDX

[Tro02] Ph. Troussel. Quelques diagnostics X pour les expériences d’interaction laser-matière. (French) [Some X diagnostics for laser–matter interaction experiments]. Annales de Physique (Paris, France), 27(3):1–??, ???? 2002. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://

www.annphys.org/articles/anphys/abs/2002/03/ann032002/
ann032002.html;   http://www.annphys.org/articles/anphys/
pdf/2002/03/ann032002.pdf.

Tignon:1995:MSS

[TVW^{+}95] J. Tignon, P. Voisin, J. Wainstain, C. Delalande, M. Voos, R. Houdré, U. Oesterle, and R. P. Stanley. Microcavité de semiconducteurs sous champ magnétique : de la règle d’or de Fermi au splitting de Rabi. (French) []. Annales de Physique (Paris, France), 20(5–6):541–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05091995/ann05091995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05091995.pdf.

vanAmerongen:2008:ODB

[vA08] A. H. van Amerongen. One-dimensional Bose gas on an atom chip. Annales de Physique (Paris, France), 33(3):1–??, ???? 2008. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/2008/02/ann032008/ann032008.html;   http://
www.annphys.org/articles/anphys/pdf/2008/02/ann032008.
pdf;   http://www.annphys.org/articles/anphys/ref/2008/
02/ann032008/ann032008.html.

Vanier:1985:POE

[Van85] J. Vanier. Le pompage optique et les étalons atomiques de fréquence. (French) []. Annales de Physique (Paris, France), 10(6):749–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1985/06/anphys_1985__10_6_749_
0/anphys_1985__10_6_749_0.html;   http://www.annphys.org/
articles/anphys/pdf/1985/06/anphys_1985__10_6_749_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1985/06/anphys_
1985__10_6_749_0/anphys_1985__10_6_749_0.html.

VanTiggelen:1989:CSD

[Van89] R. J. Van Tiggelen. Cinétique et structure des détonations. (French) []. Annales de Physique (Paris, France), 14(6):605–??, ???? 1989. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1989/06/anphys_1989__14_6_605_0/anphys_1989__14_6_
605_0.html;   http://www.annphys.org/articles/anphys/pdf/
1989/06/anphys_1989__14_6_605_0.pdf.

deWiele:2001:RMA

[Van01] J. Van de Wiele. Rotations et moments angulaires en mécanique quantique. (French) [rotations and angular momentum in quantum mechanics]. Annales de Physique (Paris, France), 26(6):1–??, ???? 2001. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/2001/06/annales06/annales06.html;   http://
www.annphys.org/articles/anphys/pdf/2001/06/annales06.
pdf.

vonBordwehr:1989:HXR

[vB89] R. Stumm von Bordwehr. A history of X-ray absorption fine structure. Annales de Physique (Paris, France), 14(4):377–??, ???? 1989. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1989/04/anphys_1989__14_4_377_0/anphys_1989__14_4_377_
0.html;   http://www.annphys.org/articles/anphys/pdf/1989/
04/anphys_1989__14_4_377_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1989/04/anphys_1989__14_4_377_0/anphys_
1989__14_4_377_0.html.

Verlot:2007:OCH

[VCB^{+}07] P. Verlot, T. Caniard, T. Briant, P.-F. Cohadon, and A. Heidmann. Optomechanical coupling in high-finesse cavities: towards the observation of quantum effects. Annales de Physique (Paris, France), 32(2–3):167–??, ???? 2007. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2007/
02/ann02200734/ann02200734.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200734.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200734/
ann02200734.html.

Verweij:1988:TOR

[Ver88] H. Verweij. Thermodynamics of the oxidation reaction of YBa_{2}Cu_{3}O_{x}. Annales de Physique (Paris, France), 13(5):349–??, ???? 1988. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1988/05/anphys_1988__13_5_349_0/anphys_1988__13_5_
349_0.html;   http://www.annphys.org/articles/anphys/pdf/
1988/05/anphys_1988__13_5_349_0.pdf;   http://www.annphys.
org/articles/anphys/ref/1988/05/anphys_1988__13_5_349_
0/anphys_1988__13_5_349_0.html.

Volino:1997:TVEf

[VGG97] F. Volino, H. Gérard, and G. Gebel. Théorie visco-élastique nonextensive VI. Application à un liquide formant une phase vitreuse: l’OrthoTerPhényl (OTP). (French) [Non-extensive visco-elastic theory VI. Application to a glass forming liquid: OrthoTerPhenyl (OTP)]. Annales de Physique (Paris, France), 22(1–2):181–??, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/

abs/1997/01/ann01081997/ann01081997.html;   http://www.
annphys.org/articles/anphys/pdf/1997/01/ann01081997.pdf.

Volino:1997:TVEb

[VGM97] F. Volino, H. Gérard, and S. Miachon. Théorie visco-élastique nonextensive II. Premiers tests expérimentaux de la théorie simplifiée à modes rotationnels. (French) [Non-extensive visco-elastic theory II. First experimental tests of the simple theory with rotational modes]. Annales de Physique (Paris, France), 22(1–2):43–??, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1997/01/ann01041997/ann01041997.html;   http://www.annphys.
org/articles/anphys/pdf/1997/01/ann01041997.pdf.

Vivien:1997:EPD

[VHBL97] C. Vivien, J. Hermann, and C. Boulmer-Leborgne. Étude de la propagation du plasma produit par ablation laser pour différentes cibles, dans le vide et dans différents gaz ambiants. (French) []. Annales de Physique (Paris, France), 22(C1):C1–133, ???? 1997. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1997/
07/19972228/19972228.html;   http://www.annphys.org/articles/
anphys/pdf/1997/07/19972228.pdf.

Villain:2003:MCM

[Vil03] J. Villain. Magnétisme des cristaux de molécules à grand spin. (French) [magnetism of crystals of giant-spin molecules]. Annales de Physique (Paris, France), 28(1):1–??, ???? 2003. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2003/
01/ann012003/ann012003.html;   http://www.annphys.org/
articles/anphys/pdf/2003/01/ann012003.pdf.

Vinet:1985:EOA

[Vin85] J.-Y. Vinet. Elasto optical antennas. Annales de Physique (Paris, France), 10(3):253–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1985/03/anphys_1985__10_3_253_
0/anphys_1985__10_3_253_0.html;   http://www.annphys.org/
articles/anphys/pdf/1985/03/anphys_1985__10_3_253_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1985/03/anphys_
1985__10_3_253_0/anphys_1985__10_3_253_0.html.

Vitel:1992:SVT

[Vit92] Y. Vitel. Source VUV à tube à éclairs. (French) []. Annales de Physique (Paris, France), ??(??):81–??, ???? 1992. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1992/
07/19921719/19921719.html;   http://www.annphys.org/articles/
anphys/pdf/1992/07/19921719.pdf.

Viladrosa:1997:SGR

[VKCP97] R. Viladrosa, A. Khacef, C. Cachoncinlle, and J. M. Pouvesle. SPHINX : générateur de rayons X pulsés, ultracompact à cadence élevée. (French) []. Annales de Physique (Paris, France), 22(C1):C1–77, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972214/
19972214.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972214.pdf.

Tiggelen:1989:EVC

[VL89] P. J. Van Tiggelen and J. C. Libouton. Évolution des variables chimiques et physiques à l’intérieur d’une maille de détonation. (French) []. Annales de Physique (Paris, France), 14(6):649–??, ???? 1989. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1989/06/anphys_1989__14_6_649_0/anphys_1989__14_6_
649_0.html;   http://www.annphys.org/articles/anphys/pdf/
1989/06/anphys_1989__14_6_649_0.pdf;   http://www.annphys.
org/articles/anphys/ref/1989/06/anphys_1989__14_6_649_
0/anphys_1989__14_6_649_0.html.

Vansteenkiste:1995:MAR

[VLL^{+}95] N. Vansteenkiste, A. Landraguin, G. Labeyrie, R. Kaiser, C. I. Westbrook, and A. Aspect. Miroir à atomes : réflexion spéculaire ou réflexion diffuse ?. (French) []. Annales de Physique (Paris, France), 20(5–6):595–??, ???? 1995. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1995/
05/ann05281995/ann05281995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05281995.pdf.

Voisin:1997:WSE

[Voi97] P. Voisin. Wannier–Stark effects in semiconductor superlattices. Annales de Physique (Paris, France), 22(6):681–??, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/1997/06/ann0601/ann0601.html;   http://www.
annphys.org/articles/anphys/pdf/1997/06/ann0601.pdf.

Volino:1997:PGF

[Vol97a] F. Volino. Présentation générale. (French) []. Annales de Physique (Paris, France), 22(1–2):3–??, ???? 1997. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1997/
01/ann01021997/ann01021997.html;   http://www.annphys.
org/articles/anphys/pdf/1997/01/ann01021997.pdf.

Volino:1997:TVEa

[Vol97b] F. Volino. Théorie visco-élastique non-extensive i. théorie simplifiée à modes rotationnels. (French) [Non-extensive visco-elastic theory I. Simple theory with rotational modes]. Annales de Physique (Paris, France), 22(1–2):7–??, ???? 1997. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1997/
01/ann01031997/ann01031997.html;   http://www.annphys.
org/articles/anphys/pdf/1997/01/ann01031997.pdf.

Volino:1997:TVEc

[Vol97c] F. Volino. Théorie visco-élastique non-extensive III. Théorie simplifiée à modes translationnels. (French) [Non-extensive viscoelastic theory III. Simple theory with translational modes]. Annales de Physique (Paris, France), 22(1–2):83–??, ???? 1997. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1997/
01/ann01051997/ann01051997.html;   http://www.annphys.
org/articles/anphys/pdf/1997/01/ann01051997.pdf.

Volino:1997:TVEd

[Vol97d] F. Volino. Théorie visco-élastique non-extensive IV. Premiers tests expérimentaux de la théorie simplifiée à modes translationnels. (French) [Non-extensive visco-elastic theory IV. First experimental tests of the simple theory with translational modes]. Annales de Physique (Paris, France), 22(1–2):121–??, ???? 1997. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1997/
01/ann01061997/ann01061997.html;   http://www.annphys.
org/articles/anphys/pdf/1997/01/ann01061997.pdf.

Volino:1997:TVEe

[Vol97e] F. Volino. Théorie visco-élastique non-extensive v. équations de diffusion, fonctions de distribution, viscosités et coefficients de diffusion, fonctions de corrélation. (French) [Non-extensive viscoelastic theory V. Diffusion equations, distribution functions viscosity and diffusion coefficients, correlation functions]. Annales de Physique (Paris, France), 22(1–2):159–??, ???? 1997. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1997/
01/ann01071997/ann01071997.html;   http://www.annphys.
org/articles/anphys/pdf/1997/01/ann01071997.pdf.

Vassoille:1985:EPM

[VP85] R. Vassoille and J. Perez. Études physiques des mélanges eaucryoprotecteurs. (French) []. Annales de Physique (Paris, France), 10(4):307–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.

org/articles/anphys/abs/1985/04/anphys_1985__10_4_307_
0/anphys_1985__10_4_307_0.html;   http://www.annphys.org/
articles/anphys/pdf/1985/04/anphys_1985__10_4_307_0.pdf;
 http://www.annphys.org/articles/anphys/ref/1985/04/anphys_
1985__10_4_307_0/anphys_1985__10_4_307_0.html.

Vanhaudenarde-Peochh:1995:DOP

[VPF95] A. Vanhaudenarde-Péoch’h and R. Frey. Différentes origines physiques de la variation d’indice de la réfraction obtenue lors du photo-noicissement des verres dopés au semi-conducteur. (French) []. Annales de Physique (Paris, France), 20(5–6):653–??, ???? 1995. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1995/05/ann05571995/ann05571995.html;   http://www.annphys.
org/articles/anphys/pdf/1995/05/ann05571995.pdf.

Wahi:1994:STH

[WBC^{+}94] T. Wahi, M. Bonnefoy, M. Chassevent, A. Fleury, J. J. Bonnet, X. Husson, and C. Hague. Spectrométrie à très haute résolution de l’émission K du néon. (French) []. Annales de Physique (Paris, France), 19(C1):C1–59, October 1994. CODEN ANPHAJ.

ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1994/07/19941918/
19941918.html;   http://www.annphys.org/articles/anphys/
pdf/1994/07/19941918.pdf.

Wuilleumier:1992:UCD

[WCB92] F. J. Wuilleumier, D. Cubaynes, and J.-M. Bizau. Utilisation combinée des deux sources de rayonnements laser et synchrotron : caractéristiques et résultats. (French) []. Annales de Physique (Paris, France), ??(??):197–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921741/
19921741.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921741.pdf.

Whitworth:1991:WDS

[Whi91]  Anthony Whitworth. Why do stars become giants?  An-

nales de Physique (Paris, France), 16(4):515–??, ???? 1991. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1991/04/anphys_1991__16_4_515_0/anphys_1991__16_4_515_
0.html;   http://www.annphys.org/articles/anphys/pdf/1991/
04/anphys_1991__16_4_515_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1991/04/anphys_1991__16_4_515_0/anphys_
1991__16_4_515_0.html.

Wiesenfeld:1994:AP

[WHPD94] L. Wiesenfeld, B. Houchmandzadeh, J. Peinke, and B. Douçot. Avant-propos. Annales de Physique (Paris, France), 19(6):R1, ???? 1994. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1994/06/anphys_1994__19_6_599_0/anphys_1994__19_6_
599_0.html;   http://www.annphys.org/articles/anphys/pdf/
1994/06/anphys_1994__19_6_599_0.pdf.

Wineland:1985:OPS

[WIB^{+}85] D. J. Wineland, W. M. Itano, J. C. Bergquist, J. J. Bollinger, and J. D. Prestage. Optical pumping of stored atomic ions. Annales de Physique (Paris, France), 10(6):737–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (elec-

tronic). URL http://www.annphys.org/articles/anphys/abs/
1985/06/anphys_1985__10_6_737_0/anphys_1985__10_6_737_
0.html;   http://www.annphys.org/articles/anphys/pdf/1985/

06/anphys_1985__10_6_737_0.pdf;   http://www.annphys.org/
articles/anphys/ref/1985/06/anphys_1985__10_6_737_0/anphys_
1985__10_6_737_0.html.

Wyart:2005:RAS

[Wya05] M. Wyart. On the rigidity of amorphous solids. Annales de Physique (Paris, France), 30(3):1–??, ???? 2005. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2005/
03/ann032005/ann032005.html;   http://www.annphys.org/
articles/anphys/pdf/2005/03/ann032005.pdf;   http://www.
annphys.org/articles/anphys/ref/2005/03/ann032005/ann032005.
html.

Wyon:1992:NMN

[Wyo92] Ch. Wyon. Nouveaux matériaux non linéaires pour la conversion de fréquence. (French) []. Annales de Physique (Paris, France), ??(??):151–??, ???? 1992. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1992/07/19921728/
19921728.html;   http://www.annphys.org/articles/anphys/
pdf/1992/07/19921728.pdf.

Xuan:2007:IKN

[XZS^{+}07] L. Le Xuan, C. Zhou, A. Slablab, D. Chauvat, and J.-F. Roch. Investigation of KTiOPO_{4} nanocrystals by means of secondharmonic light emission. Annales de Physique (Paris, France), 32(2–3):99–??, ???? 2007. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http://www.annphys.org/

articles/anphys/abs/2007/02/ann02200717/ann02200717.html;
 http://www.annphys.org/articles/anphys/pdf/2007/02/ann02200717.
pdf;   http://www.annphys.org/articles/anphys/ref/2007/
02/ann02200717/ann02200717.html.

Zeitoun:1997:LXU

[ZAC^{+}97] Ph. Zeitoun, F. Albert, A. Carillon, P. Jaeglé, G. Jamelot, A. Klisnick, D. Ros, B. Rus, and S. Sebban. Le laser X-UV au stade de ses premières applications. (French) []. Annales de Physique (Paris, France), 22(C1):C1–37, ???? 1997. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic). URL http:/

/www.annphys.org/articles/anphys/abs/1997/07/19972206/
19972206.html;   http://www.annphys.org/articles/anphys/
pdf/1997/07/19972206.pdf.

Zaghloul:1990:PTS

[ZB90] H. Zaghloul and H. A. Buckmaster. Properties of TEM standing waves with E||B. Annales de Physique (Paris, France), 15(1):21– ??, ???? 1990. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-

4838 (electronic). URL http://www.annphys.org/articles/
anphys/abs/1990/01/anphys_1990__15_1_21_0/anphys_1990_
_15_1_21_0.html;   http://www.annphys.org/articles/anphys/
pdf/1990/01/anphys_1990__15_1_21_0.pdf;   http://www.annphys.
org/articles/anphys/ref/1990/01/anphys_1990__15_1_21_0/
anphys_1990__15_1_21_0.html.

Zissis:1994:EUD

[ZD94] G. Zissis and J. J. Damelincourt. Emission UV des décharges capillaires mercure-gaz-rare basse pression. (French) []. Annales de Physique (Paris, France), 19(C1):C1–179, October 1994. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/1994/
07/19941944/19941944.html;   http://www.annphys.org/articles/
anphys/pdf/1994/07/19941944.pdf.

zuPutlitz:1985:ODR

[zP85] G. zu Putlitz. Optical double resonance and optical pumping in Heidelberg. Annales de Physique (Paris, France), 10(6):571–??, ???? 1985. CODEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838

(electronic). URL http://www.annphys.org/articles/anphys/
abs/1985/06/anphys_1985__10_6_571_0/anphys_1985__10_6_
571_0.html;   http://www.annphys.org/articles/anphys/pdf/
1985/06/anphys_1985__10_6_571_0.pdf;   http://www.annphys.
org/articles/anphys/ref/1985/06/anphys_1985__10_6_571_
0/anphys_1985__10_6_571_0.html.

Zhou:2007:TSP

[ZWJ^{+}07] C. Zhou, E. Wu, V. Jacques, J. Rabeau, F. Treussart, H. Zeng, S. Prawer, P. Grangier, and J.-F. Roch. Triggered singlephoton source based on photoluminescence of single nickelrelated colour centres in CVD-grown nanodiamonds. Annales de Physique (Paris, France), 32(2–3):159–??, ???? 2007. CO-DEN ANPHAJ. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL http://www.annphys.org/articles/anphys/abs/2007/
02/ann02200732/ann02200732.html;   http://www.annphys.
org/articles/anphys/pdf/2007/02/ann02200732.pdf;   http:
//www.annphys.org/articles/anphys/ref/2007/02/ann02200732/
ann02200732.html.